A Complete Bibliography of Publications in Isis, 1930–1939

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.07

Title word cross-reference

b [518]. ϵ̀πίσημoι [57]. πóλϵις [57].

-1455 [201]. -1619 [479].

1 [359]. 12th [866]. 1361 [799]. 13th [866, 465]. 1450 [1003]. 1583 [562]. 1642 [51]. 1650 [562]. 1670 [1147]. 1672 [759]. 1673 [1223]. 16th [623]. 1708 [943]. 1739 [432]. 1740 [397]. 1744 [764]. 1752 [880]. 1787 [586]. 1794 [985]. 17th [623, 1214, 1213, 642]. 1800 [1191, 167, 1036]. 1831 [686]. 1850 [1068]. 1855 [64]. 1860 [561]. 1865 [16]. 1875 [1039]. 18th [1019]. 1900 [531]. 1904 [622]. 1905 [1038]. 1925 [1067]. 1928 [40]. 1929 [64, 729]. 1930 [191, 193, 374, 246]. 1933 [409, 453, 435, 504]. 1934 [512, 532, 595, 987]. 1935 [636, 664, 852]. 1936 [731, 793, 829]. 1937 [1203, 1145, 1065]. 1938 [1149]. 1939 [1240].

27.XII.1931 [261].

31.X.1874 [261]. 3457 [1089].

444 [647]. 4966 [470].

6 [64].

700 [522].

817 [681]. 82 [356]. 860 [415].

90 [77].

âge [994, 658]. Âryabhatîya [101]. Âryabhata [101, 101]. Ôtani [774].

A.D [572, 681, 1003]. A.D. [866, 522, 799]. Åke [1235]. Åkerman [1235]. århundre [1029]. Årsbok [764]. Abacus [178]. ’Abbâs [1226]. Abbott [628]. Abel [434]. Aberration [174]. Abilities [1027]. Aborigines [575, 1074]. Abraham [1041]. Abram [779, 1221]. Abridged [388, 585]. Abteilung [157]. Abu [683, 1093, 230, 304, 359]. Abu’l [388, 579]. Abu’l-Farag [388]. Abu’l-Rayhan [579]. Abyssinia [1181]. académiques [237]. Academy [603]. Acanthaceae [106]. Acarologie [131]. accademia [603]. according [302, 1023, 254]. Account [613, 117]. Accounts [439, 568, 294]. achèvement [796]. Achievements [847, 853]. Acidum [836]. Acta [580, 580]. Action [1056]. Activities [803]. activity [201, 960]. ad [609, 1072]. Adam [819]. Adams [714]. Additional [544]. Additions [262]. Adelardus [712]. Adipocere [379]. Administrative [1204]. Administrator [61]. Admiral [521]. Adolf [528]. Adolphus [934]. Adoption [971]. Adrian [1193]. Adrien [182]. Advances [240, 1118]. Adventure [159]. Adventures [1112]. aedificatoria [1180, 755]. Aegypter [147, 146]. Aegypti [661]. Aerzte [1203]. af [757]. Affairs [998]. Afield [1154]. Africa [1004, 962]. Africae [661]. after [1015]. Aftermath [1073]. Afterwards [14]. Again [378, 651, 702]. against [110]. Age [1035, 984, 433, 353, 677, 769]. Ages [688, 648, 595, 1189, 556]. Agnes [1147]. Agricola [741]. Agricultural [617, 1067]. Agriculture [560, 561, 1065, 961, 953]. ägyptische [348]. ägyptischen [75]. ahbab [496]. Ahmad [45, 388, 579]. ahrām [1128]. Ain [45]. Ainsworth [189]. Air [1033]. Aires [1202]. Al [359, 525, 358, 474, 230, 304, 360, 387, 388, 579, 815, 1226, 1128, 1002, 837, 119, 474, 539, 602, 671, 835, 866, 641, 1128, 10, 360, 496, 815, 360, 608, 906]. al-ahbab [496]. al-ahrām [1128]. al-ashraf [815]. al-Awwal [1226]. al-Baladhuri [815]. al-Bannā’ [866]. al-Biruni [387, 579]. al-dakhîra [10]. Al-Farabi [359, 304, 608]. Al-fihrist [360]. al-Ghafiqi [388]. al-Ghazzālī [671, 835]. Al-Hakîm [837]. al-Hukamā’ [474, 539, 602]. Al-I‘tibâr [119]. Al-Khowarizmi [358]. al-kirām [1128]. al-Mā’ [641]. Al-morchid [525]. al-Nadim [360, 360]. al-Qiftī [474, 539, 602]. al-Qifti [474]. Al-Razi [906]. al-Tamīmī’s [641]. al-‘ulum [304]. al-Waraqī [641]. ’Alî [1226, 143]. Alan [718]. Alaska [1184, 1007]. Alaune [682]. Albert [1077, 1143, 321, 494, 557]. Alberti [755, 1180]. Alchemical [937, 1153, 1085, 750, 1089, 45]. Alchemie [416, 369, 187]. Alchemy [236, 700, 977, 277, 416, 847, 627]. Alchimy [417, 278]. Aldo [657, 1160]. Aldridge [816]. Aldrovandi [466]. Alen [837]. Alexander [1063]. Alexandra [79]. Alexandre [655, 64]. Alexandrien [1093]. Alfarabi [386]. Alfred [856, 218, 403, 613, 585, 1102, 1067]. Algebra [334, 683]. Algebraical [471]. Algorism [541]. Ali [230, 1230, 360]. Allee [1119]. Allen [408]. Allerton [988]. Allier [1166]. Allison [12, 244]. Allmittelpunkt [1142]. allocation [179]. Almquist [272]. Alois [80]. Alone [612]. Alphabet [42]. Alphonse [607]. Alpin [958]. Alte [749]. alten [951, 147, 146]. Alter [520, 353]. Altertum [786]. Altertumskunde [166]. ältesten [23]. Amateurs [1213]. Amazon [478]. Ambard [109]. America [531, 1077, 1191, 1053, 81, 1185, 613]. American [997, 481, 504, 630, 961, 155, 323, 407, 662, 16, 130, 1195, 575, 70, 808, 1129, 96, 783, 239, 491, 1078]. Americana [167]. Amethe [1058]. ami [553]. Amin [304]. among [1005, 1008, 1190, 323, 1028, 1074]. Ampère [258]. ’Amra [337, 337]. Amstelodamensium [1223]. Amsterdam [759]. Anagrammatism [751]. analyses [622]. Analysis [177, 722, 520, 762, 401, 422, 1089]. Analytic [838]. Ananda [591]. anatomicae [1223, 980]. Anatomist [120]. Anatomy [941]. anciens [921, 39]. Ancient [1076, 847, 993, 511, 1179, 443, 554, 936, 147, 544, 748, 277, 146, 189]. Ancients [217, 760]. Andalusi [230]. Anders [764]. Anderson [1048]. Andersson [555]. Andes [189]. André [258, 1037]. Andreae [980]. Andreas [980]. Andrew [663, 584]. Andrews [919]. Anfänge [484]. Anfängen [1165]. Angel [386, 107]. Angelo [511]. Anglo [541]. Anglo-Norman [541]. Animal [1119]. Animales [35]. Animalium [762]. Animals [886, 431, 960, 679]. Anna [511, 266]. Annals [720]. Annam [494]. annamite [494]. Anne [1029]. années [340]. Anniversary [76]. Annual [116, 251, 293, 567]. Ansab [815]. Anschütz [1104]. Antarctic [177]. antécédents [781]. Anthropological [887]. Anthropology [856, 788]. antico. [587]. Antike [841, 162, 1188]. antiken [615]. Antiquarische [507]. antique [826]. Antiquité [657]. Antiquity [512, 909, 814, 693]. Antoine [765]. Anton [507]. Antonini [825]. Antonio [743]. Antony [431]. août [64]. Apologia [389]. Appel [796]. Application [672]. Applied [725, 163, 263]. April [435, 532, 793, 1122]. Apuleius [854]. Aquapendente [479]. Aqueducts [589]. Arab [1141, 11, 866]. Arab-jin [1141]. Arab-Syrian [11]. arabe [866, 359, 1160]. Araber [262]. arabes [658]. Arabia [1049, 543]. Arabian [127, 15, 178, 967, 522]. Arabic [338, 45, 608, 358, 5, 474, 539, 602, 36, 143, 978, 253, 750, 171, 807]. Arabic-Latin [608]. arabicae [1072]. Arabische [474, 539, 602]. Arabs [995, 1028]. Aramaic [1090]. Arber [1147]. Archaeology [507]. archéologiques [340]. Archeology [1007]. Archibald [1043, 52].

Archimedes [221]. Architects [1161]. archives [108]. Ard [641]. Arduinus [1072]. Arepo [279, 279]. Aristotelian [269, 1132]. aristotélicienne [642]. Aristotle [34, 49]. Aristotle. [546]. Aristotlean [642]. Arithmetic [852, 1199]. Arithmetical [470]. Arithmetick [1212]. Arithmetik [75]. Arlindo [785]. Armand [921]. Armando [1181]. Armistead [245]. Armitage [623]. Army [812]. Arnold [52]. Arnost [780]. Arrhidaei [738]. Art [1024, 1145, 122, 579, 6, 591]. Artemision [973]. Arthur [110, 446, 859, 993, 663, 489, 767, 762, 401, 1009, 337, 433, 693, 627, 975]. Article [8, 43, 71, 931]. Articles [435]. Articuli [178]. Artifical [415]. Artificial [543, 1212]. Artificialis [518]. Arts [473]. Arturo [924, 190]. ärztliche [219]. as-sana‘ah [45]. as-san‘ah [45]. Ashley [1074]. Ashley-Montagu [1074]. ashraf [815]. Ashur [415]. Ashur-Nasir-Pal [415]. Asia [613, 636]. Asiatic [1237]. Asín [214]. Aslam [683]. Aspects [475, 1054]. Assaig [270]. Assyrian [822]. Assyrie [1227]. Astrolab [172]. Astrolabe [47]. Astrolabes [372, 402, 402]. astrologiae [1218]. astrologischen [581]. astrologorum [825]. Astrology [297, 1218, 579]. Astronomen [262]. Astronomers [1098]. astronomia [757]. Astronomical [548, 337, 62, 252, 803, 1214, 1001]. astronomie [213, 780]. astronomische [1214, 83]. astronomischen [111]. Astronomy [68, 975, 964, 1023, 1233, 309, 1000, 186]. Asyut [199]. at-Tabarî’s [143]. Athar [338]. Athar-i-Baqiya [338]. Athos [473]. Atlantica [1140]. Atlas [506]. Atoms [398]. Attitude [656]. Audubon [1185]. Auffassung [721]. August [991, 40, 325, 1104]. Augustus [244]. aun [45]. Aureo [1072]. Ausbreitung [187]. auspici [915]. auspiciis [609]. Australia [947]. Australian [1074]. Austryn [527, 49]. aut [585]. Author [28, 87, 114, 135, 195]. Authority [963]. Authorship [278]. Autobiographical [614]. Autobiography [1038]. autographe [224]. av [764]. Avant [232]. Avicenna [416]. Awwal [1226]. Axel [1140]. Axis [225]. Ayres [367]. Azaïs [340]. Azande [1008].

B [157, 998, 1006, 1007, 1101, 719, 121, 155, 827, 944, 14, 45, 952, 518, 1033]. B. [982]. b.’Ali [1093]. B.C [415]. B.C. [1169]. b.’Iraq [1093]. bêtes [1175]. Babel [717]. Babylonia [544, 748]. Babylonian [746, 707, 747]. Babylonie [1227]. Bachelard [317]. Back [25, 30, 58, 89, 275, 282, 289, 326, 332, 344, 375, 411, 437, 454, 460, 513, 534, 569, 597, 637, 667, 697, 735, 794, 831, 863, 895, 928, 965, 1012, 1082, 1124, 1150, 1170, 1207, 1241]. Background [447, 700, 760, 662]. Bacon [362, 254, 207, 349, 867]. Bagehot [805]. Baghdad [681, 1028]. Baker [403]. Baladhuri [815]. balistique [129]. Ballad [257]. Bammel [814]. Band [615, 106, 124]. Banks [947]. Bannā’ [866]. Baqiya [338]. Barber [754]. Barbour [552]. Barcelona [799]. Barlow [634, 485]. Barnett [1221]. Barnette [235]. Barock [404]. barométrique [781]. Baron [359, 1051, 1105]. Baroness [1105]. Barrett [1053]. Barry [55]. Bartels [692]. Bartholomew [816]. Barton [240]. Barton-Wright [240]. Bartram [481]. Basil [240]. Basile [390]. Basis [1193, 474, 539, 602]. Bate [846]. Bateson [889]. Bath [712]. Battista [1180, 755]. Battle [760]. Baumgartner [654]. Baxa [1068]. Baxter [616]. Bayet [321]. Beagle [485]. Beast [480, 1183]. beate [580]. Beatty [718]. Beaver [503]. Bede [898]. Bedford [148]. Bedouins [80]. Beebe [1238]. Beebe-Center [1238]. Beer [337]. before [531, 708, 486]. Beginn [23, 53]. Beginning [677, 1121]. Beginnings [726, 420]. Begriffsbildung [1220]. Behaviour [886, 963]. Being [1074, 859, 1009]. Beitrag [369]. Beiträge [759]. Bekämpfer [836]. Bell [999]. Bellingshausen [631]. Belon [466]. Beltrán [1202]. ben [683, 1136]. Benedetto [390]. Benevenutus [46]. Bengali [912]. Benjamin [498, 909, 159, 1114, 1115, 1224]. Bentley [401]. Berengarius [418]. Bergsoe [1190, 1005]. Bergson [320]. Bergsträsser [776, 675]. Beri [1107]. Beri-Beri [1107]. Bering [613]. Bernard [1103, 875, 758, 124]. Bernardin [550]. Berry [975]. Berthold [1195]. Bertholf [613]. Bertram [478]. Bertrand [1163, 190]. Berzelius [614]. B’estschen [775]. Betrachtung [426]. Bettina [578]. between [422, 1028, 263]. Beyond [235]. Beziehungen [721]. Bibhutibhusan [495]. Bible [849]. Bibliographical [1087, 374]. Bibliographikon [1144]. Bibliography [306, 467, 1237, 325, 16, 86, 27, 134, 194, 113, 247, 281, 286, 331, 342, 512, 409, 453, 532, 595, 636, 654, 664, 696, 731, 793, 893, 829, 861, 964, 1122, 1010, 1079, 1149, 1169, 1240, 28, 87, 114, 135, 195]. Bibliotek [764]. Bilikiewicz [404]. Bingham [167]. Bio [1144]. Bio- [1144]. biochemical [426]. biochemischen [426]. Biographical [20]. Biographik [219]. biographique [769]. Biography [362, 611, 987]. Biol [505]. Biologen [272]. Biological [112, 1010]. biologie [109]. biologischen [841, 528]. Biologist [784, 216]. Biology [942, 650, 1164]. Biometrician [767]. Bird [205]. Birds [1044, 1185]. Birgitta [580]. Birgittae [580]. Birgitte [580]. birgittina [580]. Birgittinska [580]. Biringuccio [741]. Birkenmajer [151]. Birth [1021, 972]. Birthday [435]. Birthplace [900]. Biruni [387, 579, 387]. bis [1165, 22, 23]. Biscoe [177]. Bishop [61, 871, 1033, 984]. Blachère [658]. Black [266, 848, 435, 444, 703]. Blackett [634]. blade [1015]. Blaise [1033]. Blake [606]. Blancanus [971]. Blanton [121]. Blasius [798]. Blazers [1113]. Blersch [1188]. Bligh [631]. Bliss [17]. Bloesch [941]. Blokh [20]. Board [933]. Boas [1197, 480, 693]. Boaz [1106]. Bodenheimer [84, 22]. Boerhaave [123]. Bogayevski [827]. Böhmischen [124]. Böhner [629]. Boise [589, 452]. bokora [772]. Bolton [1032]. Bolzanos [124]. Bombelli [73]. Bonaparte [820]. Bonfils [672]. Bongiorno [663]. Book [338, 588, 271, 75, 157, 314, 315, 316, 495, 615, 992, 692, 695, 754, 756, 768, 778, 783, 787, 788, 789, 790, 792, 842, 844, 849, 856, 860, 886, 887, 888, 889, 890, 891, 923, 924, 940, 1100, 1103, 1120, 941, 954, 956, 980, 983, 985, 986, 988, 989, 990, 991, 997, 998, 1004, 1005, 1006, 1007, 1008, 1032, 1034, 1038, 1043, 1047, 1073, 1076, 1161, 1167, 1168, 1190, 1193, 1197, 1198]. Book [1200, 1201, 1231, 477, 267, 373, 392, 1233, 531, 913, 310, 1077, 498, 1182, 1224, 120, 361, 685, 686, 110, 128, 160, 184, 215, 236, 306, 308, 311, 312, 371, 399, 400, 446, 447, 448, 486, 529, 558, 614, 716, 722, 723, 725, 782, 784, 847, 859, 905, 912, 914, 920, 1064, 1099, 1101, 1114, 1115, 1163, 1164, 524, 909, 951, 993, 1178, 1188, 661, 826, 958, 959, 1071, 1137, 1189, 1220, 319, 690, 777, 779, 858, 105, 109, 129, 231, 150, 265, 309, 548, 683, 689, 717, 719, 1050, 1051, 1052, 1059, 1093, 1106, 1227, 1228, 1057]. Book [1222, 492, 994, 320, 1203, 1029, 397, 527, 757, 763, 786, 881, 883, 885, 922, 1148, 182, 915, 1000, 1232, 1054, 369, 663, 617, 370, 1225, 562, 713, 878, 1030, 1031, 1158, 362, 367, 769, 854, 506, 721, 852, 121, 133, 156, 159, 188, 191, 192, 193, 218, 233, 239, 240, 241, 245, 246, 268, 269, 272, 273, 364, 365, 366, 374, 403, 404, 405, 431, 436, 480, 481, 483, 484, 485, 489, 491, 501, 503, 504, 505, 511, 546, 547, 549, 550, 551, 552, 553, 560, 561, 563, 566, 593, 594, 630, 631, 632, 633, 766]. Book [767, 841, 845, 846, 850, 851, 853, 855, 857, 947, 948, 950, 960, 961, 1039, 1040, 1042, 1066, 1075, 1105, 1113, 1116, 1118, 1119, 1159, 1165, 1179, 1183, 1184, 1185, 1186, 1187, 1191, 1192, 1194, 1196, 1221, 1234, 1235, 1236, 1237, 1238, 389, 429, 609, 132, 452, 510, 582, 583, 727, 761, 1140, 395, 152, 155, 165, 266, 322, 323, 324, 407, 408, 508, 662, 715, 762, 925, 398, 401, 499, 659, 919, 1061, 1143, 625, 1058, 451, 776, 127, 230, 235, 386, 390, 445, 995, 1230, 559, 565, 592, 244, 1001, 730, 564, 624, 628]. Book [634, 635, 656, 691, 760, 791, 963, 1035, 1056, 1199, 104, 232, 238, 242, 307, 317, 216, 36, 108, 843, 775, 848, 1162, 942, 1041, 1145, 610, 785, 1009, 337, 827, 944, 953, 1180, 981, 982, 1111, 1142, 103, 123, 153, 305, 321, 391, 526, 611, 714, 819, 879, 880, 943, 1098, 18, 20, 50, 73, 83, 161, 164, 213, 372, 402, 478, 724, 755, 814, 818, 957, 1060, 1117, 1138, 1139, 1181, 1229, 612, 984, 313, 500, 530, 613, 908, 151, 655, 106, 15, 17, 51, 74, 101, 102, 112, 122, 237, 589, 720, 877, 910, 911, 1096, 1107, 1108, 618, 619]. Book [620, 621, 688, 770, 771, 772, 773, 823, 824, 907, 916, 917, 918, 996, 1109, 1110, 1141, 318, 10, 81, 84, 85, 11, 12, 13, 14, 16, 19, 21, 22, 23, 24, 45, 46, 47, 48, 49, 52, 53, 54, 55, 56, 57, 76, 78, 79, 80, 107, 111, 125, 126, 131, 154, 158, 162, 163, 167, 168, 183, 185, 186, 187, 189, 190, 212, 214, 217, 219, 304, 339, 340, 356, 357, 359, 360, 368, 387, 388, 393, 394, 396, 406, 430, 432, 433, 434, 435, 443, 444, 449, 450, 482, 487]. Book [488, 490, 494, 496, 507, 509, 525, 554, 555, 556, 557, 578, 579, 580, 581, 584, 585, 586, 587, 590, 591, 606, 607, 608, 622, 623, 626, 629, 654, 657, 658, 681, 682, 684, 693, 711, 718, 726, 728, 729, 758, 759, 764, 774, 781, 815, 817, 820, 821, 822, 825, 828, 884, 892, 906, 946, 949, 955, 962, 987, 999, 1003, 1027, 1028, 1044, 1049, 1055, 1070, 1072, 1094, 1097, 1102, 1112, 1136, 1160, 1166, 1226, 1036, 1078, 1121, 1223, 1037, 1202, 1068, 1104, 1144, 502, 82, 124, 780, 952, 119, 358, 77, 130, 270]. Book [274, 687, 694, 712, 1048, 945, 660, 341, 493, 1033, 765, 1062, 1063, 166, 234, 616, 816, 882, 1045, 1046, 627, 363, 479, 528, 1095, 1067, 497, 921, 1053, 1065, 1069, 1074, 1146, 1147, 1195, 1002, 243, 37, 681, 20]. Books [760]. Bortolotti [18]. Botanic [1069]. Botanik [1165]. botanischen [1220]. Botanist [767, 255]. Botany [1236, 1147, 1237]. Bough [1073]. Boveri [490]. Bower [1236]. Boyer [918]. Boyle [306, 145, 256]. Bradbrook [1031]. Bradshaw [244]. Bradwardina [739]. Brahmagupta [548]. Bread [456]. Breasted [125, 443]. Breeding [1235]. Bretschneider [219]. Breves [418]. Bridgman [1167]. Brief [29, 88, 140, 196, 248, 288, 343, 459, 533, 596, 666, 734, 830, 927, 1011, 1123, 1206, 133, 1110, 457]. Briefwechsel [507]. Briet [410]. Brill [661]. Brim [849]. Britain [1236]. British [718, 616]. Britton [1003]. Bronislaw [24]. Brooks [1070]. Brown [720, 783, 526]. Browne [379, 380]. Brownell [624]. Bruce [1043]. Bruges [441]. Bruhl [857, 242]. Brunet [657, 104, 232]. Brunot [231]. Brunschvicg [310]. bruxellensis [980, 980]. Bryllion [481]. bu [619]. Bubnov [302]. Buch [265, 682, 906]. Buck [1117]. Buddhism [910]. Buddhist [911]. Buenos [1202]. Buffon [1175, 549]. Buffum [52]. Building [589]. Bull [52]. bunseki [772]. Burke [311]. Burlingame [1234]. Burr [767]. Burski [953]. Butlan [1028]. Butler [311, 365, 984, 778]. byô [1107]. Byzantium [595].

C [887, 940, 215, 1137, 1000, 1031, 240, 501, 631, 767, 1118, 1119, 132, 564, 691, 478, 1129, 85, 131, 219, 509, 1003, 358, 130, 945, 660, 1045, 1069, 672]. côtes [1184]. ca [802]. Cabins [503]. Caire [108]. Cairo [1028]. Cajal [949]. Cajori [584, 226]. Calamities [619]. Calculus [672]. Calendar [284, 1060]. calendario [1117]. Calendars [69, 199]. California [347, 508]. Californie [1184]. Cambridge [887, 954]. Camilo [785]. Campagna [511]. Campbell [984, 822, 1121, 266]. Canistris [817]. Cano [949]. Canon [799]. Canones [463]. canonizacionis [580]. canonizationis [580]. Care [1077]. Career [420]. Carl [1224, 1042, 1196, 130, 99]. Carlton [890]. Carmichael [128]. Carnap [1058]. Carnegie [295, 414, 652, 752, 938, 1134, 208, 516, 368]. Carpi [418]. Carré [307]. Carta [1181]. Carter [245]. cartògrafs [164]. Cartographica [661]. Cartography [319, 808, 1191]. Cary [913]. Casey [273, 843, 46]. Cassius [184]. Castiglioni [190, 924]. Casting [456]. catalans [164]. Catálogo [386]. Catalogue [1046]. Catalogus [825]. Catalunya [270, 164]. Catalysis [1091]. Cate [1048]. Categories [565]. Catlin [469]. Catoptrique [1178]. catoptromancie [395]. Cattell [998, 491]. Cave [152]. Cecidologie [629]. Cecil [561, 54, 1036]. Ceiling [62]. Ceiling-Decoration [62]. Celebrate [169]. Celebration [246]. Célestes [655]. Celine [988]. Cell [97]. Celli [511, 511]. Celli-Fraentzel [511]. Celsius [764]. Celtic [768]. cent [1240, 540]. centenario [972]. Centennial [246, 972]. Center [171, 1238]. Central [335, 219]. Central-Magazin [219]. Centuries [866, 1169, 8, 43, 71, 409, 453, 1030, 590, 623]. Century [70, 193, 843, 526, 16, 729, 202, 421, 1213, 45, 744, 1019, 1214, 868, 256, 466, 541, 465, 3, 636, 745, 96, 673, 144, 255, 742, 171, 642, 1198, 1222, 121, 480, 1183, 583, 942, 15, 48, 1121]. cérémonial [921]. Ch [548, 819]. Chûryô [618]. chûsei [688]. Chace [52]. Chadwick [1076, 1076]. Chain [859, 1009]. Chalmers [758]. Chambard [340]. Chamber [676]. Chandra [912]. Changes [960]. chapitre [586]. Chapter [747, 466, 422, 1147]. Chapters [407]. Charbonnier [129]. Charles [849, 983, 312, 777, 506, 132, 919, 445, 564, 635, 313, 500, 530, 76, 502, 616, 1067, 820, 902, 935, 1039]. Charles-Roux [820]. Charlotte [987]. Chartres [1045]. Chatelier [914]. Chaucer [47]. Check [424]. Check-List [424]. Chemical [575, 486, 847, 704, 180, 473, 20]. Chemie [836, 690]. cheminement [238]. Chemisch [1144]. Chemisch-Pharmazeutisches [1144]. Chemische [749]. Chemist [912, 3]. Chemistry [703, 422, 822, 1010, 836, 725, 920, 1064, 944, 765, 1063]. Chester [239, 718]. Cheynier [1037]. chez [1166, 921]. Ch’i [277]. Chia [377]. chih [617, 1027]. Chikashige [847]. Child [627]. Childe [923, 554]. Children [555]. chimie [317, 109, 1099]. Chimin [883, 396]. Chimique [1101, 123]. China [617, 475, 731, 555, 758]. Chinard [1184]. Chine [397]. Chinese [236, 883, 688, 396, 700, 1024, 677, 1133, 277]. chinesische [53]. ching [618]. chinois [1133]. chinoise [213]. Chirurgical [246]. Cholera [950]. Chou [618, 618]. Choulant [219]. Christ [836, 1217]. Christendom [1032, 1095]. Christiaan [430, 1097]. Christian [836, 473]. Christianity [1052]. Christiansen [730]. Christina [1049]. Chronological [585]. Chronologie [83]. Chronology [823, 1003]. Chung [617]. Chung-kuo [617]. Chuo [1108]. Church [1217]. Ciencias [386]. cinq [505, 340]. Circa [572]. Circle [943]. Circulation [540]. Cities [1200]. Civil [630]. civilisés [1166]. Civilization [789, 325, 86, 27, 134, 194, 113, 247, 281, 286, 331, 342, 512, 409, 453, 532, 595, 636, 664, 696, 731, 793, 893, 829, 861, 964, 1122, 1010, 1079, 931, 1149, 1169, 1240, 529, 81]. Civilizations [189]. clan [916]. Clara [365]. Clare [52, 857]. Clarence [367]. Clark [1238, 1035, 101, 13]. Clark-Kennedy [13]. Classic [341]. Classical [695]. Classification [304]. Claude [1103]. Claudii [356]. Claudius [524, 357, 987]. Clay [436]. Clifford [431]. Cline [1004]. Clinical [572]. Clipper [130]. Clocks [706]. Club [958, 958]. Coast [1191]. Coates [1193]. Codex [356]. codicum [609, 825]. Coffin [199, 983]. Coffins [252]. Cohen [998, 559, 1106]. cohérent [317]. Cole [188, 1223]. Colin [496]. Collected [312, 779, 685, 777, 846, 592, 313, 500, 530]. Collection [996, 881]. Collections [948]. Colledge [257]. College [233, 729, 573, 729]. Colleges [96]. collegii [1072]. Collier [624]. Collijn [580]. Collins [1007]. Colonel [557]. Colonial [1077, 1174, 574]. Color [63, 518, 647, 223, 960]. Color-Making [518, 647]. Coloribus [647]. Columbian [1005, 1190]. Combustion [138, 422, 94, 137]. Comenius [1172, 573]. Comet [467]. Cometis [296]. Comets. [1000]. Coming [1074]. Commemoration [267]. Comment [557]. Commentary [150, 463]. commentators [474, 539, 602]. Communion [897]. Company [1036]. Comparative [443]. Compared [145]. Compendium [773, 692]. Compendiums [143]. complètes [1097, 430]. Complex [1089]. Composaceae [106]. Composite [889]. Comptes [622]. Comte [821, 821]. Comte. [821]. Comtino [202]. con [209]. Concept [354, 805]. Conception [145, 804]. Concepts [648]. concernant [1017]. concerning [936, 618, 1017, 745]. Conductivity [354, 425]. conferenze [915]. Confucian [996, 677]. Congrès [955]. Conical [206]. connaissance [310, 397]. Conquest [1187, 50]. Conrad [490]. Conscience [443]. conservés [1017]. Consisting [435]. Constance [503]. Constantin [1105]. Constantinople [202]. constellations [147, 146]. Constructed [1020]. Contagione [152]. Contagiosis [152]. Contagium [604]. Contemporary [138, 94, 137, 1098]. Contenau [1227, 394]. Contents [29, 88, 140, 196, 248, 288, 343, 410, 459, 533, 596, 666, 734, 830, 927, 1011, 1123, 1206, 457]. Continent [177]. Continuo [739]. Contractus [172]. Contribution [138, 94, 443, 758, 1028, 826]. Contributions [695, 1235]. Controversy [1028]. conventu [1072]. conversion [1166]. Convictions [1186]. Convolvulaceae [106]. Conway [432, 726, 153]. Cook [631]. Coomaraswamy [591]. Cope [364, 364]. Copernic [655]. Copernican [801]. Copernicana [151]. Copper [1190]. Corlett [662]. corn [1015]. Cornelis [391, 879, 781, 361]. Cornelius [1114, 1115, 837]. Corpus [66, 269]. Corrections [262]. Correspondance [819, 622, 1055, 391, 879]. Correspondence [545, 229, 385, 876, 280, 1157, 100, 476, 710, 303, 813, 878, 840, 753, 442, 1177, 149, 1135, 264, 653, 118, 1219, 9, 355, 428, 523, 577, 605, 680, 979, 1026, 1092, 211, 336, 44, 72, 181, 285, 939, 392]. Corresponding [251, 293]. Cortesao [1181]. Cosmogonies [624]. Cosmogony [700]. Coulton [1201]. Council [33, 440]. Count [872, 873, 715]. Course [522]. Court [1019, 436]. Courtney [686]. Cox [398]. Craigie [949]. Cranston [193]. Creation [678]. Creative [501]. Cresca [49]. Cressy [1225]. Crew [659]. cristianizado [214]. Critical [177, 325, 86, 27, 134, 194, 113, 247, 281, 286, 331, 342, 512, 409, 453, 532, 595, 636, 664, 696, 731, 793, 893, 829, 861, 964, 1122, 1010, 1079, 1149, 1169, 1240]. Criticism [1218]. Critique [49]. Croce [390]. Crommelin [1000]. Cros [557]. crosses [37]. Crouse [632]. Crowther [997, 997, 851]. Cruces [37]. Cruise [368]. Crusades [11]. cuadrante [209]. Cubit [705]. cueillette [921]. Culpa [988]. Culross [853]. cultura [1117, 1232]. Cultural [890, 662]. Culture [889, 1200, 544, 748, 1196, 76]. Cumming [580]. Curatione [152]. Curie [987]. Curious [705]. curis [46]. cursor [209]. Curtis [564]. Curves [350]. Cusa [389, 429]. Cushing [638]. Customs [80]. Cutler [445, 130]. Cuvier [484, 643, 551]. Cyril [403]. Cytology [1118].

D [1103, 985, 1182, 128, 1000, 269, 1118, 1237, 230, 634, 720, 815]. D. [1043]. D’Acres [122]. da’Fontana [93]. dai [823]. dai-nenpyô [823]. Daisetz [910, 911]. dakhîra [10]. d’Alexandre [394]. Dalmatia [1017]. Dalmatie [1017]. Dalmatius [799]. Damme [818]. Dampier [54]. Dampier-Whetham [54]. Dance [218]. danicus [759]. Daniel [818, 586, 882]. Danish [556]. Dannemann [623]. dans [914, 242, 1160, 687]. Dantzig [1054, 185]. D’après [946]. d’archéologie [394]. D’Arcy [191, 1044]. Darlington [1118]. Darwin [485, 805, 804, 643, 678, 21]. Data [1229]. Date [543, 415, 544, 748, 798, 807]. Date-Palms [543, 415]. Dates [1229, 866]. d’atome [716]. Datta [719, 495, 45]. d’Auguste [821]. David [531, 943, 56, 79, 520, 943, 79]. David-Neel [79, 79]. Davis [271, 449]. Dawn [443]. Dawson [1159]. Dawtrey [233]. Day [297]. Dean [877]. Death [266, 498, 218]. débat [381]. Décédé [64]. decem [1180, 755]. December [281, 342, 696, 861]. Decimal [708, 542, 699, 672, 797, 1229]. Decimalization [175]. Decimals [1212]. Decline [1109]. Decoration [62]. découvertes [505]. Defence [977]. Definition [871]. d’Égypte [820]. dehors [946]. deinde [1015]. Dekane [885]. Dekansternbilder [885]. del [164, 587, 1202, 209]. Delatte [395, 921]. Delépine [716]. Delivered [486]. Della [756, 924, 603, 915, 972, 18]. Deman [589]. Democritos. [814]. Denken [1188, 363]. Denker [1111]. Deny [108]. depuis [394]. d’Erasme [818]. deren [474, 539, 602]. dérivés [395]. d’Erlanger [359]. Dermatology [408]. dernier [224]. Descartes [207, 1018]. Descartes. [819]. descobrimento [1181]. Description [1137, 718, 549]. Descriptions [341]. Desert [1049]. Dessoir [243]. Destrez [687]. Determination [425]. Determine [613]. Determining [963, 519]. Deutsch [1077]. deutsche [1139]. deutschen [111]. deux [320]. Deuxième [1084]. Deva [298]. Development [725, 403, 1235, 925, 729, 522, 789, 546]. Developments [258]. Devil [274]. d’histoire [955]. d’Història [270]. Diaet [730]. Diagnosis [828]. Diamond [790]. Diary [986, 714, 393, 487, 485, 766]. d’Ibn [1136]. Dickinson [489, 563]. Dictionary [822]. Dictionnaire [769]. Did [838]. Diderot [1100]. Diepgen [951]. Diès [622, 105, 1055]. Dietetics [377]. Dietrich [1142]. Differences [519]. Differential [299, 351]. differentiation [468, 468]. Digby [1034]. Dingle [1057]. Dingler [499]. Dingwall [692]. Dioscorides [547]. Dirac [723]. direct [468]. direkten [468]. d’Irsay [645]. Disciple [380]. discorsi [915]. discours [237, 884]. Discoverer [973]. Discoverers [631]. Discovery [174, 225, 177, 38, 520, 838, 206, 63, 644, 427, 478, 1181, 806, 97, 176, 128]. Diseases [341]. Disme [542]. Disney [403]. Dispersion [63]. Disputed [873]. Distemper [828]. Distillation [422]. Dittrich [780]. Diurnal [809]. Divinity [1221]. Division [652, 752, 938, 1134, 708, 516]. do [1181]. Dobell [431]. Dobiash [953]. Dobiash-Rozhdestvenskaya [953]. Dobzhansky [1194, 1146]. Docta [429]. doctae [389]. docteur [406]. Doctors [254]. Doctrine [123]. d’oculistique [525]. Documentary [693]. Documents [397, 946]. Dodds [1094]. Dohrn [363]. Dom [492]. Dominica [609]. Dominicans [868]. Dominis [743]. Donald [853]. Donne [983]. d’or [957]. Doren [1224]. D’Ortous [392]. Dorveaux [1130]. Doubrovnik [1017]. Douglas [765]. Dove [230]. Dr [419, 509]. Dr. [638, 380, 1155]. Drachman [850]. Drama [633]. Drawing [122]. Drawn [889]. Drebbel [361]. Drewitt [233]. Dreyer [423]. Drinker [364]. Drogenkunde [387]. Drugs [388]. Dry [422]. Dry-Distillation [422]. Dualism [110]. Dualistic [700]. Duane [919]. Dubrovnik [1017]. Ducklings [1127]. Duguet [406]. Duhem [701, 796]. Dumas [1101]. Duncan [848, 435, 444]. d’une [79]. durch [484, 841]. during [620, 684, 673, 574]. Dutcher [244]. Dwelling [507]. Dynamic [160]. Dynamidia [649]. Dynasty [62, 252].

Eagle [630]. ear [1015]. Earliest [970, 335, 1121, 807]. Early [495, 768, 686, 258, 562, 648, 155, 204, 801, 329, 382, 36, 180, 103, 1129, 490, 144, 228, 1215, 714, 954, 188, 103, 1045]. Earnest [919]. Earth [555, 225, 809]. Earthquakes [619]. East [554, 978]. Eastern [1237, 868].

Ebbell [952]. Ebers [952]. Echeneis [7]. Eckardi [609]. Eckstein [268]. éclaircissements [661]. Eclipses [724]. Ecological [1119]. Economo [1105]. edat [164]. Eddington [446]. Edessa [681, 681]. Edgar [186]. edita [609]. Edith [1186]. Edmond [392]. Edmund [23, 187]. Edo [620, 620]. Education [370, 694]. Edward [789, 524, 1225, 713, 1030, 233, 364, 974, 357, 758, 1095, 985]. Edwin [1113, 125]. Efforts [175, 613]. Egbert [1237]. Egg [176]. egritudinibus [46]. Egypt [443, 512, 566]. Egyptian [14, 977, 706, 470, 705, 1128]. Egyptians [147, 146]. Eigenschaft [179]. Eight [967]. Eighteenth [1183, 3, 409, 96]. Eighth [251, 43, 86, 409, 829, 87]. Eilhard [40]. Einleitung [1139]. Einstein [1143, 557, 1041]. Ekelöf [585]. Elastic [350]. Elected [60]. Electric [97]. Electrical [263]. Electromagnetic [258]. Element [145]. Élémentaire [1099]. Elementar [1059]. Elementar-Mathematik [1059]. Elementary [422]. Elements [377, 703, 313, 579, 1094, 150]. Elephantine [1090]. Elephants [771]. Eleventh [567, 45, 171]. Élie [129]. Eliot [729]. Elizabethans [713]. Ella [374]. Ellery [1154]. Elliot [1159]. Elmer [1237, 877]. Éloges [237]. Eloise [1158]. Emanuel [112, 585]. Emanuelis [585]. Embryologie [404]. Embryology [176, 855]. Emergence [311, 1196]. Emil [423]. Emile [884, 238, 237]. Emmanuel [892]. Emotion [1238]. Empédocle [105]. Encyc [505]. Encyclopaedia [681]. encyclopedique [946]. End [595, 636, 664, 731, 793, 893, 829, 861, 964, 1010, 1079, 1149, 96, 1122]. Enderby [177]. Engineer [489]. Engineering [228, 1225]. England [253, 1213, 1001]. English [1047, 418, 204, 144]. Enquiry [278]. Enriques [587, 1232, 73]. Entitled [417, 1181, 119, 277]. Entomological [1066]. Entomologie [22, 84]. Entomology [241, 163]. Entstehen [780]. Entstehung [1188, 6, 187]. Entwicklung [841]. Entwicklungsgang [53]. Enumeration [304]. Eorum [152]. eorumque [46]. épars [661]. Ephéméride [818]. Ephorus [754]. Epidemiographik [219]. Epiphanius [877, 606]. Epitomes [742]. époque [394, 1226]. Equilibrium [1033]. Era [803]. Erasmus [897]. erecti [1072]. Eresos [841]. erfolgreiche [836]. Erforscher [154]. Erforschung [484]. Eric [692, 323, 443, 999, 1094]. Erkenntnis [528]. erklärt [265]. Erklärung [279]. Ermengaud [798]. Ernest [986, 1161, 769, 559, 984]. Ernestus [429]. Ernst [272, 161, 1138, 1139, 57, 111]. Errata [138]. Error [1192, 1112]. Errors [254]. Erste [106]. ersten [1165]. Erster [615]. Ervin [945]. Erxleben [836, 836]. España [316]. esprit [397, 946]. Essai [588]. Essais [129]. Essay [711]. Essays [76, 856]. Essig [241]. Esther [561, 589]. étape [716]. Ethiopie [340]. Ethnological [1120]. Etsu [916]. Etsujiro [771]. Ettore [18, 102]. Étude [781]. Études [492]. Euclid [150, 474, 539, 602, 1133]. Euclide [1178, 1133]. Euclides [759]. Eugene [1032, 392, 565, 1098, 101, 56, 1048, 531]. Eugenio [98]. Euklids [474, 539, 602]. Euler [350]. Europa [209]. Europe [335, 209, 890, 1076]. European [554, 700, 789]. Evan [905]. Evans [1008]. Evans-Pritchard [1008]. Eve [987]. Evelyn [17]. Events [1066]. Ever [812]. Evolution [1143, 228, 1147, 1160, 501]. Exact [500]. exactes [109]. Exemplar [1139]. Existed [812]. Existence [399]. Expanding [446]. Expansion [324]. Expedition [887, 50]. Experiences [912]. Experimental [425, 590]. expérimentales [914]. Experimentation [1067]. Explanation [743, 542, 1132, 279]. explications [781]. Exponential [672]. Extracts [943]. Extraits [658]. Eyre [789].

F [941, 1034, 1168, 1182, 215, 306, 400, 1228, 854, 188, 233, 403, 1236, 132, 165, 401, 755, 613, 318, 84, 22, 190, 219, 509, 587, 623, 654, 815, 825, 1223, 1062]. F. [1074]. fī [1128]. Fabric [426]. Fabricius [479]. Fabrikation [23]. Fabry [941]. Face [456]. Facsimile [1181, 542]. Factors [925, 963]. Facts [127]. Faculty [246]. Fagin [481]. Fahie [51]. Fairfield [392, 364, 1098, 21]. Faksimiledruck [1139]. Famous [997]. Farabi [304, 359, 608]. Faraday [393, 487, 393, 487]. Farag [388]. Farbentheorie [223]. Farmer [127, 15, 217, 608]. Farrand [1039]. Farrington [909]. Fascicule [231]. Father [765]. Fathers [694, 624]. Faust [800]. Favre [182]. Fay [244]. Features [1108, 574]. February [636, 731, 893]. Federigo [1232, 73]. Feeling [1238]. Feibleman [1231]. Feldhaus [162]. Feldman [309]. Felicis [609]. Fellow [306]. Ferchl [1144]. Ferdinand [514, 988, 231, 507]. Fertilization [543, 415]. Feudal [1109, 916]. Fever [245]. fidem [609]. Fieri [518]. Fifteenth [48, 590, 742]. Fifth [60, 325, 696, 1149]. Fiftieth [893]. Fifty [964, 1122, 1010, 1079, 1149, 1169, 1240, 504]. Fifty-Fifth [1149]. Fifty-First [964]. Fifty-Fourth [1122]. Fifty-Second [1010]. Fifty-Seventh [1240]. Fifty-Sixth [1169]. Fifty-Third [1079]. fighters [836]. Figures [7, 1102]. fihrist [360]. Fiji [631]. fi’l [525]. fi’l-kohhl [525]. Filosofiska [761]. fin [994]. Findlay [1063]. fine [1002]. First [177, 215, 868, 1020, 180, 8, 43, 71, 194, 542, 532, 964, 195, 358, 1087, 603, 673, 1015]. Fischer [126, 356]. física [1062]. Físiques [270]. fitologia [785]. Five [466]. Flamsteed [1098]. Fletcher [1221]. Flora [106]. Florence [852]. Florian [584, 226]. Floyd [626]. Fly [222]. Foerster [81]. Folk [451]. Folkways [888]. Fontana [201, 201, 378]. Fontenelle [307]. Food [1053]. Foraminifera [405]. Förderung [841]. Forel [991]. Forgotten [432, 255]. Form [426, 465]. formation [588, 397, 821]. Formel [279]. Former [435]. Formula [330, 279]. Forschung [426]. Forschungsmethode [841]. Fort [873]. Fortieth [512]. Forty [532, 595, 636, 664, 696, 731, 793, 829, 861]. Forty-Eighth [829]. Forty-Fifth [696]. Forty-First [532]. Forty-Fourth [664]. Forty-Ninth [861]. Forty-Second [595]. Forty-Seventh [793]. Forty-Sixth [731]. Forty-Third [636]. Foster [760]. Fotheringham [1022, 903]. Foundation [445, 176, 1116]. Founder [507, 821]. Founding [729, 694]. Fournier [505]. Fourteenth [590, 541, 636]. Fourth [286, 664, 1122]. Fr [219]. fra [1029]. Fracastorii [152]. Fracastoro [582, 654]. Fractions [542, 672]. Fraentzel [511]. Fragment [761]. Fragments [814, 1066]. Frame [993]. Francais [958, 505, 482, 958]. francaise [231, 586]. France [397, 769, 232, 526, 586, 1045]. Francis [630, 948, 322, 1001, 488, 821, 362]. Francisco [316]. Francois [820]. Frank [613]. Franklin [1224, 479]. Franz [1197, 717, 981, 53, 162]. fratrum [1072]. Frauenheilkunde [951]. Frazer [1073]. Frédéric [674]. Frederick [55, 934]. Freeman [565]. Freind [901]. Fremont [902]. French [480, 1183, 866, 1017, 96, 1133, 642]. Friar [254]. Friend [1231, 801]. frije [757]. Frisius [801, 519]. Fritz [690, 786, 610, 337, 1117, 1144]. Frobisher [1158, 1158]. Frolov [990]. Front [1, 26, 32, 59, 91, 115, 136, 142, 170, 198, 220, 250, 276, 283, 291, 327, 333, 346, 376, 413, 438, 455, 462, 515, 536, 570, 599, 639, 669, 698, 737, 795, 833, 864, 896, 930, 966, 1014, 1083, 1126, 1151, 1171, 1209]. Frontier [1040, 888]. Frontiers [1006]. Frontispiece [514, 638]. Frost [245, 950]. frumentum [1015]. Fuchs [231]. Fujinami [996]. Fulbert [61]. full [1015]. Fulton [306, 132, 1034, 165, 654]. Functions [449]. Fundamental [640]. Fundamentals [546]. Funktionenlehre [124]. Further [750]. Fürtrefflichkeit [941].

G [314, 315, 754, 888, 997, 1201, 361, 562, 878, 365, 631, 767, 841, 851, 960, 1042, 1235, 1238, 1035, 819, 943, 814, 613, 10, 219, 388, 394, 587, 728, 1055, 1033]. G. [972]. Gôichi [996]. Gabriel [1021]. Gager [1069]. gaikyô [1141]. gaku [1110]. Galilée [207]. Galilei [51]. Galileo [51]. Gallico [582]. Gallicus [654]. Galloway [405]. Galt [452]. Galton [488]. Gandz [358]. ganka [1110]. ganka-gaku [1110]. Gardens [621, 1069]. Gardiner [1238, 718]. Garret [881]. Garth [1019]. Gases [425]. Gaston [317]. Gaudin [716]. Gauthier [1136]. Gave [940, 1053]. geb [261, 40]. Gedanken [749]. Gegenwart [1148, 721, 1165]. Geheimnis [906]. Geheimnisse [906]. Geiser [1040]. Gemistus [899]. Gemma [801, 519]. Gemmis [606]. genannt [1138]. Genchi [772]. Geneeskunst [982]. General [512]. generale [663]. generales [463]. Generation [678, 188]. Generatione [762]. Genesis [576]. Genetics [1194, 1146]. Genius [444, 1033, 848]. Gentile [1088]. Gentleman [11]. geodätischen [786]. Geodesy [70]. Geoffroy [484]. géographes [658]. Geographiae [356]. Geographical [899]. Geographie [219]. Geography [562, 506, 563, 1119, 524, 1189, 357]. Geology [822]. geometer [1017]. Géomètre [1017]. Geometric [301]. geometrica [39]. Geometriche [314]. Geometrie [348, 384]. Geometry [838, 299, 351, 358, 495]. Georg [759, 844, 761]. George [1073, 184, 371, 1189, 779, 156, 480, 947, 510, 127, 244, 564, 635, 15, 497, 1056, 974, 469, 1154]. Georges [1227, 496]. Georgius [899]. German [173, 201, 676, 1214, 279, 179, 426, 6, 474, 539, 602, 353, 468, 147, 223, 416, 836, 739, 146, 741]. germanischen [166]. Germany [335, 778]. Gerold [507, 994]. Geschichte [157, 615, 179, 369, 484, 353, 161, 84, 22, 219, 690, 1059, 1203, 786, 1165, 499, 23, 629]. geschichtliche [426]. Gesellschaft [1203, 507, 124]. Gesetz [858]. Gesner [466]. gest [261, 40]. Gestalten [1111]. gestalter [580]. Ghafiqi [388]. Ghazzālī [671, 835]. Ghetaldi [1017]. Ghubār [178]. Giambattista [390]. Gibbon [985]. Gidon [190]. Giftbuch [578]. Gilbert [1184, 207]. Gilding [1190]. Gilfillan [691, 660]. Gino [915]. Ginsburg [531]. Ginzberg [1106]. Ginzburg [159]. Giovanni [201, 201, 93, 378, 1088]. Girolamo [582, 654]. Gittings [1186]. Given [1085]. Glaser [969]. Glass [677]. Glasser [490]. Gliozzi [1002]. Glossary [1044, 807]. God [14, 860]. Gods [772]. Goethe [383, 154]. Goethes [363]. Gohory [970]. Goitein [815]. Gold [1005, 427]. Golden [1073]. Golder [613]. Goncal [164]. Goncalves [600]. González [386, 107]. Good [865]. Goodyer [547]. Gordon [923, 1187, 554]. Gortner [767]. Gottfried [77]. Gotthelf [776, 675]. Gouhier [821, 821]. gouvernail [826]. gouverneur [820]. Government [583]. Grace [155]. graecorum [825]. Graecus [356]. Grafton [1159]. Graminaceae [106]. Grammar [670]. grand [381, 482, 1037]. grand-père [1037]. Granhall [1235]. Grant [1049]. Grassus [46]. Graves [240]. Graveyard [873]. Gravitation [782]. Gray [561, 986]. Great [859, 905, 466, 853, 1236, 1009, 875]. Greco [470]. grecque [395, 588, 434]. Greece [778, 532]. Greek [993, 547, 183, 1028, 1044, 627, 5, 471]. Greeks [21, 838]. Green [853]. Greene [612]. Gregoriaansche [957]. Grégorien [492, 492]. Gregorius [388]. Gregory [889, 943, 684]. Gresham [257]. Greta [585]. Greyhounds [130]. Griffiths [791]. Grosse [272]. Grosseteste [296, 464]. Grousset [78]. Growth [1076]. Grund [474, 539, 602]. Grundlagen [75]. grundsätzliche [841]. Grundzüge [776]. Guide [244, 525, 728]. Guinea [889]. Gumpert [1113]. Gundel [885]. Gunnar [555]. Gunter [547]. Gunther [103, 402, 954, 714, 372, 47]. Gustav [150, 268]. Gyldne [730]. Gynaecological [692]. Gynecology [695].

H [856, 890, 891, 1120, 941, 1076, 1198, 448, 1189, 1052, 721, 374, 436, 485, 489, 766, 950, 960, 1179, 1191, 1237, 1238, 583, 925, 634, 635, 1056, 50, 478, 612, 720, 76, 219, 496, 585, 718, 828, 1055, 341, 1033, 616]. H.S.S. [638]. Hôken [1109]. ha-Middot [358, 358]. Haberly [1047]. Habit [55]. Habitat [222]. Haddon [887]. Haeser [219]. Hahnemann [982]. Haitian [940]. Hakîm [837]. Hakluyt [878]. Hakluyts [878]. Haldane [784]. Haldór [319]. Hale [1154]. Hales [13]. Half [868, 199, 673]. Hälfte [106]. Hall [686]. Halley [392, 1098, 392]. Halliwell [259]. Hallucinations [959]. Halsted [156]. Hamilton [206]. Hampton [245]. Handschrift [1214]. Handschriften [111]. Hans [1148]. Happy [480]. Harcourt [526, 720]. Hardy [846, 846]. Harland [368]. harmonie [550]. Harold [271, 449, 816]. Harris [767, 787, 508]. Harry [527, 155, 49, 1216]. Hartley [816]. Hartshorne [312, 777, 592, 313, 500, 530]. Harvard [729, 552]. Harvey [638, 663, 207, 762]. Hasidismus [775]. Haskins [76, 935]. Hassfurt [702, 800]. Hassidism [689]. Hastings [880, 1183]. Hatfield [112]. Hayashi [918]. Hayes [890, 1187]. Hayy [1136]. Hayyān [92]. Hazim [230]. Healing [773]. Health [933, 192]. Heat [919, 465]. Heath [183]. Heaton [478]. Hebrew [848, 444, 358, 42]. Hecker [219]. Hedrick [560]. Heiberg [1055, 74]. Heidel [993, 433]. Heilmittel [676]. Heimann [817]. Heinemann [692]. Heinrich [692, 581, 261]. Helen [19]. Hélène [109, 123]. hellène [105]. Helmholtz [363]. Helvetius [227]. Henderson [631]. Hennig [1071]. Henning [234]. Henri [320, 606, 821]. Henry [1007, 380, 308, 914, 133, 239, 245, 364, 766, 1191, 452, 508, 659, 127, 244, 714, 1181, 15, 17, 21, 52, 125, 217, 443, 608, 758, 1160, 1078, 419, 260]. Henry-Paul-Joseph [1160]. Henschel [219]. Heny [252]. Herausgegeben [124]. Herb [621]. Herbal [854, 547]. Herbals. [1147]. herbam [1015]. Herbarius [921]. herbe [970]. Herbert [695, 1032, 371, 858, 1057, 1066, 407, 908, 77]. Hereford [253]. Hermann [692, 558, 126, 363]. Hermannsson [319]. Hermannus [172]. Hermes [41, 839]. Heroic [433]. Herrick [630]. Hesse [1119]. Hester [1183]. Heusinger [219]. Hevelius [1098]. Hewer [240]. Hieratic [718]. Hieronymi [152]. Hieronymus [479]. Higher [449]. Hilaire [484]. Hildegard [1078]. Hilden [941]. Hill [403]. Him [435]. Himself [923]. Hindu [495, 719, 1229, 699, 797]. Hindus [936]. Hiscock [943]. Histoire [1175, 994, 826, 105, 129, 968, 586, 1084, 231, 324, 318, 78, 190, 657, 892]. Historia [316, 1072, 1202, 1062]. Historian [754, 975, 901]. Historical [652, 752, 938, 1134, 692, 506, 244, 650, 1107, 621, 516, 426, 259, 127]. Historien [902]. Historiographical [1048]. historique [509, 781]. historisch [131]. historischer [53]. History [338, 138, 295, 414, 652, 752, 938, 1134, 890, 1120, 531, 354, 425, 1077, 184, 1064, 1050, 1051, 466, 617, 133, 241, 436, 511, 546, 560, 855, 961, 1066, 1238, 132, 408, 628, 1131, 180, 422, 94, 325, 984, 808, 15, 112, 720, 688, 772, 917, 1110, 16, 19, 54, 8, 43, 71, 86, 2, 27, 163, 134, 194, 113, 208, 247, 281, 286, 331, 342, 347, 490, 512, 409, 453, 590, 623, 516, 542, 532, 595, 636, 571, 693, 664, 696, 731, 793, 751, 745, 893, 829, 861, 834, 1028, 964, 1122, 1010, 1079, 931, 1149]. History [1169, 1240, 1036, 1121, 299, 351, 424, 574, 300, 352, 1067, 497, 1147, 201, 179, 603, 353, 60, 33, 116, 251, 293, 567, 732, 862, 1081, 76, 440, 891, 630, 1069, 719, 883, 370, 561, 322, 995, 827, 396, 742]. Hitherto [1218]. Hitti [995, 11, 119]. Hjelmslev [759, 759]. Hjort [1164]. Hobart [630]. Hobbs [908]. Hoenigswald [1111]. Hoffman [429]. Hogben [956, 1199]. Holcomb [940, 193]. Holding [7]. Holtsmark [1029]. Homans [564]. Homer [76, 935]. Honain [1096]. Honesty [1152]. Honigmann [57]. Honor [1048]. Honourable [306]. Honpô [772]. Hooke [103, 714, 95]. Hoops [166]. Hope [153]. Hopkins [627]. Hopper [1228]. Horatio [1129]. Horne [949]. Horologium [67]. Horton [816]. Horton-Smith [816]. Hosier [744]. Hospice [875]. Hospital [193]. Hourticq [168]. Houston [366]. Howard [163]. Howarth [563]. Hoyt [612]. Hrotsvitha [477]. Hsi [711]. Hsiang [1153]. Hsiang-Hsien [1153]. Hsien [1153]. HSS [1080]. Huai [700]. Huai-nan-tzu [700]. Huber [1042]. Hubert [581]. Hübotter [53]. Hudson [612]. Hugh [1049]. Hugo [499, 39]. Huizinga [828, 828]. Hukamā’ [474, 539, 602]. Hukamaā [474]. Hultsch [674]. Human [1164, 1192, 1196, 963, 650, 772, 1027, 998]. Humanism [670, 184, 81, 184].

Humanist [61]. humanitaire [483]. Humphrey [206]. Hundert [507]. Hundred [788, 1233, 1225, 1063]. Hungarian [868]. Hunger [854]. Hunt [244]. Hurd [1121]. Hurd-Mead [1121]. Hurry [14]. Hurst [501]. Huxley [766, 634, 766, 962, 367, 366]. Huygens [430, 1097, 1097]. Hyatt [1053]. Hybridization [432, 726]. Hydriotaphia [379]. Hyoscyamus [1156].

I-ka [996]. I. [494]. Ibn [474, 539, 602, 641, 1096, 359, 360, 1028, 119, 358, 92, 230, 10, 11, 388, 579, 540, 866]. ibn-Munqidh [11]. Ibni [1211]. Iceland [319]. Icelandic [234]. Ichthyology [466]. Icones [980]. Iconographia [580]. Iconographic [1152]. Iconography [473]. Idea [542]. Ideale [528]. Idealism [1088]. Ideas [936, 693, 678, 1161]. Idee [981]. Ideen [528]. Idees [270, 231, 586]. identification [147]. Identifikation [147]. Identity [417]. Ie [231]. igaku [688]. Ignorantia [429]. ignorantiae [389]. ihr [223, 53]. ihre [279, 841]. ihrem [75]. Ihsa’ [304]. II [103, 402, 313, 351, 539, 137, 352, 912, 1137]. III [500, 590, 602, 152, 486]. IIIe [955]. iki [772]. iki-bokora [772]. Ill [1077]. Illustrations [8, 43, 71]. illustrée [818]. im [1188, 786, 404, 219]. images [168]. Imhotep [14]. Immanuel [612, 106, 672]. Importance [471]. Important [1089]. Impregnated [679]. Inô [774]. Incident [873]. Incipits [1046]. including [1066]. inclusive [1122, 1240]. incognitae [1071]. Incunabula [167]. Index [28, 87, 114, 135, 195]. India [664, 750]. Indian [662, 68, 69, 469]. Indian-Loving [469]. Indianapolis [1080]. Indians [1005, 1190, 323]. Indicus [869]. Individual [1167, 963]. Industries [575]. inédit [95]. inedita [585]. inédite [39, 674]. inedits [946, 397]. Infeld [1143]. Inferior [648]. Influence [256, 96, 171, 127]. Influences [1176]. Information [31, 90, 139, 141, 197, 249, 287, 290, 345, 412, 458, 461, 535, 598, 665, 668, 736, 832, 926, 929, 1013, 1125, 1205, 1208]. iniziativa [915]. Inledning [761]. Inquiry [618]. Inquisition [1201]. Inscriptions [252]. Insects [222]. Institute [931]. Instituti [609]. Institution [295, 414, 652, 752, 938, 1134, 208, 516]. Instruction [970, 579]. Instrument [744, 205]. Instrument-Maker [205]. Instrumente [786]. Integral [1020]. Intellectual [522]. intellectuelles [586]. Intelligent [1167]. Interests [913]. International [1162, 955]. internationale [1084]. Interpretation [925, 845, 444]. Interpreter [1100, 481]. interregnum [836]. Interregnums [836]. introducción [209]. Introduction [886, 1115, 273, 452, 559, 564, 475, 910, 799, 1184, 819, 209, 626, 497]. Inventing [660]. Invention [672, 691]. Inventions [628]. Inventoribus [203]. Inversion [301]. Invisible [573]. Involved [925]. Ioannes [802]. Iran [829]. Iranian [1211]. Irish [616, 3]. Iron [1234]. Isaac [1038, 763, 880, 943, 584, 611]. Isagogae [418]. Isak [580]. Ishâq [1096]. Isis [439, 733, 568, 117, 294, 292, 328, 457, 537, 894, 1239]. Islam [214, 829, 445, 1230]. islamischen [776]. Island [1007]. Islands [631]. islandsk [1029]. Israel [793]. Italia [18]. Italian [603, 972]. italienischen [1111]. I‘tibâr [119]. Itineraria [585]. iudiciorum [1218]. IV [714, 530, 590].

J [787, 849, 890, 891, 1103, 989, 997, 998, 1006, 1182, 120, 306, 725, 1064, 1101, 661, 109, 492, 188, 405, 491, 563, 851, 1105, 1187, 1196, 583, 307, 51, 74, 112, 80, 190, 219, 368, 496, 828, 1055, 1223, 270, 694, 616, 479]. J. [259, 1022, 767]. J.C. [866]. Jābir [92]. Jâhiz [35]. Jên [688]. Jackson [184, 768, 418]. Jacob [614, 689]. Jacquemont [553]. Jacques [970, 918]. Jaeger [546]. Jager [1212]. Jahn [219]. Jahre [507]. Jahrhundert [1214]. Jahrhunderts [53]. Jaina [284]. Jak [82]. Jakob [1068]. James [888, 1073, 1231, 160, 447, 663, 1185, 715, 1061, 877, 125, 443, 694, 1048]. Jamieson [14]. Jan [828, 900]. Januar [40]. January [60, 86, 512, 1079]. Janus [219]. Japan [823, 1110, 731]. Japanese [1108, 621, 917, 1109, 774]. Jaques [491]. Jardin [1175]. Jastrow [1192]. Jauregg [1105]. je [557]. Jean [108, 946, 687]. Jeannette [593]. Jeannin [492]. Jeans [447, 1061]. Jeanselme [324]. Jefferson [175]. Jekuthiel [531]. Jen [1027, 1107]. Jenner [233]. Jenseits [243]. Jensen [1148]. Jerome [1070]. jeunesse [434, 821]. Jewish [1050, 445, 978]. Jews [1023, 1051, 1106]. Jidai [620, 1141]. jigyô [823]. jin [1141]. jiseki [1141]. Jitsizô [1141]. Job [265, 681]. Joh [854]. Joh. [836]. Joham [1181]. Johan [1164, 727, 761, 555]. Johann [267, 717, 764, 514, 702, 800]. Johanne [730]. Johannes [1059, 1138, 1139, 154, 759, 166, 836, 816]. John [692, 986, 1034, 1043, 1182, 311, 784, 920, 881, 506, 1086, 547, 1185, 1238, 132, 165, 871, 85, 76, 186, 654, 1027, 901, 493, 627, 138, 973, 4, 205, 94, 137, 423, 903, 651, 983]. Johnson [236, 913, 1001, 616]. Johnsson [4]. Jonathan [1100]. Jones [1038, 760]. Jöns [614]. Jordanus [701]. José [478]. Josef [683]. Joseph [855, 947, 961, 1192, 776, 216, 703, 50, 158, 1028, 1160, 420]. Josephi [356]. Josephus [971]. Josiah [1116]. Jouannet [1037]. Journal [650, 95]. Jr [1116]. Jr. [308]. Ju [1141]. Ju-kô [1141]. Juan [949, 1202]. Judaeus [798]. Judaism [1052]. Juden [106]. Julian [766, 850, 634, 962]. Julius [1231, 369, 484, 682, 906, 1160, 363]. July [60, 964, 1149]. Junge [150]. jusqu’à [394].

kô [1141]. Kaga [916]. Kagaku [620]. Kaizo [1109]. kakke [1107]. kakke-byô [1107]. Kalender [1139, 957]. källskrifter [764]. Kamal [661, 959]. Kamil [683]. kammer [676]. Kammerer [212]. Kant [612]. Kao [1153]. Karel [124]. Karker [626]. Karl [1119, 488]. Karsten [792]. Kast [1116]. Kate [1121]. Katherine [561, 624]. Kato [772]. Katsh [1041]. Kazue [774]. Keilschrift [992]. Keilschrift-Texte [992]. Keirin [823]. Kekulé [1104]. Keller [507, 514]. Kellers [507]. Kellogg [987]. Kemp [1100]. Kenelm [1034]. kenkyû [1107, 618, 621, 772, 824]. Kennedy [13]. Kenneth [768]. Kensaburo [1110]. Kepler [267, 207]. kept [1017]. Kerotakis [1085]. Kershaw [1076]. Keyser [184]. khabar [1128]. Khandakhadyaka [548]. Khayyâm [1016]. Khowarizmi [358]. Kibre [756, 1046]. Kiernan [1049]. Kimble [1189]. King [14]. Kings [503]. Kingsley [1039]. Kinnosuke [1108, 824, 1109]. kirām [1128]. Kirwan [3]. Kitāb [641]. Kitâb [119]. Kleinhans [1072]. Klibansky [891, 609, 389, 429].

Klimata [57]. Knight [903]. Knonau [507]. Knot [42]. Knowledge [399, 904, 17, 452]. Knyveton [986]. kohhl [525]. Kokushi [619]. Kolumbus [610]. Kolumbus-Originale [610]. Kommentar [761]. Kommentatoren [474, 539, 602]. Königlichen [124]. Königsberg [1138]. Konrad [1188, 629, 969]. konst [757]. Koron [1108]. Kosmos [373]. kotoni [1141]. Koyré [655]. Kozue [917]. Kramp [1020]. Krause [1093]. Krautheimer [817]. Kritisch [131]. Kritisch-historisch [131]. Kroeber [856]. Kryptogamae [106]. Ksl [676]. Kugler [717]. Kultur [721]. Kulturgebietes [111]. kultury [82]. Kunitomo [803]. Künste [157]. Kuo [711, 617, 617]. Kuo-ting [617]. Kurt [75]. Kurtz [498]. Kuwabara [1141]. kyû [773].

L [754, 1233, 109, 1113, 972, 827, 74, 76, 163, 183, 219, 1055]. Labours [488]. Lacépède [483]. Lacroix [1102]. Lactic [644]. Laegeraad [730]. L’æuvre [509]. Lagrange [972]. Lake [507]. Lake-Dwelling [507]. L’Algebra [73]. Lamarck [552, 845]. l’Analyse [39]. l’Analysis [39]. Lancelot [956, 1199]. Land [774, 177]. Landscape [711, 481]. Lane [744]. Langdale [1221]. Langenstein [581]. Language [185, 1058]. langue [231]. Lapérouse [1184]. lapide [417]. Lappic [811]. Lärdomshistoriska [764, 764]. Largest [812]. Largiadèr [507, 507]. Larsell [614]. Larsen [234]. Last [713, 368]. Late [562]. Later [1085]. Latham [686]. Latin [269, 608, 518, 647, 616, 1095, 1015, 1210, 66, 807, 1046]. Latin-Arabic [807]. Latina [609]. Latinità [477]. Laufer [1195]. Launay [482]. Laura [245]. Laurels [853]. Lavoisier [765, 145, 1099, 329, 382]. Law [870, 790]. Lawrence [1007]. Lea [1048]. Lead [1190]. Learning [1028, 266, 791, 945]. Leben [1138]. Lebens [484]. Lebensbeschreibung [759]. Lebensgeschichte [201]. Leclainche [892]. Lecons [1101]. Lectures [486, 703]. Lee [478]. Leeuwenhoek [431]. Lefebvre [826]. Legend [274]. Légende [1175]. Legends [1106, 493]. Leibniz [468]. Leibnizens [468]. Leissner [1235]. L’elettrologia [1002]. l’Encyclopedie [946]. Lenzen [400]. Leon [1180, 755, 310, 1136]. Leona [654]. Leonardo [120]. Leonhard [844, 350, 613]. Leopold [1143, 213, 874]. l’esclavage [826]. Leslie [787]. Lesser [540]. Letters [1043, 258, 1221, 488, 144, 153]. lettre [674]. Leurs [207, 383]. Lever [221]. Levy [634, 448, 857, 242]. Lewis [1200, 529, 561]. L’expérience [781]. l’Extrême [78]. Leyburn [888]. Lhassa [79]. li [360]. Liao [1107, 688]. Libby [656]. Liber [1218, 93]. liberté [821]. Liberty [1201]. Libraries [750]. Library [756, 812, 246, 718]. Libri [429, 582, 152, 1180, 755]. Libro [35, 37]. Libros [464]. Lids [199]. Lieben [690]. Lien [883, 594, 396]. Lien-Teh [594, 883, 396]. Life [988, 991, 1043, 267, 308, 1105, 1186, 873, 103, 714, 880, 974, 24, 488, 1033, 816, 201, 866, 311, 949, 912]. Light [174, 63, 554, 709, 642]. Lignamine [854]. ligure [915]. Lilian [857]. Limits [612]. Lincei [603]. Lindsay [503]. linguae [1072]. Linguistic [401]. Link [422]. Linné [22]. L’introduction [232]. Lippmann [23, 187]. Liquids [1033]. List [60, 585, 424, 616].

Literärgeschichte [219]. Literary [1031, 444]. Literatur [219]. Literature [1076, 42, 649, 273, 850, 204, 443, 66, 1050, 244, 274]. Literatures [443]. Little [431]. Liu [1027]. Lives [853]. Living [772, 85, 1176, 502]. Livingston [663]. livre [1096]. Lloyd [206]. L’oeuvre [958]. Loewe [1052]. L’officine [494]. Logic [128, 559, 313, 500]. Logical [1114, 1058]. Logistic [625]. Lomonosov [944, 685]. London [259, 1212, 233]. long [1210]. Longitude [519]. L’Optique [1178]. Loren [1045]. Loria [314, 315, 915]. Lotharingia [171]. Louis [856, 988, 1106, 483, 549, 550, 551, 845, 611, 423, 168, 482]. Louis-Ferdinand [988]. Love [230, 718]. Love-Songs [718]. Lovejoy [693, 110, 859, 1009]. Lovers [230]. Loving [469]. Low [985, 106]. Lowie [856, 1120]. Loyd [1047]. Lucien [857, 242]. Lucretius [842]. Ludlow [52]. Ludwig [604, 1116]. Ludwik [151]. Lull [12]. lumière [642]. Luminant [905]. Lunar [1176]. Lusitanian [517]. Luteus [1156]. Luth [757]. Luther [524, 357]. Luzzi [463]. Lychnos [764]. Lychnos-Bibliotek [764]. Lychnos. [764]. Lynn [48, 590, 1046]. Lynxes [603].

M [778, 1103, 985, 998, 1043, 306, 723, 309, 1031, 192, 374, 485, 632, 766, 850, 1238, 155, 390, 634, 791, 944, 953, 391, 879, 20, 612, 19, 162, 388, 1074]. M.D. [901]. Mā’ [641]. Māyā [298]. môjin [770]. MacBride [240]. MacCallum [156]. MacCullum [1043]. Macdonald [435, 848, 435, 444]. Mackaness [947]. Mackaye [160]. MacKinney [1045]. MacPike [392, 1098]. Macy [1116]. Madame [169, 987]. Maes [553]. Magazin [219]. Magic [1008, 656, 590]. magiciens [79]. magiques [921]. Magistri [609]. Mahābhārata [538]. Mahāvīra [200]. Mahnke [1142]. Maimon [173]. Maimons [173]. Main [225]. Mairan [392]. Major [341]. Maker [205]. Makers [744]. Makes [923]. Making [563, 518, 647]. Malaria [511]. Malebranche [381]. Malinowski [24]. Malloch [1043]. Malthus [805]. Mammalian [176]. Man [1221, 662, 684, 936, 1168, 1197, 489, 305, 923, 1183, 452, 450]. Manchester [933]. Manchu [1133]. Manchurian [594]. mandchou [1133]. Manners [80]. Manning [52]. Mansur [1093]. Manual [405, 473, 183]. Manuel [394]. Manuscript [417, 1089, 1214]. Manuscripts [898, 552, 750, 1089]. manuscrits [687]. Map [521]. Maps [335, 993]. Maqbûl [45]. Marcel [716]. Marcella [477]. March [194, 286, 409, 1234]. Marci [223]. Marcis [223]. Marco [743, 1137]. Marcus [223]. Marett [451]. Margaret [1039, 452, 656, 167]. Margret [490]. Marguet [318]. Marie [258, 987, 987]. Marin [391, 879]. marine [826]. Marinus [1017, 1017]. Mario [1002]. Marjorie [153]. Mark [374]. marocain [866]. Maroccan [866]. mars [164]. Martha [1222]. Martin [1158, 1113, 1165, 712, 393]. Martyrs [783]. Marvin [821, 1168]. Marx [219]. Mary [1000, 362]. Maryland [246]. Mass [1033]. Massachusetts [574, 133, 715]. Master [364]. Masumi [847]. matematchei [915]. matematica [18, 316]. matematiche [315, 1232]. Matemàtiques [270]. matematiska [763]. Material [1069, 179]. Materialien [84, 22]. Materials [834, 882]. math [474, 539, 602]. math.-naturwiss [474, 539, 602]. Mathematical [1182, 205, 370, 1030, 838, 907, 52, 449, 584, 202, 263, 300, 352, 474, 539, 602]. Mathematical-Physical [1030]. mathematical-scientific [474, 539, 602]. mathematician [866]. Mathematicians [96]. mathématicien [866]. Mathematics [746, 531, 309, 747, 1108, 916, 917, 1109, 646, 893, 96, 673, 173, 338, 1163, 719, 401, 530, 56, 183, 999, 956]. Mathematik [157, 721, 173, 615, 1059]. Mathematiker [262]. mathématiques [210]. Mathematische [992]. mathematischen [615, 721]. Mathematischer [157, 721]. Mathesis [915]. matsu [1141]. Matteo [758]. Matter [25, 30, 58, 89, 1, 26, 32, 59, 91, 115, 136, 142, 170, 275, 282, 198, 220, 250, 276, 289, 326, 332, 344, 375, 283, 291, 327, 333, 346, 411, 437, 454, 460, 513, 376, 413, 438, 455, 462, 534, 569, 597, 637, 515, 536, 570, 599, 667, 697, 735, 794, 639, 669, 698, 737, 831, 863, 895, 928, 795, 833, 864, 896, 965, 1012, 1082, 1124, 930, 966, 1014, 1083, 1150, 1170, 1207, 1241, 1126, 1151, 1171, 1209]. Matthew [244]. Maupertuis [104]. Max [692, 722, 231, 1093, 1096, 16, 387, 525, 1028, 1160, 363, 243]. Maxime [588]. Maximilian [274]. May [113, 664, 1240]. maya [1117, 186]. Mayow [138, 94, 137]. McCaskill [1158]. McKeen [491]. McKie [720, 765]. McMurrich [120]. Mea [988]. Mead [635, 371, 1121]. Meanings [671, 835]. Means [189, 1231]. Measles [604]. measurement [6]. Measures [175, 542, 877]. Mechanical [1225, 628]. Mechanics [964, 723, 919]. Mechanism [501]. Mecque [406]. med [761]. Medallic [8, 43, 71]. médecine [892, 1227, 190]. médecins [769]. Mediaeval [269, 978, 953, 76, 1046, 1047]. Medical [812, 649, 246, 452, 508, 942, 684, 234]. medicina [924]. Medicine [133, 452, 322, 662, 925, 811, 36, 143, 688, 996, 571, 1078, 1121, 810, 901, 574, 1045, 883, 510, 14, 396, 1078, 849, 121, 323]. Medicine-Man [662]. medicinische [219]. Medico [1028, 607]. Medico-Mystical [607]. Medico-Philosophical [1028]. Medieval [270, 882, 1045, 649, 518, 647, 66, 601, 1228, 1048, 616]. Medina [478]. Medizin [53, 219]. Medizingeschichtliches [538]. Medizintexten [77]. Meeting [33, 116, 567, 1080]. Meiner [609]. Meisel [16]. Meisen [556]. Meizi [803]. Melanesia [24]. Member [50]. Members [60]. Mémoires [622, 169, 74, 339, 622, 158, 1055]. Memoirs [1017, 119, 880]. Memoranda [943]. Memorandum [843]. Memorial [486, 1042]. Memorials [51]. Memories [517]. Men [997, 491, 1066, 266, 1234, 997, 999]. Mene [747, 747]. Menelaos [1093]. Menshutkin [944]. mentalité [242]. Mentally [1077]. mercaders [164]. merilo [82]. Merriam [85]. Merrill [1237]. Mersenne [391, 879, 391, 879]. mes [661]. Mesopotamia [443, 512]. Messahalla [47]. Messkunst [6]. Mesuë [463]. Metallurgy [1004, 1005, 1190]. Metals [450]. Metaphysics [1030, 777]. Metcalf [1238]. Meteorological [1003]. Meteorologist [872, 934]. Method [519, 522, 468, 559, 19]. Methode [468, 914]. Methods [263]. Metric [708]. Métrique [182]. Metrology [601]. metsubô [1109]. Metzger [109, 123]. Meyer [1050, 507, 762, 528].

Meyerhof [388, 1160, 1096, 387, 525, 1028]. Meyerson [238]. Mich [470]. Michael [393, 487]. Michalski [1162]. Michigan [881]. Microcosmographia [472]. Microscope [419, 403, 436]. Microscopy [240]. Mid [365]. Mid-Victorian [365]. Middle [1189, 688, 648, 595]. Middot [358, 358]. Mieli [1160, 657]. Migration [1195]. Miguel [214]. Mikhail [685]. Milhaud [819]. Millàs [270]. Millennium [199, 699, 797]. Miller [779, 235]. Millikan [919, 81]. Million [956]. Milne [782]. Mind [1197, 558, 663, 1179, 1095, 1006, 635]. Ming [377]. Mingana [607, 681]. Mining [1004]. Minkin [689]. Minoan [646]. Minor [254]. minorum [1072]. Minutes [1215]. Mirandola [756]. Mirfeld [816]. Mirsky [593]. Miscellaneous [718]. Miscellany [234]. Mishnat [358]. Misprint [740]. Missing [422]. missionaires [505]. Missions [868, 1226]. missionum [1072]. Miszellankodex [77]. Mitchell [686, 724]. mitjana [164]. Mittelalter [786]. Mittelalterliche [126]. Mittelalters [162]. mittelenglischen [77]. Mizuki [917]. Möbius [1165]. Modern [271, 890, 335, 925, 659, 1145, 475, 626, 765, 176, 365]. moderne [826, 317]. Moderns [760]. Mohr [759]. Molecular [919]. molécule [716]. Monde [796, 505, 557]. Mondino [463]. mondo [587]. Monge [482]. Monk [911]. Monroe [983]. Montagu [1074]. Monte [1072, 76]. Monteiro [785]. Montgomery [266]. Monumenta [661]. Morale [321, 320]. Morbis [152]. Morbo [582]. morbus [654]. morchid [525]. Mordecai [202]. Morgan [905]. Morison [729]. Morley [728]. Mornet [586]. Morris [559, 1033, 635]. Morrow [828]. Morton [552, 976]. Mosaic [427]. Moskau [157]. Mossner [984]. Most [554]. Motion [870, 465, 867]. Motives [942]. Motley [1095]. Motte [584]. Moule [1137]. Mount [473]. mouvement [231]. Moxa [773]. Moyen [994, 769, 658]. Mozárabes [107]. Mr [417]. MS [518, 647, 472]. Mstislav [825]. Muhammad [360, 358, 230, 304, 359, 388, 579, 1226, 641]. Müller [1138, 77, 712]. Mumford [1200, 529]. Munich [647]. Munqidh [11, 119]. Munro [1076]. Müntzing [1235]. Musa [358]. Museum [676, 1145, 676, 620, 718]. Museums [157, 148]. Music [5, 15, 1061, 608]. Musical [1024, 127]. Musil [80]. musique [994, 359]. Mutual [1098]. My [991, 949]. Mycological [260]. Mystical [607]. Mystiques [79]. Myth [940]. Mythological [718].

N [1076, 719, 403, 481, 1179, 390, 1035, 944, 953, 359, 764, 518, 1048, 75]. N-Nasr [359]. Nārāyana [334]. Nô [916]. Nabatène [212]. nach [1139]. Nadim [360, 360]. Nafîs [540]. Nagel [559]. Najmīyah [641]. Nakayama [770]. nan [700]. Nantucket [148]. Napier [633]. Nara [676, 676]. Nardius [802]. nascita [972]. Nasir [415]. Nasr [1093, 304, 359]. Nathan [860]. Native [1109]. Natur [538, 154]. Natur- [538]. Natura [898]. Natural [38, 630, 850, 619, 16, 584, 681, 804, 934, 424, 742]. Naturales [712]. naturalibus [585, 93]. Naturalist [364, 155]. naturalistes [505]. Naturalists [466, 853, 1040]. Nature [768, 400, 1000, 63, 591, 483, 242, 586, 558, 1100, 549, 550, 551, 845].

Naturists [239]. Naturphilosophie [499]. Naturvetenskapens [763]. naturwiss [474, 539, 602]. naturwissenschaftliches [363]. Naval [193]. Navegants [164]. Naven [889]. navigation [318]. Near [978]. Neckring [230]. Necromancer [493]. Needham [855, 216]. Needs [634]. Neel [79, 79]. Negro [1004]. nei [18]. Nel [972]. nella [1232]. Nellis [632]. Nelson [1140]. Nemorarius [701]. nenpyô [823]. Nervous [960]. Netherlands [673]. neue [749]. neuf [1084]. Neugebauer [615, 992, 83]. Neuve [968]. Newman [510]. Newton [232, 611, 880, 943, 584, 763, 870, 39, 224, 223, 584, 1035, 123, 305]. Newtons [223]. ni [916, 1141]. Nicholas [302]. Nicolae [429]. Nicolas [655]. Nicolaus [389]. Nicolson [153]. Night [1031]. Nights [967]. Nihon [1108, 621, 770, 823, 917, 1110]. Nils [1021]. Nilsson [1235]. Nilsson-Leissner [1235]. Nine [1015]. Nineteenth [292, 328, 1121, 453]. Ninth [293, 71, 113, 453, 861, 114]. No [542, 1108, 618, 620, 621, 772, 1109, 1141]. Noah [945]. Noda [618]. Noettes [826]. Nogre [757]. Noguchi [268]. nombre [957]. non [1166]. non-civilisés [1166]. Nora [485]. Nordenmark [764]. nördlichen [147]. Nordstrom [764, 727, 761]. Norfolk [193]. Normal [1020]. Norman [81, 541]. North [593, 1187, 24]. North-Western [24]. northern [147]. Northrop [215]. Northwest [632, 1191]. Nortons [417]. nosso [1181]. Notation [302]. notations [210]. Note [648, 233, 835, 1156, 544, 748, 797]. Note-Book [233]. Notes [545, 229, 385, 876, 354, 425, 68, 280, 1157, 203, 100, 476, 710, 303, 813, 1024, 840, 1086, 753, 442, 1177, 149, 1135, 264, 653, 969, 118, 1219, 9, 355, 428, 523, 577, 605, 680, 751, 979, 1026, 1092, 211, 750, 349, 336, 44, 72, 181, 285, 1087, 1215, 939, 614, 1184, 819, 236]. notices [884]. notion [716]. novas [1181]. November [595, 1169]. Number [1228, 185]. Numelin [1075]. Numerals [178]. Nuño [600]. Nürnberg [1139]. Nutation [225]. Nuttall [948]. Nutz [941]. Nykl [230].

O [615, 992, 110, 859, 259, 1155, 1052, 506, 241, 563, 767, 1235, 1236, 1009, 23, 163, 187, 693, 727]. Observationes [1223]. Observations [1086, 572, 222, 604]. obsoleta [585]. Obstetrics [695, 407]. Occult [424]. October [134, 247, 331, 1010]. octobre [64]. oculis [46]. Oculist [843]. oeil [1096]. Oesterley [1052]. OEuvres [430, 1097]. Ogawa [1110]. Ogura [370, 1108, 824, 1109]. okeru [916, 1141]. Olaus [1140]. Old [913, 7, 234]. oldest [143]. Olive [791]. Olmsted [1103]. Olof [614]. Olsen [922]. Oluf [757]. om [730]. Omar [1016, 1226]. Omerique [39]. Omnia [389]. omnibus [93]. One [1210, 143]. Oosterhuis [982]. opdagelser [922]. Open [558]. Opera [609, 389, 585]. Ophthalmology [1110]. Opicinus [817]. opinion [39, 1193]. Oracles [1008]. Oratione [609]. Orbes [655]. Order [254]. Ordinall [417, 278]. ordinis [1072]. Organ [217]. Organic [422]. organisation [1084]. Organization [17]. Organizations [526]. Organon. [1162]. Orient [1107, 847, 78]. Oriental [554]. orientale [394]. Origin [789, 178, 1000, 1194, 302, 1146, 1147, 6, 403]. Original [878]. Originale [610]. Origine [324]. Origines [182, 231, 994, 213, 394, 586]. Origins [792, 301, 725]. Ørjan [922]. Orman [625]. Ornithology [504]. Ornstein [1222]. Osborn [364, 1066, 21]. ostiolis [479]. Other [847, 222, 718]. Otis [1053]. Otto [490]. Oudemans [131]. Our [302, 971, 624, 828]. Outline [808, 299, 351]. Overzicht [131]. Oxford [103, 714]. Oxon [306].

P [990, 1167, 498, 723, 265, 548, 505, 1119, 564, 634, 391, 879, 755, 1060, 613, 340, 388, 496]. P. [470]. Packard [322]. Pagel [942]. Painting [711, 882]. Pal [415]. Palacios [214]. Palencia [107, 386]. Palestine [443, 543]. Palms [543, 415]. Pandemic [566]. Panum [604]. Papers [392, 312, 777, 846, 592, 1204, 650, 313, 500, 530]. Pappus [150]. Papyri [1090, 718, 881]. Papyrus [75, 157, 718, 952, 52, 125]. Paracelsus [981, 982]. Paradise [143]. Paravicus [740]. Pareto [663, 564]. Paris [518]. Parisienne [79]. Parker [52, 988, 960]. Part [827, 103, 714, 619]. partie [231]. Partington [725, 1064]. Pascal [1033, 1033]. Pass [875]. Passage [632]. Passato [314]. Past [85]. Pathology [421]. Paton [891]. Patterson [580]. Paul [169, 692, 1005, 1190, 312, 951, 1137, 777, 105, 265, 592, 83, 313, 500, 530, 74, 1096, 158, 339, 622, 1055, 1160, 1062, 169, 368, 1130, 674]. Paullin [506]. Pavingtiles [1047]. Pavlov [990]. Payson [628]. Pearl [756, 1046]. Pearson [488]. Peary [908]. Peattie [853]. Pecia [687]. Peers [12]. Peet [443]. Peirce [312, 777, 592, 313, 500, 530, 312, 565, 592]. pèlerinage [406]. Pelliot [1137]. pendant [505]. Pennell [948]. Penniman [788]. Pensée [588, 238]. pensiero [587]. Pentamerone [390]. Penultimate [421]. Penzer [390]. Percival [816]. Percy [783]. père [1037]. Pérès [109, 158]. Period [620, 1109, 11, 574]. Perkin [308]. Persian [45]. Personal [1066]. Peru [50, 968]. Peter [604]. Peterborough [1215]. Peterson [366]. Pétra [212]. Pétrof [230]. Petrov [704]. Petrovski [755]. Petrum [1072]. Petty [932]. Petum [970]. peuples [1166]. Pflanzenkunde [126]. Pharmacologist [1129]. Pharmazeutisches [1144]. Phelps [1049]. Phil [854]. Philip [995, 11, 189, 119]. Philippi [738]. Philosopher [842]. Philosophers [671, 835, 996]. Philosophical [848, 681, 1028]. Philosophie [173, 1101, 307]. philosophique [397, 1136]. Philosophorum [369]. Philosophy [371, 784, 1115, 1091, 527, 1030, 1131, 325, 54, 86, 27, 134, 194, 113, 247, 281, 286, 331, 342, 512, 409, 453, 584, 623, 532, 595, 636, 664, 696, 731, 793, 893, 829, 861, 964, 1122, 1010, 1079, 1149, 1169, 1240, 173, 983, 1057, 627, 271, 891]. Phineas [1221]. ph’orum [417]. Photography [1145]. Phyllis [34]. Physical [400, 1030, 1033]. Physician [1019, 1174, 218, 14, 934, 144, 1186]. Physicians [233, 978]. Physico [704]. Physico-Chemical [704]. Physicorum [464]. Physics [919, 944, 964, 659, 1143, 626]. physiologischen [690]. Physiologist [1103]. Physiology [132, 165]. physique [381]. pi [618]. Picard [237, 884]. Pickman [1095]. Pico [756]. Pictorial [937]. Picture [889]. Pierpont [606]. Pierre [588, 550, 553, 104, 232, 657, 987]. Pierre-Maxime [588]. Piney [240]. pingue [836]. Pinot [397]. Pintard [391, 879]. Pioneer [16, 934, 901, 945]. Pioneers [192]. Pizarro [50]. Place [1000]. Plague [566, 594, 932]. Planets [648]. Plant [1235, 726, 421, 1195]. plantes [921]. Plants [432]. Platinum [1005]. Playfair [120]. plenum [1015]. Pletho [899]. Plinius [869]. Pliny [742]. Ploss [692]. pluralisme [317]. pocítalo [82]. Poem [419, 1153, 654]. Poet [842]. Poetry [768]. Poincaré [1018]. point [406]. Points [889]. Pole [1187]. Political [890, 1199]. Polo [1137]. Polycarp [836]. Polydore [203]. Polykarp [836]. Population [1199]. Porte [235]. Portrait [169]. Portraits [996]. Portuguese [1181]. portulans [959]. positivisme [821]. Possession [233]. Possit [518]. Post [974]. Potato [1195]. pour [796, 921, 105]. Power [228, 380, 419, 191]. Practice [787]. Practiced [996]. Prafulla [912]. Prag [124]. Prähistorie [384]. Pre [1005, 1190, 803]. Pre-Columbian [1005, 1190]. Pre-Meizi [803]. Preface [292, 328, 456, 517, 537, 600, 571, 670, 640, 865, 897, 967, 1015, 1210, 1127, 1152, 1172, 1084]. préhistoire [1037]. Prehistoric [555]. Prehistorical [576]. Prehistory [554]. Prelude [554, 920]. premières [1175]. Prentiss [560]. Present [302, 371]. Presente [314]. Presented [856, 435]. President [638]. preste [1181]. Preussischen [1139]. Prevention [594]. Preventive [510]. Priestley [420]. prima [603]. Primitif [1025]. Primitive [1197, 242, 242, 790]. Primitives [857]. Primitivism [693]. Primum [1015]. Prince [661, 1226]. Principali [314]. principaux [658]. principer [763]. Principle [301]. Principles [723, 1163, 584, 215]. prior [678]. prism [223]. Prisma [223]. Pritchard [1008]. Probability [1020]. Problem [426, 981, 147]. Probleme [1111]. Problems [889, 471, 470, 640]. Procedures [1030]. Proceedings [567]. Process [1190, 1085]. processus [580]. Proclus [1094]. Proctor [1000]. Profatius [798]. Professor [308, 764, 349, 1086]. Progress [838, 504]. Prominent [556]. property [179]. Prophet [366]. propos [39, 210]. Prosarezepte [77]. Prospectus [66]. Prosper [129]. Protomedicato [1202]. Pruckner [581]. Pseudo [854]. Pseudo-Apuleius [854]. psychologie [1166]. Psychology [772]. Ptolemaei [356]. Ptolemy [357, 524, 357]. pubblicati [915]. Public [192]. Publication [169, 1087]. publieé [819]. publiés [1055]. Pure [263]. Puritanism [1217]. Pursuit [498]. Pusey [408]. Pyramids [1128].

Q [82]. Qiftī [474, 539, 602]. Qifti [474]. Quadrant [798, 209]. quaedam [585]. Quaestiones [712]. Qualiter [518]. Quantum [723]. Quarterly [720]. que [1181]. Quellen [157, 157]. Quelques [661]. Quervain [941]. questions [1096]. Quetelet [383, 537]. Quilibet [518]. Quine [625]. Quirigua [728]. Quo [1210]. Qurra [10]. Qusayr [337].

R [954, 998, 1161, 1182, 128, 725, 1064, 562, 878, 133, 240, 563, 767, 1159, 322, 451, 230, 559, 1001, 307, 982, 1111, 305, 714, 340, 579, 658, 817, 822]. R. [122]. Rây [912]. rôle [1160, 1222]. Rabbinical [309]. Raccolta [102]. raccolti [915]. Rádl [112]. Rafael [792, 73]. Ragnar [1075]. Rainbow [743, 1132]. raison [307]. Ralegh [713, 1031]. Ralph [341]. Ramón [949, 1202, 12]. Ramsay [1105, 579]. Raoul [1166]. Rattlesnake [766]. Ravenel [260]. Rayhan [579]. Raymond [891, 52]. Raymundus [389, 429, 609]. Rays [806, 1087, 783, 490]. Razi [906]. re [1180, 755]. Read [920]. Readings [132]. Reagent [180]. Reallexikon [166]. Really [1231]. Reason [612, 984]. Réaumur [946]. rebus [585, 93]. Rechenbüchlein [969]. recherches [340, 921]. Rechts [776]. Reciprocals [707]. Recollections [1066, 989, 949]. Record [440]. Recorded [50]. Records [432, 953]. Redivivus [1172]. Reduction [1229]. Reference [379, 20, 512, 409, 453, 532, 595, 636, 664, 731, 793, 893, 829, 861, 964, 1010, 1079, 1149, 1045, 263, 1122, 1169, 1240]. Reflections [989]. Reform [69]. Refraction [206]. Regaud [987]. Reginald [1233, 436, 1049]. Regiomontan [1139]. Regiomontanus [1139, 1138]. Regula [738]. Reid [1186]. Reinhold [1220]. Related [693]. relatifs [397]. Relation [200, 960, 1181, 613, 263, 223]. Relations [700, 947, 54, 383, 207, 1098]. Relationships [1031]. Relativity [1041, 782]. religieux [406]. Religion [792, 54, 320, 791, 612]. Religious [1051, 942]. Remarks [745]. remedies [676]. Remora [7]. Renaissance [1001, 982, 1111, 720, 684]. renaixement [164]. Renaud [496, 1160]. rendus [622]. René [391, 879, 78]. Reparaz [164]. Report [733, 116, 251, 293, 894, 1239]. Reports [732, 862, 887]. Representations [937]. Representing [358]. reprobatione [1218]. Republic [686, 694]. Rerum [203, 898]. Research [652, 752, 938, 1134, 516, 576, 1067, 426]. Respiration [138, 94]. Revelations [580]. Review [338, 588, 271, 75, 157, 314, 315, 316, 495, 615, 992, 692, 695, 754, 756, 768, 778, 783, 787, 788, 789, 790, 792, 842, 844, 849, 856, 860, 886, 887, 888, 889, 890, 891, 923, 924, 940, 1100, 1103, 1120, 941, 954, 956, 980, 983, 985, 986, 988, 989, 990, 991, 997, 998, 1004, 1005, 1006, 1007, 1008, 1032, 1034, 1038, 1043, 1047, 1073, 1076, 1161, 1167, 1168, 1190, 1193, 1197, 1198]. Review [1200, 1201, 1231, 477, 267, 373, 392, 1233, 531, 913, 310, 1077, 498, 1182, 1224, 120, 361, 685, 686, 110, 128, 160, 184, 215, 236, 306, 308, 311, 312, 371, 399, 400, 446, 447, 448, 486, 529, 558, 614, 716, 722, 723, 725, 782, 784, 847, 859, 905, 912, 914, 920, 1064, 1099, 1101, 1114, 1115, 1163, 1164, 524, 909, 951, 993, 1178, 1188, 661, 826, 958, 959, 1071, 1137, 1189, 1220, 319, 690, 777, 779, 858, 105, 109, 129, 231, 150, 265, 309, 548, 683, 689, 717, 719, 1050, 1051, 1052, 1059, 1093, 1106, 1227, 1228, 1057]. Review [1222, 492, 994, 320, 1203, 1029, 397, 527, 757, 763, 786, 881, 883, 885, 922, 1148, 182, 915, 1000, 1232, 1054, 369, 663, 617, 370, 1225, 562, 713, 878, 1030, 1031, 1158, 362, 367, 769, 854, 506, 721, 852, 121, 133, 156, 159, 188, 191, 192, 193, 218, 233, 239, 240, 241, 245, 246, 268, 269, 272, 273, 364, 365, 366, 374, 403, 404, 405, 431, 436, 480, 481, 483, 484, 485, 489, 491, 501, 503, 504, 505, 511, 546, 547, 549, 550, 551, 552, 553, 560, 561, 563, 566, 593, 594, 630, 631, 632, 633, 766]. Review [767, 841, 845, 846, 850, 851, 853, 855, 857, 947, 948, 950, 960, 961, 1039, 1040, 1042, 1066, 1075, 1105, 1113, 1116, 1118, 1119, 1159, 1165, 1179, 1183, 1184, 1185, 1186, 1187, 1191, 1192, 1194, 1196, 1221, 1234, 1235, 1236, 1237, 1238, 389, 429, 609, 132, 452, 510, 582, 583, 727, 761, 1140, 395, 152, 155, 165, 266, 322, 323, 324, 407, 408, 508, 662, 715, 762, 925, 398, 401, 499, 659, 919, 1061, 1143, 625, 1058, 451, 776, 127, 230, 235, 386, 390, 445, 995, 1230, 559, 565, 592, 244, 1001, 730, 564, 624, 628]. Review [634, 635, 656, 691, 760, 791, 963, 1035, 1056, 1199, 104, 232, 238, 242, 307, 317, 216, 108, 843, 775, 848, 1162, 942, 1041, 1145, 610, 785, 1009, 337, 827, 944, 953, 1180, 981, 982, 1111, 1142, 103, 123, 153, 305, 321, 391, 526, 611, 714, 819, 879, 880, 943, 1098, 18, 20, 50, 73, 83, 161, 164, 213, 372, 402, 478, 724, 755, 814, 818, 957, 1060, 1117, 1138, 1139, 1181, 1229, 612, 984, 313, 500, 530, 613, 908, 151, 655, 106, 15, 17, 51, 74, 101, 102, 112, 122, 237, 589, 720, 877, 910, 911, 1096, 1107, 1108, 618, 619, 620]. Review [621, 688, 770, 771, 772, 773, 823, 824, 907, 916, 917, 918, 996, 1109, 1110, 1141, 318, 10, 81, 84, 85, 11, 12, 13, 14, 16, 19, 21, 22, 23, 24, 45, 46, 47, 48, 49, 52, 53, 54, 55, 56, 57, 76, 78, 79, 80, 107, 111, 125, 126, 131, 154, 158, 162, 163, 167, 168, 183, 185, 186, 187, 189, 190, 212, 214, 217, 219, 304, 339, 340, 356, 357, 359, 360, 368, 387, 388, 393, 394, 396, 406, 430, 432, 433, 434, 435, 443, 444, 449, 450, 482, 487, 488, 490]. Review [494, 496, 507, 509, 525, 554, 555, 556, 557, 578, 579, 580, 581, 584, 585, 586, 587, 590, 591, 606, 607, 608, 622, 623, 626, 629, 654, 657, 658, 681, 682, 684, 693, 711, 718, 726, 728, 729, 758, 759, 764, 774, 781, 815, 817, 820, 821, 822, 825, 828, 884, 892, 906, 946, 949, 955, 962, 987, 999, 1003, 1027, 1028, 1044, 1049, 1055, 1070, 1072, 1094, 1097, 1102, 1112, 1136, 1160, 1166, 1226, 1036, 1078, 1121, 1223, 1037, 1202, 1068, 1104, 1144, 502, 82, 124, 780, 952, 119, 358, 77, 130, 270, 274]. Review [687, 694, 712, 1048, 945, 660, 341, 493, 1033, 765, 1062, 1063, 166, 234, 616, 816, 882, 1045, 1046, 627, 363, 479, 528, 1095, 1067, 497, 921, 1053, 1065, 1069, 1074, 1146, 1147, 1195, 1002, 243, 1162]. Revolt [110]. Révolution [586, 826]. Revolutionibus [1173]. Révolutions [655]. Rey [1181, 434]. Rhazes [572]. Rhind [52, 75]. Rhine [1006]. Ricci [758]. ricerche [18]. Richard [940, 1071, 878, 546, 1119, 925, 398, 760, 817, 1104, 3, 816, 489]. Richmond [193]. Richtmyer [626]. Rickard [450]. Ridwan [1028]. Rignano [98]. Rigobon [477]. Rim [1032]. Rise [510, 659]. Robert [856, 1120, 988, 686, 306, 547, 919, 103, 714, 372, 402, 81, 47, 606, 256, 1212]. Robie [1174]. Robinson [244, 714, 1094, 546, 720]. Rodolphe [359]. Roentgen [783, 490, 1087]. Roger [1234, 254, 349, 867]. Rogers [663]. Roi [1175]. Rokoko [404]. Roller [919]. Romae [1072]. Roman [511, 589, 1136, 603]. romana [603]. Romance [689, 1060]. Rome [532]. Rondelet [466]. Röntgen [490]. Rose [245]. Rosebushes [222]. Rosenbaum [219]. Rosendahl [767]. Rotation [809]. Rotographs [424]. Roule [549, 550, 551, 845, 483]. Routh [1198]. Roux [820]. Rowan [206]. Royal [306, 233, 676]. Rozhdestvenskaya [953]. Rübenzucker [23]. Rudbeck [1140]. Rudolf [1058]. Rudwin [274]. Rufinus [255]. Ruijû [619]. Ruins [728]. Ruler [699, 797]. Rumford [715, 872, 873]. Rumphius [869]. Ruska [369, 682, 906, 1160]. Russell [886, 1163]. Russians [613]. Ruth [1238]. Rwala [80]. ryôhô [773]. Ryôkichi [774]. ryû [907]. Rychlík [124]. Rythme [492].

S [790, 886, 1168, 1182, 215, 689, 269, 436, 485, 766, 634, 691, 724, 84, 22, 496, 815, 1072, 660, 866]. Sô [1141]. sûgaku [916, 917, 1109, 824, 918]. Sûgaku-shi [824, 918]. Sa’îd [1226]. Sabinae [609]. sacra [785]. Sacraments [451]. saerlight [730]. Sai [619]. Sai-i [619]. Saiiki [1141]. Saiiki-jin [1141]. Saint [484, 550, 580]. Saint-Pierre [550]. Sakanishi [711]. Salih [338]. Sallet [494]. Salmonsen [374]. Salo [1051]. Salomon [817, 173]. Salts [41]. Salviani [466]. Salze [682]. Same [222]. Samfundet [764]. Samfundets [764]. Samuel [1019, 686, 1040, 729, 365]. sana‘ah [45]. san‘ah [45]. ’Sanaq [578]. Sanctae [609, 580]. Sanders [312, 777, 592, 313, 500, 530]. Sangin [825]. Sanislaw [1162]. sanitaire [406]. sanpô [907]. Sanpei [621]. Santiago [949]. Santillana [587]. Sarton [184, 497]. Sasagawa [773]. Sasuly [722]. Sator [279, 279]. Sator-Arepo-Formel [279]. Saussure [213]. Savages [24]. Savant [1025]. savants [1102]. Saxl [337]. Say [155]. Sbath [1096]. Scales [708]. Sceptical [216]. Schacht [1028, 776]. Schafer [961]. Schatz [676]. Schatz-kammer [676]. Schaumberger [717]. Schevensteen [509, 509]. Schibbye [4]. Schiff [363]. Schmidt [786, 1119]. Schoch [99]. Schöffler [77]. scholasticus [1029]. Schönen [157]. School [1031, 990]. Schrift [1148]. Schriften [581, 124]. Schuhl [588]. Schurmann [1062]. Schuster [484]. Science [60, 295, 414, 495, 783, 860, 954, 1231, 33, 116, 251, 293, 567, 732, 862, 1081, 184, 1057, 1000, 1030, 1217, 1221, 204, 871, 329, 382, 1131, 475, 103, 714, 325, 253, 720, 620, 54, 8, 43, 71, 86, 2, 27, 134, 194, 113, 208, 247, 281, 286, 331, 342, 347, 512, 409, 453, 590, 623, 532, 595, 636, 571, 664, 696, 731, 793, 745, 893, 829, 861, 834, 964, 1122, 1010, 1079, 931, 1149, 1169, 1240, 440, 1213, 424, 551, 353, 434, 1084, 1160, 652, 752, 938, 1134, 997, 259, 516, 271, 997, 184]. Science [447, 448, 1114, 1115, 1054, 159, 366, 491, 850, 851, 1113, 321, 185, 433, 497, 215, 909, 1061, 634, 1035, 81, 48]. Sciences [681, 1010, 914, 109, 304, 586, 657, 656, 17, 231, 955]. Scientific [777, 1222, 269, 846, 948, 925, 559, 55, 758, 66, 1089, 420, 171, 1046, 1215, 201, 474, 539, 602, 913, 526]. scientifico [587]. scientifique [958, 1160]. Scientifiques [622, 959, 505, 74, 158, 339, 1055, 1226, 169]. Scientist [686, 1174, 1056, 934, 1154]. Scientists [556]. Scott [784, 218, 1182]. Scritti [915]. se [82]. Sea [130]. Search [632]. secoli [315, 18]. Second [199, 247, 328, 595, 571, 640, 797, 897, 1010, 1152, 972]. secondo [972]. Secretary [116, 251, 293]. Secrets [839]. Section [295, 414, 652, 752, 938, 1134, 208, 516]. Sections [861, 1079, 1149, 1122]. Seele [243]. Sefström [1021]. Seher [154]. sei [1108]. seine [676]. seiner [981]. seit [23]. Seki [907, 907]. Seki-ryû [907]. Selected [191, 132]. Selection [38, 804]. selectiores [1223]. Self [635]. selon [483]. Semmelweis [988]. Sengupta [548]. Senhor [1181]. Senmut [62]. Senn [841]. sentetsu [996]. September [27, 453, 829]. Series [8, 43, 71]. Service [823, 594]. ses [395, 781]. Sesquicentennial [1021]. Setting [138, 94, 137]. Seven [907, 41]. Seventeenth [1222, 1030, 121, 480, 942, 526, 729, 744, 256, 409, 144]. Seventh [8, 27, 342, 793, 1240, 28]. Seventieth [435]. Sex [432]. Sextonhundratalsstudier [727]. Sexual [188, 24]. Sexus [1188]. sezione [915]. shôshi [1110]. shôzô [996]. shû [996]. Shadow [828]. shakai [823]. Shakespeare [297]. Shao [1027]. shape [426]. Shari [360]. Shaw [633]. Shearer [244]. Sheean [987]. Shepherd [582]. Sherwood [656]. shi [621, 688, 917, 770, 824, 918]. shichibu [907]. shih [617]. shihaku [1141]. shijitsu [772]. Shin [773]. Shina [1141, 688]. shinri [772]. Shinto [772]. Shio [711]. Ship [7, 130, 660]. Shipholder [7]. Shipman [244]. sho [907]. Shoemaker [945]. Short [1064, 1043]. Shoso [676, 676]. Shoso-in [676, 676]. Shryock [925, 1027]. Shuman [1113]. Siddhanta [1229]. Sidney [1186, 244]. sieben [57]. siècle [381, 231, 232, 586, 642, 994, 318, 687]. siècles [505]. Sigerist [452, 1078, 1078]. Sighting [177]. siglos [107]. signe [821]. Significance [337]. Sikes [842]. Simon [607, 441]. Simple [451, 388]. simples [921]. Simplest [530]. Sina [1211]. since [720]. Sinclair [50]. Singer [502]. Singh [719]. Sinister [582]. sino [494]. sino-annamite [494]. Sir [1031, 846, 880, 1019, 380, 932, 901, 1034, 713, 947, 1159, 584]. Sittengesetz [858]. Sitter [373]. sive [582, 654]. Sixteenth [1030, 583, 466]. Sixth [331, 731, 1169]. Sixty [1236]. Skinner [503]. Skłodowska [987]. Skłodowska-Curie [987]. Sloane [1089]. Smeaton [1058]. Smith [531, 1159, 125, 816, 56]. Smithfield [816]. Smyth [860]. Snow [950]. Sobhy [10, 388]. Social [1051, 961, 925, 634, 1035, 823, 406]. Socialized [1078]. societá [915]. Societies [1222]. Society [486, 259, 663, 440, 60, 33, 116, 251, 293, 567, 732, 862, 1081, 1215, 1167, 306, 635]. Sociologia [663]. sociologie [483]. Sociology [691, 904, 564, 821]. Soga [683]. soi [310]. Solomon [358]. Some [68, 7, 1024, 1066, 475, 574, 1087, 432]. Sommaire [108]. Son [324, 1160]. Songs [718]. sono [772]. sotto [915]. Sound [919]. Source [56]. Sources [978, 320, 616]. sourire [307]. Sous [821]. South [913]. Southern [561]. Souvenirs [1017]. Soviet [1078, 851]. sowie [474, 539, 602]. Space [398]. Spanish [209, 37]. Spargo [493]. Special [512, 409, 453, 532, 595, 664, 731, 793, 893, 829, 861, 964, 1010, 1079, 1149, 1045, 263, 636, 1122, 1169, 1240]. Species [1194, 1146]. Spectrochemical [520]. Sphaere [1142]. Sphaerik [1093]. spica [1015]. spicam [1015]. Spiegelberg [858]. Spiers [1033]. Spillman [1105]. Spinoza [227, 527]. Spirit [1075]. Spiritual [828]. Spontaneous [678]. Spot [1086]. Śrīdhara [200]. St [1007, 875, 816]. Staatlichen [157]. Staatsbibliothek [1139, 647]. Stace [399]. Stahl [123]. State [246, 560]. States [175, 561, 1067, 506, 322]. Statistical [19]. Statistics [722]. Statistik [219]. Stefansson [1158, 1112]. Stejneger [844, 613]. Steller [613, 844]. Stendhal [553]. Stephen [645, 13]. Sternbilder [147, 146]. Sterndienst [717]. Sternkunde [161, 717]. Stevenson [524, 357]. Stevin [441, 542]. Stewart [156]. Stiernhielm [761]. Stillman [365]. Stillwell [167]. Stockholmer [77]. Stoff [179, 426]. Stone [323]. store [922]. storia [603, 915, 1232, 18, 315, 924, 587]. Story [1192, 718, 967, 502, 130, 508]. Straits [887]. Strand [205]. Strauss [578]. Streicher [610]. Stroh [585]. Stromata [151]. Strömberg [1220]. Strong [1030]. Structure [1114, 782]. Strunz [981]. Struve [157]. Stuart [562, 1069]. Student [349]. Students [435]. Studi [18]. Studien [157, 1220, 92, 641, 775, 581]. Studier [764]. Studies [555, 1048, 493, 601, 223, 1199, 768, 850, 19, 76]. studii [1072]. Study [495, 805, 236, 1085, 1030, 1031, 630, 760, 984, 1107, 621, 347, 443, 1027, 522, 1089, 886]. Stukeley [880]. Sturt [362]. stycker [757]. suan [618]. Subdivision [705]. Sublime [235]. successful [836]. sugaku [1108]. Sugar [644]. Suggested [889]. Sugimura [774]. Sulba [495]. sulla [915, 18]. Summa [464]. Summary [439, 568]. Sun [1086, 724]. Sunamoto [771]. Super [609]. Supernatural [857]. Supplement [1073]. Supplementary [236]. Surgeon [638, 986, 1036, 156]. Surgical [125]. surnaturel [242]. Survey [889, 374]. Surveyor [774]. Surya [1229]. Surya-Siddhanta [1229]. Suter [262]. Suzuki [910, 911]. Swammerdam [900]. Swedenborg [585, 585]. Swedenborgii [585]. Swedish [1235]. Sydney [821]. Sylvanus [728]. Symbolism [1228]. Symposium [1199]. Syntax [1058]. Syphilidis [582]. Syphilis [940, 324, 654]. Syrian [11, 1049]. System [708, 302, 625, 17, 584]. Système [182, 796]. Szczygiel [265].

T [271, 788, 954, 399, 1189, 231, 854, 547, 398, 714, 372, 402, 478, 443, 449, 450]. Tô [1141]. Tôyô [1107]. Tabarî’s [143]. Tabelle [75]. Table [29, 88, 140, 196, 248, 288, 343, 410, 459, 533, 596, 666, 734, 830, 927, 1011, 1123, 1206, 1020, 457]. Tables [707, 1229, 799, 449]. Tadataka [774]. Tadeusz [404]. Taibutheh [607]. taisei [773]. Takakazu [907]. Tamīmī’s [641]. Tanaka [916]. tandem [1210]. Taniyama [823]. Tannery [169, 105, 74, 158, 622, 1055, 169, 339]. Tao [905]. Tapas [298]. Ta’rīkh [474, 539, 602]. Ta’rīkHal [474]. Ta’rīkHal-Hukamaā [474]. Tarascon [672]. Tarkhan [359]. Taro [770]. Tasman [631]. Tätigkeit [201]. Taught [681]. Taylor [562, 76, 878]. Teacher [61]. Teaching [852, 2, 347, 96]. Teatro [477]. Technics [827, 529]. Technik [162]. techniques [231]. technologie [902]. Technology [1005, 623, 1010]. Tedin [1235]. Teeple [186]. Teh [594, 883, 396]. Teikyo [1141]. Teikyo-shihaku [1141]. Teitaro [910, 911]. Tekel [747]. Telescope [971]. Temple [999]. Tenth [638, 843]. Tenth-Century [843]. Teorie [314]. Tercentenary [267]. Terrae [1071]. Terre [968]. Testimonials [814]. Tetsukichi [916]. Text [202, 296]. Texte [992]. Texts [718]. Th [219]. Thâbit [10]. Thalès [105]. Thayer [1186]. Their [1077, 1000, 943, 1098, 1147, 474, 539, 602]. thenn [757]. Theodore [103, 612, 47]. Theodosius [1194, 1146]. Theology [1094]. Théophile [994]. Theophrast [841]. Theophrastea [1220]. Theophrastus [981]. théorème [1018]. théorie [642]. Theorien [749]. Theories [188, 302, 232, 1238, 112]. Theorists [5]. Theory [787, 38, 399, 937, 258, 221, 465, 801, 809, 540, 1041, 804, 263, 223, 642, 1120, 400]. There [838]. Thermal [354, 425]. Thibet [79]. Thierfelder [219]. Things [1176, 502]. Third [566, 870, 281, 636, 699, 748, 797, 1079, 931, 1172, 199]. Thirteenth [868, 255]. Thirtieth [134, 135]. Thirty [572, 325, 194, 247, 281, 286, 331, 342, 409, 453, 1210, 195]. Thirty-Eighth [409]. Thirty-Fifth [325]. Thirty-First [194, 195]. Thirty-Fourth [286]. Thirty-Ninth [453]. Thirty-One [1210]. Thirty-Second [247]. Thirty-Seventh [342]. Thirty-Sixth [331]. Thirty-Third [281]. Thirty-Three [572]. Thofaïl [1136]. Thomas [913, 436, 552, 766, 948, 662, 634, 1181, 183, 393, 739, 1174, 132, 583, 155]. Thompson [744, 713, 561, 715, 822, 1044, 882, 1048]. Thoms [695, 407]. Thomson [989, 150]. Thorndike [1046, 48, 590]. Thorp [1039]. Thought [256, 480, 1183, 1001, 48, 984, 55]. Three [889, 1158, 572, 1098]. Tiahuanacu [1117]. t’ien [617]. t’ien-chih [617]. Tierie [361]. Time [7, 415, 867, 300, 352, 828, 398]. Timepieces [519]. Timerding [721]. Times [1121, 1077, 511]. ting [617]. Titley [489]. To-Morrow [828]. Tobacco [970, 1070]. Tobias [1054, 185]. Together [542]. tokushu [1108]. tokushu-sei [1108]. Tokyo [620]. Toledo [107]. Tolhopf [651]. Tomb [62, 1016]. Tome [359]. Tomooki [773]. Toribio [478]. Torlais [946]. Torres [887]. Torrey [445]. Torsten [775]. Toulouse [169]. Toussoun [1226]. Town [583]. Tractatus [739, 518, 647]. Tracts [467]. Tracy [239]. Trail [1113]. Trail-Blazers [1113]. Training [911]. Traité [1099]. Transference [805]. Transformation [591]. Translation [298, 5, 418, 582, 798, 45]. Translator [740]. translators [474, 539, 602]. Transmutation [227]. Transporting [519]. Trattato [663]. Trattner [1161]. Travels [948]. travers [505, 998]. Treasure [676, 36]. Treasurer [116, 251, 293]. Treasures [681]. Treatise [970, 899, 548, 36, 519, 867, 277, 1128, 45, 877]. Treatises [467, 907, 1033]. Treatment [1077]. Trenchard [611]. Trend [722]. Tres [429, 582]. Trevithick [489]. Triangulation [519]. Tribe [889]. Tribute [435]. Trier [472]. Tropfke [1059]. Trotula [65]. True [1067]. Tsûshô [1141]. Ts’an [277]. Tsuruichi [918]. Tudor [562, 562]. Tuhfat [1128, 496]. T’ung [277]. Turajeff [157]. Turba [369]. Turk [1211]. Turkish [810]. turques [108]. Turtle [973]. Twentieth [1198, 453, 745]. Twenty [86, 27, 113, 967, 1015, 28, 87, 114]. Twenty-Eight [967]. Twenty-Eighth [86, 87]. Twenty-Nine [1015]. Twenty-Ninth [113, 114]. Twenty-Seventh [27, 28]. Two [258, 878, 640]. Tyler [1216]. typographica [580]. Tyranny [778]. tzu [700].

Übersetzer [539, 602, 474]. übersetzt [265]. Ugly [1127]. ulum [304]. Ulysses [560]. Umail [641]. Unendliche [1142]. Unidentified [93]. Unification [865]. Union [1078]. Unique [1108]. United [175, 506, 561, 322, 1067]. Universe [448, 1173, 160, 446]. universitaires [687]. University [881, 729, 347]. Unnoticed [867, 1218]. Unpublished [258]. Untersuchungen [223]. Unveiling [1049]. Upharsin [747]. Uplavici [1155]. upon [456, 96]. Uppsala [764]. Urbinas [356]. Usâmah [11].

Usâmah’s [119]. Usamah [119]. Use [1020]. Usher [628]. usité [921]. usque [1210]. úsvite [82]. utgivna [764]. Uthmân [304]. Uyeda [621]. Uzlagh [359].

V [923, 1198, 944, 755, 554, 556, 764, 882]. Vallicrosa [270]. Value [1164]. Vasilij [704]. Vasilyevich [685]. Vassilieff [64]. Vassilievic [64]. Velocity [709]. venarum [479]. Vera [316]. Verbesserung [1093]. Verbreitung [721]. Verein [219]. Verfahren [786]. Verga [102]. Vergangenheit [1148]. Vergil [203]. Verhältnis [223]. Vernon [987]. Verona [654]. Verrill [1053]. Version [388, 358]. Versions [269]. versus [571, 670]. Vertebrate [273]. Verzeichnis [111]. Vesalii [980]. Vesalius [980]. vétérinaire [892]. Vetter [82]. VI [777, 231, 103, 685]. VI-VII [685]. Viatorum [67]. Victor [400, 553, 83]. Victorian [365]. vie [818, 866, 168, 821]. vieram [1181]. Viets [133]. view [426, 889, 962]. Views [137]. VII [103, 1240, 685]. VIII [464]. Vilfredo [663]. Vilhjalmur [1158, 1112]. Villamil [305]. Vincent [1228, 987]. Vinci [120]. Vinciana [102]. vingt [1084]. Virdung [702, 800]. Virgil [493]. Virgile [397]. Virgilian [493]. Virginia [121]. Vis [606]. Vision [1168]. Vitamins [374, 787]. Vizier [14]. Vladimirovic [704]. vocabulaires [231]. Vogel [75]. vois [557]. Vol [169, 459, 533, 596, 666, 734, 830, 927, 1011, 1123, 1206, 486, 912, 1137, 505, 761, 827, 103, 714, 83, 622, 1152, 270, 157]. Vol.I [587]. Vols [685, 457]. Volta [1002]. Voltaire [656]. Volume [29, 88, 31, 140, 196, 90, 139, 141, 248, 197, 249, 288, 343, 287, 290, 345, 410, 412, 458, 461, 535, 598, 665, 668, 736, 832, 926, 929, 1013, 1125, 1205, 1208, 315, 887, 1076, 777, 779, 617, 402, 313, 500, 530, 292, 328, 456, 517, 537, 600, 571, 693, 670, 640, 865, 897, 967, 1015, 1210, 1127, 1172, 1084, 1042]. Volumes [590]. Vom [243]. Vorgriechische [615]. Vorlesungen [615]. Vorwort [387]. Voss [721]. Voyage [485, 766, 79, 1184]. Voyages [1158, 613, 505]. vue [406].

W [157, 998, 1167, 1231, 399, 309, 1051, 1052, 1030, 854, 240, 489, 566, 948, 950, 1119, 635, 714, 943, 957, 1060, 1229, 613, 720, 219, 260, 1053]. Wên [688]. wa [45]. Waard [391, 879, 781]. Wade [245]. Wagel [1078]. Wagner [1105, 1191]. Wagner-Jauregg [1105]. Wakil [566]. wa’l [641]. Waley [236]. Walker [192, 1237, 19]. Wall [1036]. Wallis [1182]. Walsh [694]. Walter [1004, 1216, 713, 1031, 942, 714, 101]. Walther [154]. Wan [617]. Wandering [1075]. Waraqī [641]. Warden [1196]. Warren [1159]. Warrior [11]. Warthin [218]. Was [348, 1211, 521]. Washington [652, 752, 938, 1134, 516, 812]. Water [706, 122]. Water-Drawing [122]. Waterfield [1233]. Waters [456]. Watson [403, 919]. Waxman [1050]. Weather [633]. Webster [945, 493]. Weight [1033]. Weights [175, 877]. Weinberg [683]. Weiss [155, 312, 777, 592, 313, 500, 530]. Welfare [1035]. Welt [951]. Weltanschauung [981]. Wen [1107]. Wen-jen [1107]. Wentworth [1044]. Werke [474, 539, 602]. Werner [546, 758]. Wesen [1188]. Western [953, 402, 24]. Westfall [1048]. Wethered [1179]. Weyl [558]. Whaling [148]. Wheeler [552, 976]. Whetham [54]. White [880]. Who [1020, 996, 521, 940]. Wickersheimer [769]. Wiedemann [40]. Wieleitner [261]. Wien [1203]. Wijk [957, 957, 1229]. Wilfred [403]. Wilhelm [844, 941, 885, 490]. Wilkins [871]. Will [865]. Willard [625]. Willem [373]. William [692, 308, 993, 150, 4, 156, 552, 846, 1186, 582, 408, 662, 762, 244, 1056, 932, 880, 908, 54, 433, 580, 976, 206, 156, 481]. Wilmer [152]. Wilson [631, 1060]. Wind [679]. Wingate [269]. Winter [881]. Winthrop [1086]. Wirken [1138, 363]. Wisdom [143]. Wislizenus [934]. Wissens [721]. Wissenschaft [353]. Wissenschaften [615, 124]. wissenschaftlichen [201, 484]. Wistar [1038, 1038]. Witch [684]. Witchcraft [1008]. within [612]. Wodlinger [374]. Wolf [623]. Wolfson [527, 49]. Woman. [692]. Women [1121]. Wong [883, 396]. Wood [1040, 973, 273, 843, 1129, 46]. Word [971, 616]. Word-List [616]. Work [988, 1182, 308, 203, 329, 382, 704, 103, 714, 607, 260, 93, 798, 991, 267, 1105]. Worke [417]. Works [685, 779, 944, 974, 823, 585, 474, 539, 602, 816]. World [940, 1137, 402, 1069, 558, 782, 1179, 1230, 372, 402, 443, 584, 1053]. World-Structure [782]. Worship [772]. Wright [506, 240, 152, 80, 579]. Writers [1085]. Writings [878, 191, 1022, 608, 1046]. Wu [883, 594, 396, 1027]. Wyck [582]. Wyndham [121].

X [150, 714, 806, 77, 169, 505]. X-Rays [806]. Xaver [717]. XII [107, 622]. XIIe [866]. XIII [622, 29, 107]. XIIIe [866, 687]. XIIIth [15, 540]. XIth [252]. XIV [1055, 88]. XIVe [994, 687]. XIX [343]. XIXth [8, 43, 71]. XV [202, 140]. XVe [505, 318]. XVI [315, 18, 196]. XVII [248, 542, 315, 18]. XVIIe [642]. XVIII [288]. XVIIIe [231, 232, 586, 381]. XVIIIth [62]. XVIth [673]. XX [53, 410, 457]. XXe [505, 318]. XXI [459, 456]. XXII [533, 517]. XXIII [596, 537, 571]. XXIV [666, 600, 640]. XXIX [1123]. XXth [8, 43, 71]. XXV [734, 670]. XXVI [830]. XXVII [927, 865, 897]. XXVIII [1011]. XXX [1152, 1206, 1127, 1172].

yakuen [621]. yakuen-shi [621]. Yaqdhân [1136]. Year [986, 251, 293, 504, 1191]. Yearbook [1065]. Years [788, 1043, 1233, 1225, 1236, 799, 1063, 852]. Yeldham [852]. Yellow [555, 245]. York [560]. Yougoslave [1017]. Younger [463]. Youssouf [661, 959]. Youthful [872]. Ysander [775]. Yugoslav [1017]. Yverbornes [727].

Zô. [771]. Zaki [1230]. Zeitalter [404]. Zeiten [23]. Zeitschrift [219]. Zeki [338]. Zen [910, 911]. Zeuthen [1055]. Ziegler [155]. Zilboorg [684]. Zinner [161, 1138, 111, 1139]. Zirkle [432, 726]. Zodiac [337]. Zoology [273]. Zoser [14]. Zosimus [1085]. Zoubov [755]. Zrozeni [780]. Zuckererzeugung [1068]. Zuckers [23]. zum [775, 23, 147]. Zuordnung [179]. zur [157, 1220, 369, 721, 1165, 84, 22, 387, 759, 201, 179]. Zürich [507]. Zusammenhang [75]. Zweite [106].

References

Anonymous:1930:FMa

[1] Anonymous. Front matter. Isis, 13(2):269–271, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224645.

Sarton:1930:THS

[2] George Sarton. The teaching of the history of science. Isis, 13(2):272– 297, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224646.

Reilly:1930:RKI

[3] J. Reilly and N. O’Flynn. Richard Kirwan, an Irish chemist of the Eighteenth Century. Isis, 13(2):298–319, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224647.

Hansen:1930:JWS

[4] Axel Hansen. John William Schibbye Johnsson. Isis, 13(2):320–324, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224648.

Farmer:1930:GTM

[5] Henry George Farmer. Greek theorists of music in Arabic translation. Isis, 13(2):325–333, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224649.

Fettweis:1930:EMG

[6] Ewald Fettweis. Über die Entstehung der Messkunst. (German) [On the origin of the art of measurement]. Isis, 13(2):334–339, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224650.

Gudger:1930:SOT

[7] E. W. Gudger. Some old time figures of the shipholder, Echeneis or Remora, holding the ship. Isis, 13(2):340–352, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224651.

Sarton:1930:MIHa

[8] George Sarton. Medallic illustrations of the history of science (first series: XIXth and XXth Centuries) seventh article. Isis, 13(2):353–355, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224652.

Sarton:1930:NC

[9] George Sarton, Miguel Asín Palacios, and E. Kremers. Notes and correspondence. Isis, 13(2):356–363, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224653.

Sarton:1930:BRBa

[10] George Sarton. Book review: The Book of al-dakhîra by Thâbit ibn Qurra; G. Sobhy. Isis, 13(2):364–365, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224654.

Sarton:1930:BRBb

[11] George Sarton. Book review: An Arab-Syrian Gentleman and Warrior in the Period of the Crusades by Usâmah ibn-Munqidh; Philip K. Hitti. Isis, 13(2):365–368, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224655.

Sarton:1930:BRBc

[12] George Sarton. Book review: Ramon Lull by E. Allison Peers. Isis, 13 (2):368–370, February 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224656.

Sarton:1930:BRBd

[13] Georges Sarton. Book review: Stephen Hales by A. E. Clark-Kennedy. Isis, 13(2):370–373, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224657.

Sarton:1930:BRBe

[14] George Sarton. Book review: Imhotep, the Vizier and Physician of King Zoser and Afterwards the Egyptian God of Medicine by Jamieson B. Hurry. Isis, 13(2):373–375, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224658.

S:1930:BRBa

[15] G. S. Book review: A History of Arabian Music to the XIIIth Century by Henry George Farmer. Isis, 13(2):375–376, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224659.

Sarton:1930:BRBf

[16] George Sarton. Book review: A Bibliography of American Natural History. The Pioneer Century, 1769–1865 by Max Meisel. Isis, 13(2):376– 378, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224660.

S:1930:BRBb

[17] G. S. Book review: The Organization of Knowledge and the System of the Sciences by Henry Evelyn Bliss. Isis, 13(2):378–379, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224661.

Pogo:1930:BRBa

[18] A. Pogo. Book review: Studi e ricerche sulla storia della matematica in Italia nei secoli XVI e XVII by Ettore Bortolotti. Isis, 13(2):379–382, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224662.

Sarton:1930:BRBg

[19] George Sarton. Book review: Studies in the History of Statistical Method by Helen M. Walker. Isis, 13(2):382–383, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224663.

Pogo:1930:BRBb

[20] A. Pogo. Book review: Chemical Biographical Reference Book by M. A. Blokh. Isis, 13(2):383–386, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224664.

Sarton:1930:BRBh

[21] George Sarton. Book review: From the Greeks to Darwin by Henry Fairfield Osborn. Isis, 13(2):386–388, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224665.

Sarton:1930:BRBi

[22] George Sarton. Book review: Materialien zur Geschichte der Entomologie bis Linné by F. S. Bodenheimer. Isis, 13(2):388–392, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224666.

Sarton:1930:BRBj

[23] George Sarton. Book review: Geschichte des Zuckers seit den ältesten Zeiten bis zum Beginn der Rübenzucker–Fabrikation by Edmund O. von Lippmann. Isis, 13(2):393–395, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224667.

Sarton:1930:BRBk

[24] George Sarton. Book review: The Sexual Life of Savages in North-Western Melanesia by Bronislaw Malinowski. Isis, 13(2):395–397, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224668.

Anonymous:1930:BMa

[25] Anonymous. Back matter. Isis, 13(2):??, February 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224669.

Anonymous:1930:FMb

[26] Anonymous. Front matter. Isis, 13(3):399, March 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224776.

Sarton:1930:TSC

[27] George Sarton. Twenty-seventh critical bibliography of the history and philosophy of science and the history of civilization. (to September 1929). Isis, 13(3):400–586, March 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224777.

Siegel:1930:AITa

[28] Frances Siegel. Authors’ index to the twenty-seventh bibliography. Isis, 13(3):587–602, March 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224778.

Anonymous:1930:BTCa

[29] Anonymous. Brief table of contents of volume XIII (1929–30). Isis, 13(3):603–604, March 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224779.

Anonymous:1930:BMb

[30] Anonymous. Back matter. Isis, 13(3):??, March 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224780.

Anonymous:1930:VI

[31] Anonymous. Volume information. Isis, 14(1):1, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224375.

Anonymous:1930:FMc

[32] Anonymous. Front matter. Isis, 14(1):3–5, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224376.

B:1930:CMH

[33] F. E. B. Council meeting of the History of Science Society. Isis, 14(1): 6–7, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224377.

Sarton:1930:AP

[34] George Sarton. Aristotle and Phyllis. Isis, 14(1):8–19, May 1930. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224378.

Palacios:1930:LAJ

[35] Miguel Asin Palacios. El “libro de los animales” de Jâh.iz.. Isis, 14(1): 20–54, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224379.

Meyerhof:1930:BTE

[36] Max Meyerhof. The “book of treasure”, an early Arabic treatise on medicine. Isis, 14(1):55–76, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224380.

Perez:1930:LCS

[37] José A. Sánchez Pérez. El Libro de Las Cruces. (Spanish) [The book of the crosses]. Isis, 14(1):77–132, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224381.

Darwin:1930:DTN

[38] Charles Darwin, Alfred Russel Wallace, George Sarton, Charles Lyell, and Jos. D. Hooker. Discovery of the theory of natural selection. Isis, 14(1):133–154, May 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224382.

Pelseneer:1930:OIN

[39] Jean Pelseneer. Une opinion inédite de Newton sur “l’Analyse des Anciens” à propos de l’Analysis geometrica de Hugo de Omerique. Isis, 14(1):155–165, May 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224383.

Seemann:1930:EWG

[40] H. J. Seemann. Eilhard Wiedemann geb. 1. August 1852–gest. 7. Januar 1928. Isis, 14(1):166–186, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224384.

Thorndike:1930:SSH

[41] Lynn Thorndike. Seven salts of Hermes. Isis, 14(1):187–188, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224385.

Gandz:1930:KHL

[42] Solomon Gandz. The knot in Hebrew literature, or from the knot to the alphabet. Isis, 14(1):189–214, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224386.

Sarton:1930:MIHb

[43] George Sarton. Medallic illustrations of the history of science (first series: XIXth and XXth Centuries) eighth article. Isis, 14(1):215–218, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224387.

Thorndike:1930:NCa

[44] Lynn Thorndike, Jean Capart, Hugh Richardson, George Sarton, and J. Pelseneer. Notes and correspondence. Isis, 14(1):219–228, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224388.

Sarton:1930:BRBl

[45] George Sarton. Book review: A Persian translation of the eleventh century Arabic alchemical treatise ‘Ain as.-s.an‘ah wa ‘aun as.-s.ana‘ah by Maqbûl Ah.mad; B. B. Datta. Isis, 14(1):229–232, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224389.

Sarton:1930:BRBm

[46] George Sarton. Book review: De oculis eorumque egritudinibus et curis by Benevenutus Grassus; Casey A. Wood. Isis, 14(1):232–233, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224390.

Sarton:1930:BRBn

[47] George Sarton. Book review: Chaucer and Messahalla on the Astrolabe by Robert Theodore Gunther. Isis, 14(1):233–235, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224391.

Sarton:1930:BRBo

[48] George Sarton. Book review: Science and Thought in the Fifteenth Century by Lynn Thorndike. Isis, 14(1):235–240, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224392.

Sarton:1930:BRBp

[49] George Sarton. Book review: Crescas’ Critique of Aristotle by Harry Austryn Wolfson. Isis, 14(1):240–244, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224393.

Pogo:1930:BRBc

[50] Alexander Pogo. Book review: The Conquest of Peru as Recorded by a Member of the Pizarro Expedition by Joseph H. Sinclair. Isis, 14(1): 244–250, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224394.

S:1930:BRBc

[51] G. S. Book review: Memorials of Galileo Galilei, 1564–1642 by J. J. Fahie. Isis, 14(1):250–251, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224395.

Sarton:1930:BRBq

[52] George Sarton. Book review: The Rhind Mathematical Papyrus by Arnold Buffum Chace; Ludlow Bull; Henry Parker Manning; Raymond Clare Archibald. Isis, 14(1):251–255, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224396.

Sarton:1930:BRBr

[53] George Sarton. Book review: Die chinesische Medizin zu Beginn des XX. Jahrhunderts und ihr historischer Entwicklungsgang by Franz Hübotter. Isis, 14(1):255–263, May 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224397.

Sarton:1930:BRBs

[54] George Sarton. Book review: A History of Science and Its Relations with Philosophy and Religion by William Cecil Dampier-Whetham. Isis, 14(1):263–265, May 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224398.

Sarton:1930:BRBt

[55] George Sarton. Book review: The Scientific Habit of Thought by Frederick Barry. Isis, 14(1):265–268, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224399.

Sarton:1930:BRBu

[56] George Sarton. Book review: A Source Book in Mathematics by David Eugene Smith. Isis, 14(1):268–270, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224400.

Sarton:1930:BRBv

[57] George Sarton. Book review: Die sieben Klimata und die πóλϵις ϵ̀πίσημoι by Ernst Honigmann. Isis, 14(1):270–276, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224401.

Anonymous:1930:BMc

[58] Anonymous. Back matter. Isis, 14(1):??, May 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224402.

Anonymous:1930:FMd

[59] Anonymous. Front matter. Isis, 14(2):277–279, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224675.

Anonymous:1930:FLM

[60] Anonymous. Fifth list of members of the History of Science Society members elected from January 1, 1929 to July 1, 1930. Isis, 14(2):280– 284, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224676.

Mackinney:1930:BFT

[61] Loren C. Mackinney. Bishop Fulbert: Teacher, administrator, humanist. Isis, 14(2):285–300, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224677.

Pogo:1930:ACD

[62] A. Pogo. The astronomical ceiling-decoration in the tomb of Senmut (XVIIIth Dynasty). Isis, 14(2):301–325, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224678.

Newton:1930:DDL

[63] Isaac Newton and George Sarton. Discovery of the dispersion of light and of the nature of color (1672). Isis, 14(2):326–341, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224679.

Rainoff:1930:AVV

[64] T. Rainoff. Alexandre Vassilievič Vassilieff. (Né le 6 août 1855. Décédé le 6 octobre 1929.). Isis, 14(2):342–348, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224680.

Hurd-Mead:1930:T

[65] Kate Campbell Hurd-Mead. Trotula. Isis, 14(2):349–367, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224681.

Thorndike:1930:PCM

[66] Lynn Thorndike. Prospectus for a corpus of Medieval scientific literature in Latin. Isis, 14(2):368–384, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224682.

Thorndike:1930:HV

[67] Herrn L. Thorndike and Ernst Zinner. Horologium viatorum. Isis, 14 (2):385–387, October 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224683.

Das:1930:SNI

[68] Sukumar Ranjan Das’. Some notes on Indian astronomy. Isis, 14(2):388– 402, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224684.

Krishnaswami:1930:RIC

[69] G. V. Krishnaswami. Reform of the Indian calendars. Isis, 14(2):403– 410, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224685.

Cajori:1930:CAG

[70] Florian Cajori. A century of American geodesy. Isis, 14(2):411–416, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224686.

Sarton:1930:MIHc

[71] George Sarton. Medallic illustrations of the history of science (first series: XIXth and XXth Centuries) ninth article. Isis, 14(2):417–419, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224687.

Thorndike:1930:NCb

[72] Lynn Thorndike, George Sarton, and F. M. Feldhaus. Notes and correspondence. Isis, 14(2):420–424, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224688.

Pogo:1930:BRBd

[73] A. Pogo. Book review: L’Algebra by Rafael Bombelli; Federigo Enriques. Isis, 14(2):425–426, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224689.

S:1930:BRBd

[74] G. S. Book review: Mémoires scientifiques by Paul Tannery; J. L. Heiberg. Isis, 14(2):426–430, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224690.

Archibald:1930:BRB

[75] R. C. Archibald. Book review: Die Grundlagen der ägyptischen Arithmetik in ihrem Zusammenhang mit der 2:n — Tabelle des Papyrus Rhind by Kurt Vogel. Isis, 14(2):430–433, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224691.

Sarton:1930:BRBw

[76] George Sarton. Book review: Studies in Mediaeval Culture by Charles Homer Haskins; Anniversary Essays in Mediaeval History by Charles H. Taylor; John L. La Monte. Isis, 14(2):433–436, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224692.

Thorndike:1930:BRB

[77] Lynn Thorndike. Book review: Aus mittelenglischen Medizintexten; Die Prosarezepte des Stockholmer Miszellankodex X, 90 by Gottfried Müller; Herbert Schöffler. Isis, 14(2):436–437, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224693.

Sarton:1930:BRBx

[78] George Sarton. Book review: Histoire de l’Extrême–Orient by René Grousset. Isis, 14(2):437–441, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224694.

Sarton:1930:BRBy

[79] George Sarton. Book review: Mystiques et magiciens du Thibet by Alexandra David-Neel; Voyage d’une Parisienne à Lhassa by Alexandra David-Neel. Isis, 14(2):441–444, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224695.

Sarton:1930:BRBz

[80] George Sarton. Book review: The Manners and Customs of the Rwala Bedouins by Alois Musil; J. K. Wright. Isis, 14(2):444–446, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224696.

Sarton:1930:BRBaa

[81] George Sarton. Book review: Science and the New Civilization by Robert A. Millikan; Humanism and America by Norman Foerster. Isis, 14(2): 446–449, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224697.

Struik:1930:BRB

[82] Ruth Struik. Book review: Jak se pocítalo a merilo na úsvite kultury by Q. Vetter. Isis, 14(2):449–450, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224698.

Pogo:1930:BRBe

[83] A. Pogo. Book review: Astronomische Chronologie. Vol. I by Paul Victor Neugebauer. Isis, 14(2):450–454, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224699.

Sarton:1930:BRBab

[84] George Sarton. Book review: Materialien zur Geschichte der Entomologie by F. S. Bodenheimer. Isis, 14(2):454–456, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224700.

Sarton:1930:BRBac

[85] George Sarton. Book review: The Living Past by John C. Merriam. Isis, 14(2):456–458, October 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224701.

Sarton:1930:TEC

[86] George Sarton. Twenty-eighth critical bibliography of the history and philosophy of science and of the history of civilization. (to January 1930). Isis, 14(2):459–569, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224702.

Siegel:1930:AITb

[87] Frances Siegel. Author’s index to the twenty-eighth bibliography. Isis, 14(2):570–578, October 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224703.

Anonymous:1930:BTCb

[88] Anonymous. Brief table of contents of volume XIV (1930). Isis, 14 (2):579–581, October 1930. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224704.

Anonymous:1930:BMd

[89] Anonymous. Back matter. Isis, 14(2):??, October 1930. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224705.

Anonymous:1931:VIa

[90] Anonymous. Volume information. Isis, 15(1):1, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224566.

Anonymous:1931:FMa

[91] Anonymous. Front matter. Isis, 15(1):3–6, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224567.

Kraus:1931:SJI

[92] Paul Kraus. Studien zu Jābir ibn Hayyān. Isis, 15(1):7–30, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224568.

Thorndike:1931:UWG

[93] Lynn Thorndike. An unidentified work by Giovanni da’Fontana: Liber de omnibus rebus naturalibus. Isis, 15(1):31–46, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224569.

Patterson:1931:JMCa

[94] T. S. Patterson. John Mayow in contemporary setting. A contribution to the history of respiration and combustion. Isis, 15(1):47–96, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224570. See errata [138].

Pelseneer:1931:JIH

[95] Jean Pelseneer. Un journal inédit de Hooke. Isis, 15(1):97–103, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224571.

Simons:1931:IFM

[96] Lao G. Simons. The influence of French mathematicians at the end of the Eighteenth Century upon the teaching of mathematics in American colleges. Isis, 15(1):104–123, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224572.

Volta:1931:DEC

[97] Alexander Volta and George Sarton. The discovery of the electric Cell (1800). Isis, 15(1):124–157, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224573.

Sarton:1931:ER

[98] George Sarton. Eugenio Rignano (1870–1930). Isis, 15(1):158–162, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224574.

Pogo:1931:CS

[99] Alexander Pogo. Carl Schoch (1873–1929). Isis, 15(1):163–169, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224575.

G:1931:NC

[100] L. G. and George Sarton. Notes and correspondence. Isis, 15(1):170– 172, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224576.

S:1931:BRBa

[101] G. S. Book review: The Âryabhat.îya of Âryabhat.a by Âryabhat.a; Walter Eugene Clark. Isis, 15(1):173–174, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224577.

S:1931:BRBb

[102] G. S. Book review: Raccolta Vinciana by Ettore Verga. Isis, 15(1):174, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224578.

Pelseneer:1931:BRBa

[103] Jean Pelseneer. Book review: Early Science in Oxford, Vol. VI, The Life and Work of Robert Hooke (Part I) by Robert Theodore Gunther; Early Science in Oxford, Vol. VII, The Life and Work of Robert Hooke (Part II). Isis, 15(1):174–177, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224579.

Metzger:1931:BRB

[104] Hélène Metzger. Book review: Maupertuis by Pierre Brunet. Isis, 15 (1):177–179, February 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224580.

G:1931:BRBa

[105] L. G. Book review: Pour l’histoire de la hellène. De Thalès à Empédocle by Paul Tannery; A. Diès. Isis, 15(1):179–180, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224581.

Ruska:1931:BRB

[106] Julius Ruska. Book review: Die Flora der Juden, Band I, Erste Hälfte: Kryptogamae. Acanthaceae — Composaceae. — Zweite Hälfte: Convolvulaceae — Graminaceae by Immanuel Löw. Isis, 15(1):181–182, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224582.

Sarton:1931:BRBa

[107] George Sarton. Book review: Los Mozárabes de Toledo en los siglos XII y XIII by Angel González Palencia. Isis, 15(1):183–187, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224583.

Meyerhof:1931:BRB

[108] Max Meyerhof. Book review: Sommaire des archives turques du Caire by Jean Deny. Isis, 15(1):187–188, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224584.

G:1931:BRBb

[109] L. G. Book review: Les sciences exactes by J. Pérès; La chimie by Hélène Metzger; La biologie by L. Ambard. Isis, 15(1):188–190, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224585.

Davis:1931:BRBa

[110] H. T. Davis. Book review: The Revolt against Dualism by Arthur O. Lovejoy. Isis, 15(1):190–193, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224586.

Sarton:1931:BRBb

[111] George Sarton. Book review: Verzeichnis der astronomischen Handschriften des deutschen Kulturgebietes by Ernst Zinner. Isis, 15(1):193– 195, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224587.

S:1931:BRBc

[112] G. S. Book review: The History of Biological Theories by Emanuel Rádl; E. J. Hatfield. Isis, 15(1):195–196, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224588.

Sarton:1931:TNC

[113] George Sarton. Twenty-ninth critical bibliography of the history and philosophy of science and of the history of civilization. (to may 1930). Isis, 15(1):197–319, February 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224589.

Siegel:1931:AITa

[114] Frances Siegel. Author’s index to the twenty-ninth bibliography. Isis, 15 (1):320–328, February 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224590.

Anonymous:1931:FMb

[115] Anonymous. Front matter. Isis, 15(2):329–330, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224440.

Brasch:1931:AMH

[116] Frederick E. Brasch. Annual meeting of the History of Science Society and report of the Secretary and Treasurer for 1930. Isis, 15(2):331– 335, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224441.

Sarton:1931:BA

[117] George Sarton. Isis account for 1930. Isis, 15(2):336–337, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224442.

S:1931:NC

[118] G. S. Notes and correspondence. Isis, 15(2):338–340, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224443.

Thomson:1931:BRB

[119] William Thomson. Book review: Usâmah’s Memoirs Entitled Kitâb Al-I‘tibâr by Usamah Ibn Munqidh; Philip K. Hitti. Isis, 15(2):341–342, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224444.

Corner:1931:BRB

[120] George W. Corner. Book review: Leonardo da Vinci the Anatomist by J. Playfair McMurrich. Isis, 15(2):342–344, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224445.

Kofoid:1931:BRBa

[121] Charles A. Kofoid. Book review: Medicine in Virginia in the Seventeenth Century by Wyndham B. Blanton. Isis, 15(2):345–349, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224446.

S:1931:BRBd

[122] G. S. Book review: R. D’Acres’s The Art of Water-Drawing. Isis, 15(2): 349–350, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224447.

Pelseneer:1931:BRBb

[123] J. Pelseneer. Book review: Newton, Stahl, Boerhaave et la Doctrine Chimique by Hélène Metzger. Isis, 15(2):351–353, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224448.

Struik:1931:BRB

[124] Ruth Struik. Book review: Herausgegeben von der Königlichen Böhmischen Gesellschaft der Wissenschaften in Prag. Band I. Funktionenlehre by Bernard Bolzanos Schriften; Karel Rychlík. Isis, 15(2):353– 355, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224449.

Sarton:1931:BRBc

[125] George Sarton. Book review: The Edwin Smith Surgical Papyrus by James Henry Breasted. Isis, 15(2):355–367, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224450.

Sarton:1931:BRBd

[126] George Sarton. Book review: Mittelalterliche Pflanzenkunde by Hermann Fischer. Isis, 15(2):367–370, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224451.

Macdonald:1931:BRB

[127] D. B. Macdonald. Book review: Historical Facts for the Arabian Musical Influence by Henry George Farmer. Isis, 15(2):370–372, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224452.

Davis:1931:BRBb

[128] H. T. Davis. Book review: The Logic of Discovery by R. D. Carmichael. Isis, 15(2):373–376, April 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224453.

G:1931:BRBc

[129] L. G. Book review: Essais sur l’histoire de la balistique by Prosper Élie Charbonnier. Isis, 15(2):376–380, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224454.

Thorndike:1931:BRB

[130] Lynn Thorndike. Book review: Greyhounds of the Sea: The Story of the American Clipper Ship by Carl C. Cutler. Isis, 15(2):381, April 1931.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224455.

Sarton:1931:BRBe

[131] George Sarton. Book review: Kritisch-historisch Overzicht der Acarologie by A. C. Oudemans. Isis, 15(2):381–386, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224456.

Larkey:1931:BRB

[132] S. V. Larkey and G. S. Book review: Selected Readings in the History of Physiology by John F. Fulton; Charles C. Thomas. Isis, 15(2):386– 388, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224457.

Kofoid:1931:BRBb

[133] Charles A. Kofoid. Book review: A Brief History of Medicine in Massachusetts by Henry R. Viets. Isis, 15(2):388–391, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224458.

Sarton:1931:TCB

[134] George Sarton. Thirtieth critical bibliography of the history and philosophy of science and of the history of civilization (to October 1930). Isis, 15(2):392–492, April 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224459.

Siegel:1931:AITb

[135] Frances Siegel. Author’s index to the thirtieth bibliography. Isis, 15(2): 493–501, April 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224460.

Anonymous:1931:FMc

[136] Anonymous. Front matter. Isis, 15(3):503, September 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224671.

Patterson:1931:JMCb

[137] T. S. Patterson. John Mayow. In contemporary setting. II. — Mayow’s views on combustion. Isis, 15(3):504–546, September 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224672.

Anonymous:1931:EJM

[138] Anonymous. Errata: John Mayow in contemporary setting. a contribution to the history of respiration and combustion. Isis, 15(3):546, September 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224673. See [94].

Anonymous:1931:VIb

[139] Anonymous. Volume information. Isis, 15(3):547–628, September 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224670.

Anonymous:1931:BTCa

[140] Anonymous. Brief table of contents of volume XV (1931). Isis, 15(3): 629–630, September 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224674.

Anonymous:1931:VIc

[141] Anonymous. Volume information. Isis, 16(1):1, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224346.

Anonymous:1931:FMd

[142] Anonymous. Front matter. Isis, 16(1):3–5, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224347.

Meyerhof:1931:ATB

[143] Max Meyerhof. ‘Alî at-Tabarî’s “Paradise of Wisdom”, one of the oldest Arabic compendiums of medicine. Isis, 16(1):6–54, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224348.

Symcotts:1931:LEP

[144] John Symcotts, Gervase Fullwood, Edw. Johnson, and Frederick George Marcham. Letters of an English physician in the early Seventeenth Century. Isis, 16(1):55–81, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224350.

Davis:1931:BCE

[145] Tenney L. Davis. Boyle’s conception of element compared with that of Lavoisier. Isis, 16(1):82–91, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224351.

Zinner:1931:SAA

[146] Ernst Zinner. Die Sternbilder der alten Aegypter. (German) [The constellations of the Ancient Egyptians]. Isis, 16(1):92–101, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224352.

Pogo:1931:PIN

[147] A. Pogo. Zum Problem der Identifikation der nördlichen Sternbilder der alten Aegypter. (German) [On the problem of identification of the northern constellations of Ancient Egyptians]. Isis, 16(1):102–114, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224353.

Sarton:1931:WMN

[148] George Sarton. The Whaling Museums of New Bedford and Nantucket. Isis, 16(1):115–123, July 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224354.

Mayer:1931:NC

[149] Joseph Mayer and G. S. Notes and correspondence. Isis, 16(1):124– 131, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224355.

Gandz:1931:BRB

[150] Solomon Gandz. Book review: The Commentary of Pappus on Book X of Euclid’s Elements by Gustav Junge; William Thomson. Isis, 16(1): 132–136, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224356.

Rose:1931:BRB

[151] William J. Rose. Book review: Stromata Copernicana by Ludwik A. Birkenmajer. Isis, 16(1):136–138, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224357.

Leake:1931:BRBa

[152] Chauncey D. Leake. Book review: Hieronymi Fracastorii De Contagione et Contagiosis Morbis et Eorum Curatione, Libri III by Wilmer Cave Wright. Isis, 16(1):138–141, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224358.

Pelseneer:1931:BRBc

[153] J. Pelseneer. Book review: Conway Letters by Marjorie Hope Nicolson. Isis, 16(1):141–143, July 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224359.

Sarton:1931:BRBf

[154] George Sarton. Book review: Goethe als Seher und Erforscher der Natur by Johannes Walther. Isis, 16(1):143–145, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224360.

Leake:1931:BRBb

[155] Chauncey D. Leake. Book review: Thomas Say: Early American Naturalist by Harry B. Weiss; Grace M. Ziegler. Isis, 16(1):145–146, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224361.

Kofoid:1931:BRBc

[156] Charles A. Kofoid. Book review: William Stewart Halsted, Surgeon by William George MacCallum. Isis, 16(1):146–148, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224362.

Archibald:1931:BRB

[157] R. C. Archibald. Book review: Quellen und Studien zur Geschichte der Mathematik, Abteilung A: Quellen, vol. I by Mathematischer Papyrus des Staatlichen Museums der Schönen Künste in Moskau; W. W. Struve; B. A. Turajeff. Isis, 16(1):148–155, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224363.

Sarton:1931:BRBg

[158] George Sarton. Book review: Mémoires scientifiques by Paul Tannery; Joseph Pérès. Isis, 16(1):155–157, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224364.

Kofoid:1931:BRBd

[159] Charles A. Kofoid. Book review: The Adventure of Science by Benjamin Ginzburg. Isis, 16(1):157–158, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224365.

Davis:1931:BRBc

[160] H. T. Davis. Book review: The Dynamic Universe by James Mackaye. Isis, 16(1):158–161, July 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224366.

Pogo:1931:BRBa

[161] A. Pogo. Book review: Die Geschichte der Sternkunde by Ernst Zinner. Isis, 16(1):161–167, July 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224367.

Sarton:1931:BRBh

[162] George Sarton. Book review: Die Technik der Antike und des Mittelalters by Franz M. Feldhaus. Isis, 16(1):167–169, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224368.

Sarton:1931:BRBi

[163] George Sarton. Book review: A History of Applied Entomology by L. O. Howard. Isis, 16(1):169–173, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224369.

Pogo:1931:BRBb

[164] A. Pogo. Book review: Catalunya a les mars. Navegants, mercaders i cartògrafs catalans de l’edat mitjana i del renaixement by Gonçal de Reparaz. Isis, 16(1):173–174, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224370.

Leake:1931:BRBc

[165] Chauncey D. Leake. Book review: Physiology by John F. Fulton. Isis, 16(1):174–176, July 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224371.

Welborn:1931:BRB

[166] Mary Catherine Welborn. Book review: Reallexikon der germanischen Altertumskunde by Johannes Hoops. Isis, 16(1):176–178, July 1931. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224372.

Sarton:1931:BRBj

[167] George Sarton. Book review: Incunabula and Americana, 1450–1800 by Margaret Bingham Stillwell. Isis, 16(1):178–181, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224373.

Sarton:1931:BRBk

[168] George Sarton. Book review: La vie des images by Louis Hourticq. Isis, 16(1):181–184, July 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224374.

Anonymous:1931:PPT

[169] Anonymous. Portrait of Paul Tannery (1843–1904), to celebrate the publication of vol. X of his Mémoires Scientifiques by Madame Paul Tannery (Toulouse, 1930). Isis, 16(1):??, July 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224349.

Anonymous:1931:FMe

[170] Anonymous. Front matter. Isis, 16(2):185–187, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224706.

Welborn:1931:LCA

[171] Mary Catherine Welborn. Lotharingia as a center of Arabic and scientific influence in the Eleventh Century. Isis, 16(2):188–199, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224707.

Drecker:1931:HCA

[172] J. Drecker. Hermannus contractus Über das astrolab. Isis, 16(2):200–219, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224708.

Bergmann:1931:SMP

[173] Hugo Bergmann. Salomon Maimons Philosophie der Mathematik. (German) [Salomon Maimon’s philosophy of mathematics]. Isis, 16(2):220– 232, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224709.

Bradley:1931:DAL

[174] James Bradley, Edmond Halley, and George Sarton. Discovery of the aberration of light. Isis, 16(2):233–239+241–265, November 1931. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224710.

Hellman:1931:JET

[175] C. Doris Hellman. Jefferson’s efforts towards the decimalization of united states weights and measures. Isis, 16(2):266–314, November 1931. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224711.

vonBaer:1931:DME

[176] K. E. von Baer and George Sarton. The discovery of the mammalian egg and the foundation of modern embryology. Isis, 16(2):315–377, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224712.

Dalton:1931:FSA

[177] Edward W. Dalton. The first sighting of the Antarctic continent: A critical analysis of Biscoe’s discovery of Enderby Land. Isis, 16(2):378– 392, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224713.

Gandz:1931:OGN

[178] Solomon Gandz. The origin of the Ghubār numerals, or the Arabian abacus and the articuli. Isis, 16(2):393–424, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224714.

Farber:1931:GZS

[179] Eduard Färber. Zur Geschichte der Zuordnung von Stoff und Eigenschaft. (German) [Toward a history of the allocation of material and property]. Isis, 16(2):425–438, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224715.

Nierenstein:1931:EHF

[180] M. Nierenstein. The early history of the first chemical reagent. Isis, 16 (2):439–446, November 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224716.

Thorndike:1931:NC

[181] Lynn Thorndike, L. G., and Yoshio Mikami. Notes and correspondence. Isis, 16(2):447–448, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224717.

Hellman:1931:BRB

[182] C. D. Hellman. Book review: Les Origines du Système Métrique by Adrien Favre. Isis, 16(2):449–450, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224718.

Sarton:1931:BRBl

[183] George Sarton. Book review: A Manual of Greek Mathematics by Thomas L. Heath. Isis, 16(2):450–451, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224719.

Davis:1931:BRBd

[184] H. T. Davis. Book review: Humanism and Science by Cassius Jackson Keyser; The History of Science and the New Humanism by George Sarton. Isis, 16(2):451–455, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224720.

Sarton:1931:BRBm

[185] George Sarton. Book review: Number, the Language of Science by Tobias Dantzig. Isis, 16(2):455–459, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224721.

Sarton:1931:BRBn

[186] George Sarton. Book review: Maya Astronomy by John Edgar Teeple. Isis, 16(2):459–462, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224722.

Sarton:1931:BRBo

[187] George Sarton. Book review: Entstehung und Ausbreitung der Alchemie by Edmund O. von Lippmann. Isis, 16(2):462–463, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224723.

Kofoid:1931:BRBe

[188] Charles A. Kofoid. Book review: Early Theories of Sexual Generation by F. J. Cole. Isis, 16(2):463–465, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224724.

Sarton:1931:BRBp

[189] George Sarton. Book review: Ancient Civilizations of the Andes by Philip Ainsworth Means. Isis, 16(2):465–468, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224725.

Sarton:1931:BRBq

[190] George Sarton. Book review: Histoire de la Médecine by Arturo Castiglioni; J. Bertrand; F. Gidon. Isis, 16(2):468–471, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224726.

Kofoid:1931:BRBf

[191] Charles A. Kofoid. Book review: Selected Writings, 1877–1930 by D’Arcy Power. Isis, 16(2):471–472, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224727.

Kofoid:1931:BRBg

[192] Charles A. Kofoid. Book review: Pioneers of Public Health by M. E. M. Walker. Isis, 16(2):472–474, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224728.

Kofoid:1931:BRBh

[193] Charles A. Kofoid. Book review: A Century with Norfolk Naval Hospital 1830–1930 by Richmond Cranston Holcomb. Isis, 16(2):474–475, November 1931. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224729.

Sarton:1931:TFC

[194] George Sarton. Thirty-first critical bibliography of the history and philosophy of science and of the history of civilization (to March 1931). Isis, 16 (2):476–575, November 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224730.

Siegel:1931:AITc

[195] Frances Siegel. Author’s index to the thirty-first bibliography. Isis, 16 (2):576–584, November 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224731.

Anonymous:1931:BTCb

[196] Anonymous. Brief table of contents of volume XVI (1931). Isis, 16(2): 585–586, November 1931. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224732.

Anonymous:1932:VIa

[197] Anonymous. Volume information. Isis, 17(1):1, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224622.

Anonymous:1932:FMa

[198] Anonymous. Front matter. Isis, 17(1):3–5, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224623.

Pogo:1932:CCL

[199] A. Pogo. Calendars on coffin lids from Asyut (second half of the Third Millennium). Isis, 17(1):6–24, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224624.

Datta:1932:RMS

[200] Bibhutibhusan Datta. On the relation of Mahāvīra to Śrīdhara. Isis, 17 (1):25–33, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224625.

Birkenmajer:1932:LWT

[201] Alexander Birkenmajer. Zur Lebensgeschichte und wissenschaftlichen Tätigkeit von Giovanni Fontana (1395?-1455?). (German) [On the life history and scientific activity of Giovanni Fontana (1395?–1455?)]. Isis, 17(1):34–53, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224626.

Schub:1932:MTM

[202] Pincus Schub. A mathematical text by Mordecai Comtino (Constantinople, XV Century). Isis, 17(1):54–70, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224627.

Ferguson:1932:NWP

[203] John Ferguson and Elizabeth H. Alexander. Notes on the work of polydore vergil “de inventoribus rerum”. Isis, 17(1):71–93, ???? 1932. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224628.

Maynard:1932:SEE

[204] Katharine Maynard. Science in early English literature 1550 to 1650. Isis, 17(1):94–126, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224629.

Hellman:1932:JBM

[205] C. D. Hellman. John Bird (1709–1776) mathematical instrument-maker in the Strand. Isis, 17(1):127–153, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224630.

Lloyd:1932:DCR

[206] Humphrey Lloyd and George Sarton. Discovery of conical refraction by William rowan Hamilton and Humphrey Lloyd (1833). Isis, 17(1):154– 170, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224631.

Pelseneer:1932:GBG

[207] J. Pelseneer. Gilbert, Bacon, Galilée, Képler, Harvey et Descartes: Leurs relations. Isis, 17(1):171–208, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224632.

Sarton:1932:SHS

[208] George Sarton. Section of the history of science of the Carnegie institution. Isis, 17(1):209–217, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224633.

Vallicrosa:1932:IDC

[209] J. Millàs Vallicrosa. La introducción del cuadrante con cursor en Europa. (spanish) [The introduction of the quadrant with cursor in Europe]. Isis, 17(1):218–258, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224634.

Pelseneer:1932:PNM

[210] J. Pelseneer. A propos de notations mathématiques. Isis, 17(1):259, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224635.

Seydl:1932:NC

[211] Otto Seydl, Mary Catherine Welborn, Y. Mikami, and G. S. Notes and correspondence. Isis, 17(1):260–264, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224636.

Sarton:1932:BRBa

[212] G. Sarton. Book review: Pétra et la Nabatène by A. Kammerer. Isis, 17(1):265–267, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224637.

Pogo:1932:BRB

[213] Alexander Pogo. Book review: Les origines de l’astronomie chinoise by Léopold de Saussure. Isis, 17(1):267–271, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224638.

Sarton:1932:BRBb

[214] George Sarton. Book review: El Islam cristianizado by Miguel Asín Palacios. Isis, 17(1):271–273, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224639.

Davis:1932:BRBa

[215] H. T. Davis. Book review: Science and First Principles by F. S. C. Northrop. Isis, 17(1):273–277, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224640.

Meyer:1932:BRB

[216] Adolf Meyer. Book review: The Sceptical Biologist by Joseph Needham. Isis, 17(1):277–278, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224641.

Sarton:1932:BRBc

[217] George Sarton. Book review: The Organ of the Ancients by Henry George Farmer. Isis, 17(1):278–282, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224642.

Kofoid:1932:BRBa

[218] Charles A. Kofoid. Book review: The Physician of the Dance of Death by Alfred Scott Warthin. Isis, 17(1):282–283, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224643.

Sarton:1932:BRBd

[219] G. Sarton. Book review: Zeitschrift für Geschichte und Literatur der Medizin im Verein mit (1846–48) by Janus; L. Choulant; H. Haeser; J. F. C. Hecker; C. F. Heusinger; F. Jahn; J. C. Marx; J. Rosenbaum; A. W. E. Th. Henschel; Central-Magazin für Geschichte und Literärgeschichte der Medizin, ärztliche Biographik, Epidemiographik, medicinische Geographie und Statistik (1851–52) by Janus H. Bretschneider; A. W. E. Th. Henschel; C. Fr. Heusinger; J. G. Thierfelder. Isis, 17(1):283–284, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224644.

Anonymous:1932:FMb

[220] Anonymous. Front matter. Isis, 17(2):285–287, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224537.

Lenzen:1932:ATL

[221] V. F. Lenzen. Archimedes’ theory of the lever. Isis, 17(2):288–289, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224538.

Vallisnieri:1932:OFR

[222] Antonio Vallisnieri, Mab Bigelow, and R. P. Bigelow. Observations on the fly of rosebushes and other insects of the same habitat. Isis, 17(2): 290–324, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224539.

Rosenfeld:1932:MMU

[223] L. Rosenfeld. Marcus Marcis Untersuchungen über das Prisma und ihr Verhältnis zu Newtons Farbentheorie. (German) [Marcus Marci’s studies of the prism and its relation to Newton’s color theory]. Isis, 17(2):325– 330, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224540.

Pelseneer:1932:DAN

[224] J. Pelseneer. Le dernier autographe de Newton. Isis, 17(2):331, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224541.

Bradley:1932:DMN

[225] James Bradley and George Sarton. Discovery of the main nutation of the Earth’s axis. Isis, 17(2):333–383, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224542.

Archibald:1932:FC

[226] Raymond Clare Archibald. Florian Cajori 1859–1930. Isis, 17(2):384– 407, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224543.

Nierenstein:1932:HST

[227] M. Nierenstein. Helvetius, Spinoza, and transmutation. Isis, 17(2):408– 411, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224544.

Vowles:1932:EEP

[228] Hugh P. Vowles. Early evolution of power engineering. Isis, 17(2):412– 420, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224545.

Birkenmajer:1932:NC

[229] A. Birkenmajer, S. Dickstein, Issei Yamamoto, L. S., Alex. C. Burr, C. W. Adams, and L. G. Notes and correspondence. Isis, 17(2):421– 429, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224546.

Macdonald:1932:BRBa

[230] D. B. Macdonald. Book review: The Dove’s Neckring about Love and Lovers by Abu Muh.ammad ‘Ali ibn H .azim al Andalusi; D. K. Pétrof; A. R. Nykl. Isis, 17(2):430–431, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224547.

G:1932:BRB

[231] L. G. Book review: Histoire de la langue française des origines à 1900. T. VI. Le XVIIIe siècle. Ie partie. Le mouvement des idees et les vocabulaires techniques. Fascicule 2. La langue des sciences by Ferdinand Brunot; Max Fuchs. Isis, 17(2):431–433, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224548.

Metzger:1932:BRBa

[232] Hélène Metzger. Book review: L’introduction des théories de Newton en France au XVIIIe siècle. (Avant 1738) by Pierre Brunet. Isis, 17(2): 433–435, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224549.

Kofoid:1932:BRBb

[233] Charles A. Kofoid. Book review: The Note-Book of Edward Jenner in the Possession of the Royal College of Physicians of London by F. Dawtrey Drewitt. Isis, 17(2):435–436, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224550.

Welborn:1932:BRB

[234] Mary Catherine Welborn. Book review: An Old Icelandic Medical Miscellany by Henning Larsen. Isis, 17(2):436–438, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224551.

Macdonald:1932:BRBb

[235] D. B. Macdonald. Book review: Beyond the Sublime Porte by Barnette Miller. Isis, 17(2):439–440, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224552.

Davis:1932:BRBb

[236] Tenney L. Davis. Book review: Notes on Chinese Alchemy (Supplementary to Johnson’s A Study of Chinese Alchemy) by A. Waley. Isis, 17(2): 440–443, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224553.

S:1932:BRB

[237] G. S. Book review: Éloges et discours académiques by Émile Picard. Isis, 17(2):443–444, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224554.

Metzger:1932:BRBb

[238] Hélène Metzger. Book review: Du cheminement de la pensée by Émile Meyerson. Isis, 17(2):444–445, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224555.

Kofoid:1932:BRBc

[239] Charles A. Kofoid. Book review: American Naturists by Henry Chester Tracy. Isis, 17(2):445–446, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224556.

Kofoid:1932:BRBd

[240] Charles A. Kofoid. Book review: Recent Advances in Microscopy by A. Piney; Basil Graves; E. W. MacBride; R. R. Hewer; E. C. Barton-Wright. Isis, 17(2):446–447, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224557.

Kofoid:1932:BRBe

[241] Charles A. Kofoid. Book review: A History of Entomology by E. O. Essig. Isis, 17(2):447–450, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224558.

Metzger:1932:BRBc

[242] Hélène Metzger. Book review: Le surnaturel et la nature dans la mentalité primitive by Lucien Lévy–Bruhl; La mentalité primitive. Isis, 17 (2):450–453, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224559.

vKlinckowstroem:1932:BRB

[243] Graf Carl v. Klinckowstroem. Book review: Vom Jenseits der Seele by Max Dessoir. Isis, 17(2):453, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224560.

Mayer:1932:BRB

[244] Joseph Mayer. Book review: A Guide to Historical Literature by George Matthew Dutcher; Henry Robinson Shipman; Sidney Bradshaw Fay; Augustus Hunt Shearer; William Henry Allison. Isis, 17(2):453–464, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224561.

Kofoid:1932:BRBf

[245] C. A. Kofoid. Book review: Yellow Fever by Henry Rose Carter; Laura Armistead Carter; Wade Hampton Frost. Isis, 17(2):464–466, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224562.

Kofoid:1932:BRBg

[246] Charles A. Kofoid. Book review: Celebration of the Centennial of the Library of the Medical and Chirurgical Faculty of the State of Maryland, 1830–1930. Isis, 17(2):466–467, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224563.

Sarton:1932:TSC

[247] George Sarton, Alexander Pogo, and Frances Siegel. Thirty-second critical bibliography of the history and philosophy of science and of the history of civilization (to October 1931). Isis, 17(2):468–634, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224564.

Anonymous:1932:BTC

[248] Anonymous. Brief table of contents of volume XVII (1932). Isis, 17(2): 635–636, ???? 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224565.

Anonymous:1932:VIb

[249] Anonymous. Volume information. Isis, 18(1):1, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224473.

Anonymous:1932:FMc

[250] Anonymous. Front matter. Isis, 18(1):3–4, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224474.

Brasch:1932:EAR

[251] F. E. Brasch. Eighth annual report of the Corresponding Secretary and Treasurer of the History of Science Society for the year 1931. Isis, 18 (1):5–6, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224475.

Pogo:1932:AIC

[252] A. Pogo. The astronomical inscriptions on the coffins of Heny (XIth Dynasty?). Isis, 18(1):7–13, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224476.

Russell:1932:HAS

[253] Josiah C. Russell. Hereford and Arabic science in England about 1175– 1200. Isis, 18(1):14–25, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224477.

Welborn:1932:EDA

[254] Mary Catherine Welborn. The errors of the doctors according to Friar Roger Bacon of the Minor Order. Isis, 18(1):26–62, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224478.

Thorndike:1932:RFB

[255] Lynn Thorndike. Rufinus: A forgotten botanist of the Thirteenth Century. Isis, 18(1):63–76, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224479.

Fulton:1932:RBH

[256] J. F. Fulton. Robert Boyle and his influence on thought in the Seventeenth Century. Isis, 18(1):77–102, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224480.

Stimson:1932:BGC

[257] Dorothy Stimson. “Ballad of Gresham Colledge”. Isis, 18(1):103–117, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224481.

Guillemin:1932:EDE

[258] Ernst Adolph Guillemin. Early developments in electromagnetic theory: With two unpublished letters by André Marie Ampère. Isis, 18(1):118– 126, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224482.

Dickinson:1932:JHH

[259] H. W. Dickinson. J. O. Halliwell and the Historical Society of Science (London 1841). Isis, 18(1):127–132, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224483.

Stevens:1932:MWH

[260] Neil E. Stevens. The mycological work of Henry W. Ravenel. Isis, 18(1): 133–149, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224484.

Ruska:1932:HWG

[261] Julius Ruska. Heinrich Wieleitner (geb. 31.X.1874, gest. 27.XII.1931). Isis, 18(1):150–165, July 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224485.

Renaud:1932:ACS

[262] H. P. J. Renaud. Additions et corrections à suter “die mathematiker und astronomen der araber”. Isis, 18(1):166–183, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224486.

Windred:1932:RBP

[263] G. Windred. The relation between pure and applied electrical theory: With special reference to mathematical methods. Isis, 18(1):184–190, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224487.

Partington:1932:NC

[264] J. R. Partington. Notes and correspondence. Isis, 18(1):191–192, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224488.

Gandz:1932:BRB

[265] Solomon Gandz. Book review: Das Buch Job, übersetzt und erklärt by P. Paul Szczygiel. Isis, 18(1):193–195, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224489.

Leake:1932:BRB

[266] C. D. Leake. Book review: The Black Death and Men of Learning by Anna Montgomery Campbell. Isis, 18(1):195–197, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224490.

Bobrovnikoff:1932:BRB

[267] N. T. Bobrovnikoff. Book review: A Tercentenary Commemoration of His Life and Work by Johann Kepler. Isis, 18(1):197–200, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224491.

Kofoid:1932:BRBh

[268] Charles A. Kofoid. Book review: Noguchi by Gustav Eckstein. Isis, 18(1): 200–202, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224492.

Kofoid:1932:BRBi

[269] C. A. Kofoid. Book review: The Mediaeval Latin Versions of the Aristotelian Scientific Corpus by S. D. Wingate. Isis, 18(1):202–203, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224493.

Thorndike:1932:BRBa

[270] Lynn Thorndike. Book review: Assaig d’Història de les Idees Físiques i Matemàtiques a la Catalunya Medieval, Vol. I by J. Millàs Vallicrosa. Isis, 18(1):203–204, July 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224494.

Anger:1932:BRB

[271] Carol Jane Anger. Book review: Philosophy and Modern Science by Harold T. Davis. Isis, 18(1):204–206, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224495.

Kofoid:1932:BRBj

[272] Charles A. Kofoid. Book review: Grosse Biologen by Ernst Almquist. Isis, 18(1):206–207, July 1932. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224496.

Kofoid:1932:BRBk

[273] C. A. Kofoid. Book review: An Introduction to the Literature of Vertebrate Zoology by Casey A. Wood. Isis, 18(1):207–208, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224497.

Thorndike:1932:BRBb

[274] Lynn Thorndike. Book review: The Devil in Legend and Literature by Maximilian Rudwin. Isis, 18(1):208, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224498.

Anonymous:1932:BMa

[275] Anonymous. Back matter. Isis, 18(1):??, July 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224499.

Anonymous:1932:FMd

[276] Anonymous. Front matter. Isis, 18(2):209, October 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224530.

Wu:1932:ACT

[277] Lu-Chiang Wu, Tenney L. Davis, and Wei Po-Yang. An Ancient Chinese treatise on alchemy entitled Ts’an T’ung Ch’i. Isis, 18(2):210–289, October 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224531.

Nierenstein:1932:EAB

[278] M. Nierenstein and P. F. Chapman. Enquiry into the authorship of the Ordinall of Alchimy. Isis, 18(2):290–321, October 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224532.

Darmstaedter:1932:SAF

[279] Ernst Darmstaedter. Die Sator-Arepo-Formel und ihre Erklärung. (German) [The Sator–Arepo formula and its explanation]. Isis, 18(2):322– 329, October 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224533.

Dickstein:1932:NC

[280] A. Dickstein and L. S. Notes and correspondence. Isis, 18(2):330–333, October 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224534.

Sarton:1932:TTC

[281] George Sarton, Alexander Pogo, and Frances Siegel. Thirty-third critical bibliography of the history and philosophy of science and of the history of civilization (to December 1931). Isis, 18(2):334–429, October 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224535.

Anonymous:1932:BMb

[282] Anonymous. Back matter. Isis, 18(2):??, October 1932. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224536.

Anonymous:1933:FMa

[283] Anonymous. Front matter. Isis, 18(3):431, January 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224462.

Das:1933:JC

[284] Sukumar Ranjan Das. The Jaina calendar. Isis, 18(3):432–437, January 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224463.

Webb:1933:NC

[285] R. S. Webb and L. S. Notes and correspondence. Isis, 18(3):438–439, January 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224464.

Sarton:1933:TFC

[286] George Sarton, Alexander Pogo, and Frances Siegel. Thirty-fourth critical bibliography of the history and philosophy of science and of the history of civilization (to March 1932). Isis, 18(3):440–540, January 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224465.

Anonymous:1933:VIa

[287] Anonymous. Volume information. Isis, 18(3):541–612, January 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224461.

Anonymous:1933:BTCa

[288] Anonymous. Brief table of contents of volume XVIII (1932). Isis, 18 (3):613–614, January 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224466.

Anonymous:1933:BMa

[289] Anonymous. Back matter. Isis, 18(3):??, January 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224467.

Anonymous:1933:VIb

[290] Anonymous. Volume information. Isis, 19(1):1, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225179.

Anonymous:1933:FMb

[291] Anonymous. Front matter. Isis, 19(1):3–5, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225180.

Sarton:1933:PNV

[292] George Sarton. Preface to the nineteenth volume of Isis. Isis, 19(1): 6–10, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225181.

Brasch:1933:NAR

[293] F. E. Brasch. Ninth annual report of the Corresponding Secretary and Treasurer of the History of Science Society for the year 1932. Isis, 19(1): 11–12, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225182.

Sarton:1933:BA

[294] George Sarton. Isis’ accounts for 1931 and 1932. Isis, 19(1):13–14, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225183.

Anonymous:1933:SHS

[295] Anonymous. Section of the history of science of the Carnegie institution. Isis, 19(1):15–18, April 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225184.

Thomson:1933:TGC

[296] S. Harrison Thomson. The text of Grosseteste’s de cometis. Isis, 19(1): 19–25, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225185.

Camden:1933:ASD

[297] Carroll Camden, Jr. Astrology in Shakespeare’s day. Isis, 19(1):26– 73, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225186.

Coomaraswamy:1933:TMD

[298] Ananda K. Coomaraswamy. On translation: Māyā, deva, tapas. Isis, 19 (1):74–91, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225187.

Struik:1933:OHDa

[299] D. J. Struik. Outline of a history of differential geometry: I. Isis, 19(1): 92–120, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225188.

Windred:1933:HMTa

[300] G. Windred. The history of mathematical time: I. Isis, 19(1):121– 153, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225189.

Patterson:1933:OGP

[301] Boyd C. Patterson. The origins of the geometric principle of inversion. Isis, 19(1):154–180, April 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225190.

Lattin:1933:OOP

[302] Harriet Pratt Lattin. The origin of our present system of notation according to the theories of Nicholas Bubnov. Isis, 19(1):181–194, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225191.

Graf:1933:NC

[303] Adolf Graf, G. S., and L. G. Notes and correspondence. Isis, 19(1):195– 200, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225192.

Sarton:1933:BRBa

[304] George Sarton. Book review: Ih.s.a’ al-‘ulum. (Enumeration or Classification of the sciences.) by Abu Nasr al-Farabi; Uthmân Muh.ammad Amin. Isis, 19(1):201–203, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225193.

Pelseneer:1933:BRBa

[305] J. Pelseneer. Book review: Newton: The Man by R. de Villamil. Isis, 19(1):203–204, April 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225194.

Davis:1933:BRBa

[306] T. L. Davis. Book review: A Bibliography of the Honourable Robert Boyle, Fellow of the Royal Society by J. F. Fulton; M. A. Oxon. Isis, 19(1):204–205, April 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225195.

Metzger:1933:BRBa

[307] Hélène Metzger. Book review: La philosophie de Fontenelle ou le sourire de la raison by J. R. Carré. Isis, 19(1):205–207, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225196.

Davis:1933:BRBb

[308] Tenney L. Davis. Book review: The Life and Work of Professor William Henry Perkin, Jr. Isis, 19(1):207–208, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225197.

Gandz:1933:BRB

[309] Solomon Gandz. Book review: Rabbinical Mathematics and Astronomy by W. M. Feldman. Isis, 19(1):208–212, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225198.

Brunet:1933:BRB

[310] Pierre Brunet. Book review: De la connaissance de soi by Léon Brunschvicg. Isis, 19(1):212–214, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225199.

Davis:1933:BRBc

[311] H. T. Davis. Book review: The Emergence of Life by John Butler Burke. Isis, 19(1):214–217, April 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225200.

Davis:1933:BRBd

[312] H. T. Davis. Book review: The Collected Papers of Charles Sanders Peirce by Charles Sanders Peirce; Charles Hartshorne; Paul Weiss. Isis, 19(1):217–220, April 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225201.

Quine:1933:BRB

[313] W. V. Quine. Book review: Collected Papers of Charles Sanders Peirce. — Volume II: Elements of Logic by Charles Hartshorne; Paul Weiss. Isis, 19(1):220–229, April 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225202.

Archibald:1933:BRBa

[314] R. C. Archibald. Book review: Il Passato e il Presente delle Principali Teorie Geometriche by G. Loria. Isis, 19(1):229–231, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225203.

Archibald:1933:BRBb

[315] R. C. Archibald. Book review: Storia delle matematiche. Volume 2, I secoli XVI e XVII by G. Loria. Isis, 19(1):231, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225204.

Archibald:1933:BRBc

[316] R. C. Archibald. Book review: Historia de la matemática en España by Francisco Vera. Isis, 19(1):232–233, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225205.

Metzger:1933:BRBb

[317] H. Metzger. Book review: Le pluralisme cohérent de la chimie moderne by Gaston Bachelard. Isis, 19(1):233–235, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225206.

Sapin:1933:BRB

[318] L. Sapin. Book review: Histoire de la navigation du XVe au XXe siècle by F. Marguet. Isis, 19(1):235–237, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225207.

Einarsson:1933:BRB

[319] Stefán Einarsson. Book review: The Cartography of Iceland by Haldór Hermannsson. Isis, 19(1):237–240, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225208.

Guinet:1933:BRB

[320] L. Guinet. Book review: Les deux sources de la morale et de la religion by Henri Bergson. Isis, 19(1):240–241, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225209.

Pelseneer:1933:BRBb

[321] J. Pelseneer. Book review: La Morale de la Science by Albert Bayet. Isis, 19(1):241–245, April 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225210.

Leake:1933:BRBa

[322] C. D. Leake. Book review: History of Medicine in the United States by Francis R. Packard. Isis, 19(1):245–247, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225211.

Leake:1933:BRBb

[323] C. D. Leake. Book review: Medicine among the American Indians by Eric Stone. Isis, 19(1):247–248, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225212.

Leake:1933:BRBc

[324] C. D. Leake. Book review: Histoire de la Syphilis: Son Origine, Son Expansion by E. Jeanselme. Isis, 19(1):249–252, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225213.

Pogo:1933:TFC

[325] Alexander Pogo, George Sarton, and Frances Siegel. Thirty-fifth critical bibliography of the history and philosophy of science and of the history of civilization (to August 1932). Isis, 19(1):253–320, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225214.

Anonymous:1933:BMb

[326] Anonymous. Back matter. Isis, 19(1):??, April 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225215.

Anonymous:1933:FMc

[327] Anonymous. Front matter. Isis, 19(2):321, June 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225103.

Sarton:1933:SPN

[328] George Sarton. Second preface to the nineteenth volume of Isis. Isis, 19(2):322–329, June 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225104.

Meldrum:1933:LEW

[329] A. N. Meldrum. Lavoisier’s early work in science 1763–1771. Isis, 19(2): 330–363, June 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225105.

Birkenmajer:1933:F

[330] Alexander Birkenmajer. Formula. Isis, 19(2):364–378, June 1933. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225106.

Sarton:1933:TSCa

[331] George Sarton, Alexander Pogo, and Frances Siegel. Thirty-sixth critical bibliography of the history and philosophy of science and of the history of civilization (to October 1932). Isis, 19(2):379–470, June 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225107.

Anonymous:1933:BMc

[332] Anonymous. Back matter. Isis, 19(2):??, June 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225108.

Anonymous:1933:FMd

[333] Anonymous. Front matter. Isis, 19(3):471, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224868.

Datta:1933:AN

[334] Bibhutibhusan Datta. The algebra of Nārāyana. Isis, 19(3):472–485, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224869.

Durand:1933:EMM

[335] Dana B. Durand. The earliest modern maps of Germany and central Europe. Isis, 19(3):486–502, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224870.

Talman:1933:NC

[336] C. F. Talman and G. S. Notes and correspondence. Isis, 19(3):503, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224871.

P:1933:BRB

[337] A. P. Book review: The Zodiac of Qus.ayr ’Amra by Fritz Saxl; The Astronomical Significance of the Zodiac of the Qus.ayr ’Amra by Arthur Beer. Isis, 19(3):504–506, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224872.

Adnan:1933:BRB

[338] A. Adnan. Book review: Athar-i-Baqiya, a History of Arabic Mathematics by S.alih. Zeki. Isis, 19(3):506–515, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224873.

Sarton:1933:BRBb

[339] Georges Sarton. Book review: Mémoires scientifiques by Paul Tannery. Isis, 19(3):515–516, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224874.

Sarton:1933:BRBc

[340] George Sarton. Book review: Cinq années de recherches archéologiques en Ethiopie by R. P. Azaïs; R. Chambard. Isis, 19(3):516–518, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224875.

Viets:1933:BRB

[341] Henry R. Viets. Book review: Classic Descriptions of Diseases by Ralph H. Major. Isis, 19(3):518–520, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224876.

Sarton:1933:TSCb

[342] George Sarton and Frances Siegel. Thirty-seventh critical bibliography of the history and philosophy of science and of the history of civilization (to December 1932). Isis, 19(3):521–619, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224877.

Anonymous:1933:BTCb

[343] Anonymous. Brief table of contents of volume XIX (1933). Isis, 19(3): 620–621, September 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224878.

Anonymous:1933:BMd

[344] Anonymous. Back matter. Isis, 19(3):??, September 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224879.

Anonymous:1933:VIc

[345] Anonymous. Volume information. Isis, 20(1):1, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224880.

Anonymous:1933:FMe

[346] Anonymous. Front matter. Isis, 20(1):3–5, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224881.

Sarton:1933:TSH

[347] George Sarton. The teaching and study of the history of science at the University of California. Isis, 20(1):6–14, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224882.

Luckey:1933:WAG

[348] P. Luckey. Was ist ägyptische geometrie? Isis, 20(1):15–52, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224883.

Steele:1933:RBP

[349] Robert Steele. Roger Bacon as professor. A student’s notes. Isis, 20 (1):53–71, November 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224884.

Oldfather:1933:LEE

[350] W. A. Oldfather, C. A. Ellis, and Donald M. Brown. Leonhard Euler’s elastic curves. Isis, 20(1):72–160, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224885.

Struik:1933:OHDb

[351] D. J. Struik. Outline of a history of differential geometry (ii). Isis, 20 (1):161–191, November 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224886.

Windred:1933:HMTb

[352] G. Windred. The history of mathematical time: II. Isis, 20(1):192–219, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224887.

Kotsovsky:1933:AGW

[353] D. Kotsovsky. Das Alter in der Geschichte der Wissenschaft. (German) [Age in the history of science]. Isis, 20(1):220–245, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224888.

Burr:1933:NHC

[354] Alex. C. Burr. Notes on the history of the concept of thermal conductivity. Isis, 20(1):246–259, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224889.

Sarton:1933:NC

[355] George Sarton and L. S. Notes and correspondence. Isis, 20(1):260–266, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224890.

Sarton:1933:BRBd

[356] George Sarton. Book review: Geographiae Codex Urbinas Graecus 82 by Claudii Ptolemaei; Josephi Fischer. Isis, 20(1):267–270, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224891.

Sarton:1933:BRBe

[357] George Sarton. Book review: Geography of Claudius Ptolemy by Ptolemy; Edward Luther Stevenson. Isis, 20(1):270–274, November 1933.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224892.

Thomson:1933:BRB

[358] William Thomson. Book review: The Mishnat ha-Middot, the First Hebrew Geometry of about 150 C. E., and the Geometry of Muhammad Ibn Musa Al-Khowarizmi, the first Arabic Geometry (c. 820), Representing the Arabic Version of the Mishnat ha-Middot by Solomon Gandz. Isis, 20 (1):274–280, November 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224893.

Sarton:1933:BRBf

[359] George Sarton. Book review: La musique arabe Tome 1 by Baron Rodolphe d’Erlanger; Al-Farabi; Abu N-Nas.r Muh.ammad Ibn Muh.ammad Ibn Tarkhan Ibn Uzlagh. Isis, 20(1):280–283, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224894.

Sarton:1933:BRBg

[360] George Sarton. Book review: Al-fihrist li Ibn al-Nadim by Ibn al-Nadim; Shari Muhammad Ali. Isis, 20(1):283–285, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224895.

D:1933:BRB

[361] T. L. D. Book review: Cornelis Drebbel (1572–1633) by G. Tierie. Isis, 20 (1):285–286, November 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224896.

K:1933:BRBa

[362] C. A. K. Book review: Francis Bacon, a Biography by Mary Sturt. Isis, 20(1):286–287, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224897.

Wolf:1933:BRB

[363] Ernst Wolf. Book review: Goethes naturwissenschaftliches Denken und Wirken by Hermann von Helmholtz; Max Dohrn; Julius Schiff. Isis, 20 (1):287–290, November 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224898.

Kofoid:1933:BRBa

[364] C. A. Kofoid. Book review: Cope: Master Naturalist by Henry Fairfield Osborn; Edward Drinker Cope. Isis, 20(1):291–293, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224899.

Kofoid:1933:BRBb

[365] C. A. Kofoid. Book review: Samuel Butler, a Mid-Victorian Modern by Clara G. Stillman. Isis, 20(1):294–296, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224900.

Kofoid:1933:BRBc

[366] C. A. Kofoid. Book review: Huxley, Prophet of Science by Houston Peterson. Isis, 20(1):297–298, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224901.

K:1933:BRBb

[367] C. A. K. Book review: Huxley by Clarence Ayres. Isis, 20(1):298–300, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224902.

Sarton:1933:BRBh

[368] George Sarton. Book review: The Last Cruise of the Carnegie by J. Harland Paul. Isis, 20(1):300–302, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224903.

Holmyard:1933:BRB

[369] E. J. Holmyard. Book review: Turba Philosophorum: Ein Beitrag zur Geschichte der Alchemie by Julius Ruska. Isis, 20(1):302–305, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224904.

Ikehara:1933:BRB

[370] S. Ikehara. Book review: History of Mathematical Education by K. Ogura. Isis, 20(1):305–307, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224905.

Davis:1933:BRBe

[371] H. T. Davis. Book review: The Philosophy of the Present by George Herbert Mead. Isis, 20(1):307–310, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224906.

Pogo:1933:BRB

[372] Alexander Pogo. Book review: The Astrolabes of the World by Robert T. Gunther. Isis, 20(1):310–316, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224907.

Bobrovnikoff:1933:BRB

[373] N. T. Bobrovnikoff. Book review: Kosmos by Willem de Sitter. Isis, 20 (1):316–318, November 1933. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224908.

Kofoid:1933:BRBd

[374] C. A. Kofoid. Book review: Bibliographical Survey of Vitamins, 1650– 1930 by Mark H. Wodlinger; Ella M. Salmonsen. Isis, 20(1):318–319, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224909.

Anonymous:1933:BMe

[375] Anonymous. Back matter. Isis, 20(1):??, November 1933. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224910.

Anonymous:1934:FMa

[376] Anonymous. Front matter. Isis, 20(2):321–323, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225251.

Chang:1934:CME

[377] T. T. Chang. Chia Ming’s elements of dietetics. Isis, 20(2):324–334, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225252.

Thorndike:1934:GFA

[378] Lynn Thorndike. Giovanni da Fontana again. Isis, 20(2):335–336, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225253.

Barnes:1934:BBR

[379] William H. Barnes. Browne’s “Hydriotaphia” with a reference to Adipocere. Isis, 20(2):337–343, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225254.

Cowles:1934:DHPa

[380] Thomas Cowles. Dr. Henry Power, disciple of Sir Thomas Browne. Isis, 20(2):344–366, January 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225255.

Brunet:1934:GDP

[381] Pierre Brunet. Un grand débat sur la physique de Malebranche au XVI-IIe siècle. Isis, 20(2):367–395, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225256.

Meldrum:1934:LEW

[382] A. N. Meldrum. Lavoisier’s early work in science 1763–1771. Isis, 20 (2):396–425, January 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225257.

Collard:1934:GQL

[383] Auguste Collard. Goethe et Quetelet. Leurs relations de 1829 à 1832. Isis, 20(2):426–435, January 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225258.

Lietzmann:1934:GP

[384] W. Lietzmann. Geometrie und Prähistorie. Isis, 20(2):436–439, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225259.

Birkenmajer:1934:NC

[385] Alexander Birkenmajer, Abbé A. Rome, Gino Loria, George Sarton, Edward Kremers, A. Pogo, Lynn Thorndike, Eduard Färber, and F. M. Feldhaus. Notes and correspondence. Isis, 20(2):440–449, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225260.

Macdonald:1934:BRBa

[386] D. B. Macdonald. Book review: Catálogo de las Ciencias by Alfarabi; Angel González Palencia. Isis, 20(2):450–451, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225261.

Sarton:1934:BRBa

[387] George Sarton. Book review: Das Vorwort zur Drogenkunde des Biruni by al-Biruni; Max Meyerhof. Isis, 20(2):451–454, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225262.

Sarton:1934:BRBb

[388] George Sarton. Book review: The Abridged Version of “The Book of Simple Drugs” by Ah.mad ibn Muh.ammad al-Ghafiqi; Gregorius Abu’l-Farag; M. Meyerhof; G. P. Sobhy. Isis, 20(2):454–457, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225263.

Kurland:1934:BRBa

[389] S. Kurland. Book review: Opera Omnia by Nicolaus de Cusa; Apologia doctae ignorantiae by Raymundus Klibansky. Isis, 20(2):457–464, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225264.

Macdonald:1934:BRBb

[390] D. B. Macdonald. Book review: The Pentamerone by Giambattista Basile; Benedetto Croce; N. M. Penzer. Isis, 20(2):464–467, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225265.

Pelseneer:1934:BRB

[391] J. Pelseneer. Book review: Correspondance du P. Marin Mersenne by Mersenne, M.; Cornelis de Waard; René Pintard. Isis, 20(2):467–470, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225266.

Bobrovnikoff:1934:BRB

[392] N. T. Bobrovnikoff. Book review: Correspondence and Papers of Edmond Halley by Halley; D’Ortous de Mairan; Eugene Fairfield MacPike. Isis, 20(2):470–472, January 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225267.

Sarton:1934:BRBc

[393] G. Sarton. Book review: Faraday’s Diary by Michael Faraday; Thomas Martin. Isis, 20(2):472–474, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225268.

Sarton:1934:BRBd

[394] George Sarton. Book review: Manuel d’archéologie orientale depuis les origines jusqu’à l’époque d’Alexandre by G. Contenau. Isis, 20(2):474– 478, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225269.

Lattin:1934:BRB

[395] Harriet P. Lattin. Book review: La catoptromancie grecque et ses dérivés by A. Delatte. Isis, 20(2):478–480, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225270.

Sarton:1934:BRBe

[396] George Sarton. Book review: History of Chinese Medicine by K. Chimin Wong; Wu Lien-teh. Isis, 20(2):480–482, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225271.

Hamilton:1934:BRB

[397] Clarence H. Hamilton. Book review: La Chine et la formation de l’esprit philosophique en France. (1640–1740) by Virgile Pinot; Documents inédits relatifs à la connaissance de la Chine en France de 1685 à 1740. Isis, 20(2):482–484, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225272.

Lenzen:1934:BRBa

[398] V. F. Lenzen. Book review: Time, Space and Atoms by Richard T. Cox. Isis, 20(2):484, January 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225273.

Davis:1934:BRBa

[399] H. T. Davis. Book review: The Theory of Knowledge and Existence by W. T. Stace. Isis, 20(2):485–488, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225274.

Davis:1934:BRBb

[400] H. T. Davis. Book review: The Nature of Physical Theory by Victor F. Lenzen. Isis, 20(2):488–491, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225275.

Lenzen:1934:BRBb

[401] V. F. Lenzen. Book review: Linguistic Analysis of Mathematics by Arthur F. Bentley. Isis, 20(2):491–492, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225276.

Pogo:1934:BRBa

[402] Alexander Pogo. Book review: The Astrolabes of the World by Robert T. Gunther; The Astrolabes of the World Volume II. The Western Astrolabes. Isis, 20(2):492–495, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225277.

Kofoid:1934:BRBa

[403] C. A. Kofoid. Book review: Origin and Development of the Microscope by Alfred N. Disney; Cyril F. Hill; Wilfred E. Watson Baker. Isis, 20 (2):495–497, January 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225278.

Kofoid:1934:BRBb

[404] C. A. Kofoid. Book review: Die Embryologie im Zeitalter des Barock und des Rokoko by Tadeusz Bilikiewicz. Isis, 20(2):497–498, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225279.

Kofoid:1934:BRBc

[405] Charles A. Kofoid. Book review: A Manual of Foraminifera by J. J. Galloway. Isis, 20(2):498–499, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225280.

Sarton:1934:BRBf

[406] George Sarton. Book review: Le pèlerinage de la Mecque au point de vue religieux, social et sanitaire by Le docteur Duguet. Isis, 20(2):499– 503, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225281.

Leake:1934:BRBa

[407] C. D. Leake. Book review: Chapters in American Obstetrics by Herbert Thoms. Isis, 20(2):503–504, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225282.

Leake:1934:BRBb

[408] C. D. Leake. Book review: The History of Dermatology by William Allen Pusey. Isis, 20(2):504–505, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225283.

Sarton:1934:TEC

[409] George Sarton and Frances Siegel. Thirty-eighth critical bibliography of the history and philosophy of science and of the history of civilization (to March 1933, — with special reference to the Seventeenth and Eighteenth Centuries). Isis, 20(2):506–626, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225284.

Anonymous:1934:BTCa

[410] Anonymous. Briet table of contents of volume XX (1933). Isis, 20(2): 627–628, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225285.

Anonymous:1934:BMa

[411] Anonymous. Back matter. Isis, 20(2):??, January 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225286.

Anonymous:1934:VIa

[412] Anonymous. Volume information. Isis, 21(1):1, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224801.

Anonymous:1934:FMb

[413] Anonymous. Front matter. Isis, 21(1):3–4, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224802.

Anonymous:1934:SHS

[414] Anonymous. Section of the history of science of the Carnegie institution. Isis, 21(1):5–7, April 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224803.

Sarton:1934:AFD

[415] George Sarton. The artifical fertilization of date-palms in the time of Ashur-Nasir-Pal B.C. 885–860. Isis, 21(1):8–13, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224804.

Ruska:1934:AAG

[416] Julius Ruska. Die Alchemie des Avicenna. (German) [The alchemy of Avicenna]. Isis, 21(1):14–51, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224805.

Nierenstein:1934:IME

[417] M. Nierenstein and Frances M. Price. The identity of the manuscript entitled “Mr Nortons Worke, de lapide ph’orum” with the Ordinall of Alchimy. Isis, 21(1):52–56, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224806.

Jackson:1934:JET

[418] Henry Jackson, Sanford V. Larkey, and Linda tum Suden. Jackson’s English translation of Berengarius of Carpi’s “Isagogae Breves”, 1660 and 1664. Isis, 21(1):57–70, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224807.

Cowles:1934:DHPb

[419] Thomas Cowles. Dr. Henry Power’s poem on the microscope. Isis, 21(1): 71–80, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224808.

Walker:1934:BSC

[420] W. Cameron Walker. The beginnings of the scientific career of Joseph Priestley. Isis, 21(1):81–97, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224809.

Stevens:1934:PPP

[421] Neil E. Stevens. Plant pathology in the penultimate century. Isis, 21(1): 98–122, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224810.

Nierenstein:1934:MCH

[422] M. Nierenstein. A missing chapter in the history of organic chemistry: The link between elementary analysis by dry-distillation and combustion. Isis, 21(1):123–130, April 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224811.

Sampson:1934:JLE

[423] R. A. Sampson and Frederick E. Brasch. John Louis Emil Dreyer (1852– 1926). Isis, 21(1):131–144, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224812.

Thorndike:1934:CLR

[424] Lynn Thorndike. Check-list of rotographs in the history of natural and occult science. Isis, 21(1):145–168, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224813.

Burr:1934:NHE

[425] Alex. C. Burr. Notes on the history of the experimental determination of the thermal conductivity of gases. Isis, 21(1):169–186, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224814.

Farber:1934:SFP

[426] Eduard Färber. Stoff und Form als Problem der biochemischen Forschung. Eine geschichtliche Betrachtung. (German) [Fabric and shape as a problem in biochemical research. a historical view]. Isis, 21(1):187– 202, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224815.

Partington:1934:DMG

[427] J. R. Partington. The discovery of Mosaic gold. Isis, 21(1):203–206, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224816.

Sarton:1934:NC

[428] George Sarton and L. G. Notes and correspondence. Isis, 21(1):207– 210, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224817.

Kurland:1934:BRBb

[429] S. Kurland. Book review: De Docta Ignorantia Libri Tres by Nicolae de Cusa; Ernestus Hoffman; Raymundus Klibansky. Isis, 21(1):211– 213, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224818.

Sarton:1934:BRBg

[430] George Sarton. Book review: OEuvres complètes by Christiaan Huygens. Isis, 21(1):213–215, April 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224819.

Kofoid:1934:BRBd

[431] Charles A. Kofoid. Book review: Antony van Leeuwenhoek and His “Little Animals” by Clifford Dobell. Isis, 21(1):216–218, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224820.

Sarton:1934:BRBh

[432] George Sarton. Book review: Some Forgotten Records of Hybridization and Sex in Plants, 1716–1739 by Conway Zirkle. Isis, 21(1):218–220, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224821.

Sarton:1934:BRBi

[433] George Sarton. Book review: The Heroic Age of Science by William Arthur Heidel. Isis, 21(1):220–224, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224822.

Sarton:1934:BRBj

[434] George Sarton. Book review: La jeunesse de la science grecque by Abel Rey. Isis, 21(1):224–226, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224823.

Sarton:1934:BRBk

[435] George Sarton. Book review: A Tribute to Duncan Black Macdonald, Consisting of Articles by Former Students, Presented to Him on His Seventieth Birthday, April 9, 1933 by Macdonald. Isis, 21(1):226–227, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224824.

Kofoid:1934:BRBe

[436] C. A. Kofoid. Book review: The History of the Microscope by Reginald S. Clay; Thomas H. Court. Isis, 21(1):227–230, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224825.

Anonymous:1934:BMb

[437] Anonymous. Back matter. Isis, 21(1):??, April 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224826.

Anonymous:1934:FMc

[438] Anonymous. Front matter. Isis, 21(2):231–232, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225109.

Anonymous:1934:SBA

[439] Anonymous. Summary of Isis accounts for 1933. Isis, 21(2):233, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225110.

Simons:1934:HSS

[440] Lao G. Simons and F. E. Brasch. The History of Science Society Council Record. Isis, 21(2):234–240, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225111.

Sarton:1934:SSB

[441] George Sarton. Simon Stevin of Bruges (1548–1620). Isis, 21(2):241– 303, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225112.

Kremers:1934:NC

[442] E. Kremers, L. G., and G. S. Notes and correspondence. Isis, 21(2): 304, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225113.

Sarton:1934:BRBl

[443] George Sarton. Book review: A Comparative Study of the Literatures of Egypt, Palestine, and Mesopotamia. Egypt’s Contribution to the Literature of the Ancient World by T. Eric Peet; The Dawn of Conscience by James Henry Breasted. Isis, 21(2):305–316, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225114.

Sarton:1934:BRBm

[444] George Sarton. Book review: The Hebrew Literary Genius. An Interpretation by Duncan Black Macdonald. Isis, 21(2):316–319, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225115.

Macdonald:1934:BRBc

[445] D. B. Macdonald. Book review: The Jewish Foundation of Islam by Charles Cutler Torrey. Isis, 21(2):319–322, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225116.

Davis:1934:BRBc

[446] H. T. Davis. Book review: The Expanding Universe by Arthur Eddington. Isis, 21(2):322–326, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225117.

Davis:1934:BRBd

[447] H. T. Davis. Book review: The New Background of Science by James Jeans. Isis, 21(2):326–328, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225118.

Davis:1934:BRBe

[448] H. T. Davis. Book review: The Universe of Science by H. Levy. Isis, 21(2):328–330, July 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225119.

Sarton:1934:BRBn

[449] George Sarton. Book review: Tables of the Higher Mathematical Functions by Harold T. Davis. Isis, 21(2):330–334, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225120.

Sarton:1934:BRBo

[450] George Sarton. Book review: Man and Metals by T. A. Rickard. Isis, 21(2):334–336, July 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225121.

Lowie:1934:BRB

[451] Robert H. Lowie. Book review: Sacraments of Simple Folk by R. R. Marett. Isis, 21(2):336–337, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225122.

Larkey:1934:BRBa

[452] S. V. Larkey. Book review: Man and Medicine. An Introduction to Medical Knowledge by Henry E. Sigerist; Margaret Galt Boise. Isis, 21(2): 337–338, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225123.

Sarton:1934:TNC

[453] George Sarton and Frances Siegel. Thirty-ninth critical bibliography of the history and philosophy of science and of the history of civilization (to September 1933, — with special reference to the Nineteenth and Twentieth Centuries). Isis, 21(2):339–486, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225124.

Anonymous:1934:BMc

[454] Anonymous. Back matter. Isis, 21(2):??, July 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225125.

Anonymous:1934:FMd

[455] Anonymous. Front matter. Isis, 21(3):487, September 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225007.

Sarton:1934:PVX

[456] George Sarton. Preface to volume XXI: Casting bread upon the face of the waters. Isis, 21(3):488–501, September 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225008.

Sarton:1934:BTC

[457] George Sarton. Brief table of contents of vols. I to XX of Isis. Isis, 21 (3):502–618, September 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225009.

Anonymous:1934:VIb

[458] Anonymous. Volume information. Isis, 21(3):619–698, September 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225006.

Anonymous:1934:BTCb

[459] Anonymous. Brief table of contents of vol. XXI (1934). Isis, 21(3): 699–700, September 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225010.

Anonymous:1934:BMd

[460] Anonymous. Back matter. Isis, 21(3):??, September 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225011.

Anonymous:1934:VIc

[461] Anonymous. Volume information. Isis, 22(1):1, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225316.

Anonymous:1934:FMe

[462] Anonymous. Front matter. Isis, 22(1):3–7, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225317.

Welborn:1934:MLC

[463] Mary Catherine Welborn. Mondino de’ Luzzi’s commentary on the Canones generales of Mesuë the Younger. Isis, 22(1):8–11, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225319.

Thomson:1934:SBL

[464] S. Harrison Thomson. The summa in VIII Libros Physicorum of Grosseteste. Isis, 22(1):12–18, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225320.

Marx:1934:CTH

[465] Rose S. Marx. A 13th Century theory of heat as a form of motion. Isis, 22(1):19–20, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225321.

Gudger:1934:FGN

[466] E. W. Gudger. The five great naturalists of the Sixteenth Century: Belon, Rondelet, Salviani, Gesner and Aldrovandi: A chapter in the history of ichthyology. Isis, 22(1):21–40, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225322.

Hellman:1934:BTT

[467] C. Doris Hellman. A bibliography of tracts and treatises on the comet of 1577. Isis, 22(1):41–68, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225323.

Petronievics:1934:LMD

[468] Braanislav Petronievics. Über Leibnizens Methode der direkten Differentiation. (German) [On Leibniz’s method of direct differentiation]. Isis, 22(1):69–76, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225324.

Sarton:1934:GCI

[469] George Sarton. George Catlin “Indian-Loving Catlin” (1796–1872). Isis, 22(1):77–94, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225325.

Robbins:1934:GEA

[470] Frank Egleston Robbins. Greco–Egyptian arithmetical problems: P. Mich. 4966. Isis, 22(1):95–103, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225326.

Karpinski:1934:IGA

[471] Louis C. Karpinski. Importance of the Greek algebraical problems. Isis, 22(1):104–105, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225327.

Williams:1934:MTM

[472] John R. Williams. The microcosmographia of trier ms. 1041. Isis, 22 (1):106–135, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225328.

Partington:1934:CAM

[473] J. R. Partington. Chemical arts in the Mount Athos manual of Christian iconography. Isis, 22(1):136–149, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225329.

Kapp:1934:AUK

[474] A. G. Kapp. Arabische Übersetzer und Kommentatoren Euklids, sowie deren math.-naturwiss. Werke auf Grund des Ta’rīkH .al-Hukamaā des Ibn al-Qifti: I. (German) [Arabic translators and commentators of Euclid, and their mathematical-scientific works on the basis of Ta’rīkh al-H .ukamā’ des ibn al-Qift.ī: I]. Isis, 22(1):150–172, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225330.

Peake:1934:SAI

[475] Cyrus H. Peake. Some aspects of the introduction of modern science into China. Isis, 22(1):173–219, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225331.

Gandz:1934:NC

[476] Solomon Gandz, Harry Friedenwald, George Sarton, Max Meyerhof, Thomas F. Dunn, A. P., Lynn Thorndike, J. R. Partington, and Morris C. Leikind. Notes and correspondence. Isis, 22(1):220–232, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225332.

Bevenot:1934:BRB

[477] Hugh Bévenot. Book review: Il Teatro e la Latinità di Hrotsvitha by Marcella Rigobon. Isis, 22(1):233–234, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225333.

Pogo:1934:BRBb

[478] A. Pogo. Book review: The Discovery of the Amazon by José Toribio Medina; Bertram T. Lee; H. C. Heaton. Isis, 22(1):234–238, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225334.

Wolf:1934:BRB

[479] E. Wolf. Book review: De venarum ostiolis 1603 of Hieronymus Fabricius of Aquapendente (1533?-1619) by K. J. Franklin. Isis, 22(1):238–240, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225335.

Kofoid:1934:BRBf

[480] C. A. Kofoid. Book review: The Happy Beast in French Thought of the Seventeenth Century by George Boas. Isis, 22(1):240–241, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225336.

Kofoid:1934:BRBg

[481] Charles A. Kofoid. Book review: William Bartram, Interpreter of the American Landscape by N. Bryllion Fagin. Isis, 22(1):241–242, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225337.

Sarton:1934:BRBp

[482] George Sarton. Book review: Un grand français: Monge by Louis de Launay. Isis, 22(1):242–245, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225338.

Kofoid:1934:BRBh

[483] Charles A. Kofoid. Book review: Lacépède et la sociologie humanitaire selon la nature by Louis Roule. Isis, 22(1):245–246, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225339.

Kofoid:1934:BRBi

[484] Charles A. Kofoid. Book review: Die Anfänge der wissenschaftlichen Erforschung der Geschichte des Lebens durch Cuvier und Geoffroy Saint Hilaire by Julius Schuster. Isis, 22(1):246–248, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225340.

Kofoid:1934:BRBj

[485] Charles A. Kofoid. Book review: Diary of the Voyage of H. M. S. “Beagle.” by Charles Darwin; Nora Barlow. Isis, 22(1):248–251, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225341.

Davis:1934:BRBf

[486] Tenney L. Davis. Book review: Memorial Lectures Delivered before the Chemical Society, 1914–1932. Vol. III. Isis, 22(1):251–252, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225342.

Sarton:1934:BRBq

[487] G. Sarton. Book review: Faraday’s Diary by Michael Faraday. Isis, 22 (1):252–253, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225343.

Sarton:1934:BRBr

[488] George Sarton. Book review: The Life, Letters and Labours of Francis Galton by Karl Pearson. Isis, 22(1):253–255, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225344.

Kofoid:1934:BRBk

[489] Charles A. Kofoid. Book review: Richard Trevithick (1771–1833), the Engineer and the Man by H. W. Dickinson; Arthur Titley. Isis, 22(1): 255–256, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225345.

Sarton:1934:BRBs

[490] George Sarton. Book review: Wilhelm Conrad Röntgen and the Early History of the Roentgen Rays by Otto Glasser; Margret Boveri. Isis, 22 (1):256–259, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225346.

Kofoid:1934:BRBl

[491] Charles A. Kofoid. Book review: American Men of Science by Cattell, J. McKeen; Cattell, Jaques. Isis, 22(1):259–261, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225347.

Greene:1934:BRB

[492] Henry Copley Greene. Book review: Études sur le Rythme Grégorien by Dom J. Jeannin; Rythme Grégorien. Isis, 22(1):261–265, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225348.

W:1934:BRB

[493] M. C. W. Book review: Virgil the Necromancer, Studies in Virgilian Legends by John Webster Spargo. Isis, 22(1):265–267, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225349.

Sarton:1934:BRBt

[494] George Sarton. Book review: L’officine sino-annamite en Annam. I. by Albert Sallet. Isis, 22(1):267–272, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225350.

Archibald:1934:BRB

[495] R. C. Archibald. Book review: The Science of the Sulba. A Study in Early Hindu Geometry by Bibhutibhusan Datta. Isis, 22(1):272–277, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225351.

Sarton:1934:BRBu

[496] George Sarton. Book review: Tuh.fat al-ah.bab by H. P. J. Renaud; Georges S. Colin. Isis, 22(1):277–279, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225352.

Zinner:1934:BRB

[497] Ernst Zinner. Book review: Introduction to the History of Science by George Sarton. Isis, 22(1):279–280, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225353.

Chalmers:1934:BRB

[498] Gordon Keith Chalmers. Book review: The Pursuit of Death by Benjamin P. Kurtz. Isis, 22(1):280–284, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225354.

Lenzen:1934:BRBc

[499] V. F. Lenzen. Book review: Geschichte der Naturphilosophie by Hugo Dingler. Isis, 22(1):284–285, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225355.

Quine:1934:BRB

[500] Williard V. Quine. Book review: Collected Papers of Charles Sanders Peirce.– Volume III: Exact Logic by Charles Sanders Peirce; Charles Hartshorne; Paul Weiss. Isis, 22(1):285–297, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225356.

Kofoid:1934:BRBm

[501] C. A. Kofoid. Book review: The Mechanism of Creative Evolution by C. C. Hurst. Isis, 22(1):297–298, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225357.

Stimson:1934:BRB

[502] Dorothy Stimson. Book review: The Story of Living Things by Charles Singer. Isis, 22(1):298–300, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225358.

Kofoid:1934:BRBn

[503] C. A. Kofoid. Book review: Beaver, Kings and Cabins by Constance Lindsay Skinner. Isis, 22(1):300–301, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225359.

Kofoid:1934:BRBo

[504] Charles A. Kofoid. Book review: Fifty Years’ Progress of American Ornithology, 1883–1933. Isis, 22(1):301–303, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225360.

Kofoid:1934:BRBp

[505] Charles A. Kofoid. Book review: Voyages et découvertes scientifiques des missionaires naturalistes français à travers le monde pendant cinq siècles (XVe à XXe siècles) Encyc. Biol. Vol X by P. Fournier. Isis, 22 (1):303–305, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225361.

Karpinski:1934:BRB

[506] Louis C. Karpinski. Book review: Atlas of the Historical Geography of the United States by Charles O. Paullin; John K. Wright. Isis, 22(1): 305–308, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225362.

Sarton:1934:BRBv

[507] George Sarton. Book review: The Founder of Lake-Dwelling Archaeology by Gerold Meyer von Knonau; Ferdinand Keller; Aus dem Briefwechsel Ferdinand Kellers by Anton Largiadèr; Hundert Jahre Antiquarische Gesellschaft in Zürich by Anton Largiadèr. Isis, 22(1):308–311, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225363.

Leake:1934:BRBc

[508] C. D. Leake. Book review: California’s Medical Story by Henry Harris. Isis, 22(1):312–315, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225364.

Sarton:1934:BRBw

[509] Gearge Sarton. Book review: L’æuvre historique du Dr. Van Schevensteen by A. F. C. Van Schevensteen. Isis, 22(1):315–317, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225365.

Larkey:1934:BRBb

[510] Sanford V. Larkey. Book review: The Rise of Preventive Medicine by George Newman. Isis, 22(1):318–319, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225366.

Kofoid:1934:BRBq

[511] C. A. Kofoid. Book review: The History of Malaria in the Roman Campagna from Ancient Times by Angelo Celli; Anna Celli-Fraentzel. Isis, 22 (1):319–321, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225367.

Sarton:1934:FCB

[512] George Sarton and F. S. Fortieth critical bibliography of the history and philosophy of science and of the history of civilization (to January 1934, — with special reference to antiquity in general, Egypt, and Mesopotamia). Isis, 22(1):322–431, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225368.

Anonymous:1934:BMe

[513] Anonymous. Back matter. Isis, 22(1):??, December 1934. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225369.

Anonymous:1934:FJF

[514] Anonymous. Frontispiece: Johann Ferdinand Keller (1800–1881). Isis, 22(1):??, December 1934. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225318.

Anonymous:1935:FMa

[515] Anonymous. Front matter. Isis, 22(2):433–434, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225126.

Sarton:1935:CIW

[516] George Sarton, Alexander Pogo, Mary Catherine Welborn, and Frances Siegel. Carnegie Institution of Washington. Division of Historical Research. Section of the History of Science. Isis, 22(2):435–439, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225127.

Sarton:1935:PVXa

[517] George Sarton. Preface to volume XXII: Lusitanian memories. Isis, 22 (2):440–455, February 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225128.

Thompson:1935:MCM

[518] Daniel V. Thompson, Jr. Medieval color-making: Tractatus Qualiter Quilibet Artificialis Color Fieri Possit from Paris, B. N., MS. Latin 6749^{b}. Isis, 22(2):456–468, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225129.

Pogo:1935:GFH

[519] A. Pogo. Gemma frisius, his method of determining differences of longitude by transporting timepieces (1530), and his treatise on triangulation (1533). Isis, 22(2):469–506, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225130.

Hamor:1935:DAD

[520] William A. Hamor. David alter and the discovery of spectrochemical analysis. Isis, 22(2):507–510, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225131.

Stanton:1935:AMW

[521] Samuel Mc Coskry Stanton. The Admiral’s map what was it? And who the Admiral? Isis, 22(2):511–515, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225132.

Sorokin:1935:CAI

[522] Pitirim A. Sorokin and Robert K. Merton. The course of Arabian intellectual development, 700–1300 A.D. a study in method. Isis, 22(2): 516–524, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225133.

Sarton:1935:NCa

[523] George Sarton and L. G. Notes and correspondence. Isis, 22(2):525–532, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225134.

Diller:1935:BRB

[524] Aubrey Diller. Book review: Geography of Claudius Ptolemy by Edward Luther Stevenson. Isis, 22(2):533–539, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225135.

Sarton:1935:BRBa

[525] George Sarton. Book review: (Al-morchid fi’l-kohhl) ou Le guide d’oculistique by Max Meyerhof. Isis, 22(2):539–542, February 1935. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225136.

Pelseneer:1935:BRBa

[526] J. Pelseneer. Book review: Scientific Organizations in Seventeenth Century France (1620–1680) by Harcourt Brown. Isis, 22(2):542–543, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225137.

Harap:1935:BRB

[527] Louis Harap. Book review: The Philosophy of Spinoza by Harry Austryn Wolfson. Isis, 22(2):543–546, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225138.

Wolf:1935:BRB

[528] Ernst Wolf. Book review: Ideen und Ideale der biologischen Erkenntnis by Adolf Meyer. Isis, 22(2):546–548, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225139.

Davis:1935:BRBa

[529] H. T. Davis. Book review: Technics and Civilization by Lewis Mumford. Isis, 22(2):548–551, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225140.

Quine:1935:BRB

[530] W. V. Quine. Book review: Collected Papers of Charles Sanders Peirce.– Volume IV: The Simplest Mathematics by Charles Hartshorne; Paul Weiss. Isis, 22(2):551–553, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225141.

Brasch:1935:BRB

[531] Frederick E. Brasch. Book review: A History of Mathematics in America before 1900 by David Eugène Smith; Jekuthiel Ginsburg. Isis, 22(2):553– 556, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225142.

Sarton:1935:FFC

[532] George Sarton and Frances Siegel. Forty-first critical bibliography of the history and philosophy of science and of the history of civilization (to April 1934, — with special reference to Greece and Rome). Isis, 22 (2):557–619, February 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225143.

Anonymous:1935:BTCa

[533] Anonymous. Brief table of contents of vol. XXII (1934–1935). Isis, 22 (2):620–622, February 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225144.

Anonymous:1935:BMa

[534] Anonymous. Back matter. Isis, 22(2):??, February 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225145.

Anonymous:1935:VIa

[535] Anonymous. Volume information. Isis, 23(1):1, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225216.

Anonymous:1935:FMb

[536] Anonymous. Front matter. Isis, 23(1):3–5, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225217.

Sarton:1935:PVXb

[537] George Sarton. Preface to volume XXIII of Isis (Quetelet). Isis, 23(1): 6–24, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225218.

Muller:1935:NMM

[538] Rheinhold F. G. Müller. Natur- und medizingeschichtliches aus dem Mahābhārata. Isis, 23(1):25–53, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225219.

Kapp:1935:AUKa

[539] A. G. Kapp. Arabische Übersetzer und Kommentatoren Euklids, sowie deren math.-naturwiss. Werke auf Grund des Ta’rīkh al-H .ukamā’ des Ibn al-Qift.ī II. (German) [Arabic translators and commentators of Euclid, and their mathematical-scientific works on the basis of Ta’rīkh al-H .ukamā’ des ibn al-Qift.ī II]. Isis, 23(1):54–99, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225220.

Meyerhof:1935:INX

[540] Max Meyerhof. Ibn An–Nafîs (XIIIth Cent.) and his theory of the lesser circulation. Isis, 23(1):100–120, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225221.

Karpinski:1935:ANA

[541] Louis C. Karpinski and Charles N. Staubach. An Anglo-Norman algorism of the Fourteenth Century. Isis, 23(1):121–152, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225222.

Sarton:1935:FED

[542] George Sarton. The first explanation of decimal fractions and measures (1585). Together with a history of the decimal idea and a facsimile (no. XVII) of Stevin’s Disme. Isis, 23(1):153–244, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225223.

Gandz:1935:AFD

[543] Solomon Gandz. Artificial fertilization of date-palms in Palestine and Arabia. Isis, 23(1):245–250, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225224.

Sarton:1935:AND

[544] George Sarton. Additional note on date culture in Ancient Babylonia. Isis, 23(1):251–252, June 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225225.

Bevenot:1935:NC

[545] Hugh Bévenot, Chauncey D. Leake, Edward Kremers, G. S., and Herbert M. Evans. Notes and correspondence. Isis, 23(1):253–259, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225226.

Kofoid:1935:BRBa

[546] Charles A. Kofoid. Book review: Aristotle. Fundamentals of the History of His Development by Werner Jaeger; Richard Robinson. Isis, 23(1): 260–261, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225227.

Kofoid:1935:BRBb

[547] Charles A. Kofoid. Book review: The Greek Herbal of Dioscorides by Robert T. Gunter; John Goodyer. Isis, 23(1):261–262, June 1935. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225228.

Gandz:1935:BRB

[548] Salomon Gandz. Book review: The Khand.akhadyaka, an Astronomical Treatise of Brahmagupta by P. Ch. Sengupta. Isis, 23(1):262–263, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225229.

Kofoid:1935:BRBc

[549] C. A. Kofoid. Book review: Buffon et la déscription de la nature by Louis Roule. Isis, 23(1):264–265, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225230.

Kofoid:1935:BRBd

[550] Charles A. Kofoid. Book review: Bernardin de Saint-Pierre et l’harmonie de la nature by Louis Roule. Isis, 23(1):265–266, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225231.

Kofoid:1935:BRBe

[551] C. A. Kofoid. Book review: Cuvier et la science de la nature by Louis Roule. Isis, 23(1):266–267, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225232.

Kofoid:1935:BRBf

[552] C. A. Kofoid. Book review: The Lamarck Manuscripts at Harvard by William Morton Wheeler; Thomas Barbour. Isis, 23(1):267–268, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225233.

Kofoid:1935:BRBg

[553] C. A. Kofoid. Book review: Un ami de Stendhal by Pierre Maes; Victor Jacquemont. Isis, 23(1):268, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225234.

Sarton:1935:BRBb

[554] George Sarton. Book review: New Light on the Most Ancient East. The Oriental Prelude to European Prehistory by V. Gordon Childe. Isis, 23 (1):269–271, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225235.

Sarton:1935:BRBc

[555] George Sarton. Book review: Children of the Yellow Earth. Studies in Prehistoric China by Johan Gunnar Andersson. Isis, 23(1):271–275, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225236.

Sarton:1935:BRBd

[556] George Sarton. Book review: Prominent Danish Scientists through the Ages by V. Meisen. Isis, 23(1):276–278, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225237.

Sarton:1935:BRBe

[557] George Sarton. Book review: Comment je vois le monde by Albert Einstein; Colonel Cros. Isis, 23(1):278–280, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225238.

Davis:1935:BRBb

[558] H. T. Davis. Book review: The Open World by Hermann Weyl; Mind and Nature. Isis, 23(1):281–284, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225239.

Malik:1935:BRBa

[559] Charles Mâlik. Book review: An Introduction to Logic and Scientific Method by Morris R. Cohen; Ernest Nagel. Isis, 23(1):284–287, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225240.

Kofoid:1935:BRBh

[560] Charles A. Kofoid. Book review: A History of Agriculture in the State of New York by Ulysses Prentiss Hedrick. Isis, 23(1):287–288, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225241.

Kofoid:1935:BRBi

[561] Charles A. Kofoid. Book review: History of Agriculture in the Southern United States to 1860 by Lewis Cecil Gray; Esther Katherine Thompson. Isis, 23(1):289, June 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225242.

Johnson:1935:BRB

[562] Francis R. Johnson. Book review: Tudor Geography, 1485–1583 by E. G. R. Taylor; Late Tudor and Early Stuart Geography, 1583–1650. Isis, 23(1):289–294, June 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225243.

Kofoid:1935:BRBj

[563] Charles A. Kofoid. Book review: The Making of Geography by R. E. Dickinson; O. J. R. Howarth. Isis, 23(1):294–295, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225244.

Merton:1935:BRBa

[564] Robert K. Merton. Book review: An Introduction to Pareto: His Sociology by George C. Homans; Charles P. Curtis. Isis, 23(1):295–296, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225245.

Malik:1935:BRBb

[565] Charles Mâlik. Book review: The Categories of Charles Peirce by Eugene Freeman. Isis, 23(1):296–297, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225246.

Kofoid:1935:BRBk

[566] Charles A. Kofoid. Book review: The Third Pandemic of Plague in Egypt by A. W. Wakil. Isis, 23(1):297–298, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225247.

Brasch:1935:PEA

[567] Frederick E. Brasch and Lao G. Simons. Proceedings of the eleventh annual meeting of the History of Science Society. Isis, 23(1):299–306, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225248.

S:1935:SBA

[568] G. S. Summary of Isis accounts for 1934. Isis, 23(1):307–308, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225249.

Anonymous:1935:BMb

[569] Anonymous. Back matter. Isis, 23(1):??, June 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225250.

Anonymous:1935:FMc

[570] Anonymous. Front matter. Isis, 23(2):309–458, September 1935. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224948.

Sarton:1935:SPV

[571] George Sarton. Second preface to volume XXIII: The history of science versus the history of medicine. Isis, 23(2):313–320, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224949.

Meyerhof:1935:TTC

[572] Max Meyerhof. Thirty-three clinical observations by rhazes (circa 900 a.d.). Isis, 23(2):321–372, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224950.

Stimson:1935:CIC

[573] Dorothy Stimson. Comenius and the invisible college. Isis, 23(2):373–388, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224951.

Viets:1935:SFH

[574] Henry R. Viets. Some features of the history of medicine in Massachusetts during the colonial period (1620–1770). Isis, 23(2):389–405, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224952.

Browne:1935:CIA

[575] C. A. Browne. The chemical industries of the American Aborigines. Isis, 23(2):406–424, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224953.

Shorr:1935:GPR

[576] Philip Shorr. The Genesis of prehistorical research. Isis, 23(2):425–443, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224954.

Sarton:1935:NCb

[577] George Sarton, Albert Tiberghien, and Edward Kremers. Notes and correspondence. Isis, 23(2):444–446, September 1935. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224955.

Sarton:1935:BRBf

[578] George Sarton. Book review: Das Giftbuch des ’Sanaq by Bettina Strauss. Isis, 23(2):447–448, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224956.

Sarton:1935:BRBg

[579] George Sarton. Book review: The Book of Instruction in the Elements of the Art of Astrology by Abu’l-Rayh.an Muh.ammad ibn Ah.mad al-Biruni; R. Ramsay Wright. Isis, 23(2):448–450, September 1935. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224957.

Sarton:1935:BRBh

[580] George Sarton. Book review: Acta et processus canonizacionis beate Birgitte by Isak Collijn; Acta et processus canonizationis Sanctae Birgittae; Iconographia birgittina typographica by Isak Collijn; Birgittinska gestalter by Isak Collijn; The Revelations of Saint Birgitta by William Patterson Cumming. Isis, 23(2):450–452, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224958.

Sarton:1935:BRBi

[581] George Sarton. Book review: Studien zu den astrologischen Schriften des Heinrich von Langenstein by Hubert Pruckner. Isis, 23(2):452–454, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224959.

Larkey:1935:BRBa

[582] S. V. Larkey. Book review: The Sinister Shepherd: A Translation of Girolamo Fracastoro’s Syphilidis sive de Morbo Gallico Libri Tres by William van Wyck. Isis, 23(2):454–455, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224960.

Larkey:1935:BRBb

[583] S. V. Larkey. Book review: Town Government in the Sixteenth Century by J. H. Thomas. Isis, 23(2):455–456, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224961.

Sarton:1935:BRBj

[584] George Sarton. Book review: Sir Isaac Newton’s Mathematical Principles of Natural Philosophy and His System of the World by Andrew Motte; Florian Cajori; Isaac Newton. Isis, 23(2):456–457, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224962.

Sarton:1935:BRBk

[585] George Sarton. Book review: Opera quaedam aut inedita aut obsoleta de rebus naturalibus by Emanuel Swedenborg; Itineraria by Emanuelis Swedenborgii; An Abridged Chronological List of the Works of Emanuel Swedenborg by Alfred H. Stroh; Greta Ekelöf. Isis, 23(2):459–463, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224963.

Sarton:1935:BRBl

[586] George Sarton. Book review: Les sciences de la nature en France, au XVIIIe siècle. Un chapitre de l’histoire des idées by Daniel Mornet; Les origines intellectuelles de la Révolution française, 1715–1787. Isis, 23 (2):463–467, September 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224964.

Sarton:1935:BRBm

[587] George Sarton. Book review: Storia del pensiero scientifico. Vol.I: Il mondo antico. by F. Enriques; G. de Santillana. Isis, 23(2):467–469, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224965.

Allen:1935:BRB

[588] James T. Allen. Book review: Essai sur la formation de la Pensée grecque by Pierre-Maxime Schuhl. Isis, 23(2):469–470, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224966.

S:1935:BRB

[589] G. S. Book review: The Building of the Roman Aqueducts by Esther Boise Van Deman. Isis, 23(2):470–471, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224967.

Sarton:1935:BRBn

[590] George Sarton. Book review: A History of Magic and Experimental Science. Volumes III and IV: Fourteenth and Fifteenth Centuries by Lynn Thorndike. Isis, 23(2):471–475, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224968.

Sarton:1935:BRBo

[591] George Sarton. Book review: The Transformation of Nature in Art by Ananda K. Coomaraswamy. Isis, 23(2):475–477, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224969.

Malik:1935:BRBc

[592] Charles Mâlik. Book review: Collected Papers of Charles Sanders Peirce by Charles Sanders Peirce; Charles Hartshorne; Paul Weiss. Isis, 23(2): 477–483, September 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224970.

Kofoid:1935:BRBl

[593] C. A. Kofoid. Book review: To the North! by Jeannette Mirsky. Isis, 23 (2):483–485, September 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224971.

Kofoid:1935:BRBm

[594] Charles A. Kofoid. Book review: Manchurian Plague Prevention Service by Wu Lien-Teh. Isis, 23(2):485–487, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224972.

Sarton:1935:FSC

[595] George Sarton and Frances Siegel. Forty-second critical bibliography of the history and philosophy of science and of the history of civilization (to end of November 1934, — with special reference to the Middle Ages and Byzantium). Isis, 23(2):488–627, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224973.

Anonymous:1935:BTCb

[596] Anonymous. Brief table of contents of vol. XXIII (1935). Isis, 23(2): 628–629, September 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224974.

Anonymous:1935:BMc

[597] Anonymous. Back matter. Isis, 23(2):??, September 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224975.

Anonymous:1935:VIb

[598] Anonymous. Volume information. Isis, 24(1):1, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224827.

Anonymous:1935:FMd

[599] Anonymous. Front matter. Isis, 24(1):3–6, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224828.

Sarton:1935:PVXc

[600] George Sarton. Preface to volume XXIV: Nuño Goncalves. Isis, 24(1):7– 14, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224830.

Welborn:1935:SMM

[601] Mary Catherine Welborn. Studies in Medieval metrology. Isis, 24(1):15– 36, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224831.

Kapp:1935:AUKb

[602] A. G. Kapp. Arabische Übersetzer und Kommentatoren Euklids, sowie deren math.-naturwiss. Werke auf Grund des Ta’rīkh al-H .ukamā’ des Ibn al-Qift.ī: III. (German) [Arabic translators and commentators of Euclid, and their mathematical-scientific works on the basis of Ta’rīkh al-H .ukamā’ des ibn al-Qift.ī III]. Isis, 24(1):37–79, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224832.

Gabrieli:1935:PSD

[603] Giuseppe Gabrieli. Per la storia della prima romana accademia dei Lincei. (Italian) [On the history of the first Roman Academy of the Lynxes]. Isis, 24(1):80–89, December 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224833.

Gafafer:1935:PLP

[604] William M. Gafafer. Peter Ludwig Panum’s “Observations on the Contagium of Measles”. Isis, 24(1):90–101, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224834.

Sarton:1935:NCc

[605] George Sarton, Henri Terrasse, H. P. J. Renaud, Ralph E. Ockenden, F. Szinnyei, Valeriu Bologa, O. Stein, Arnold C. Klebs, J. R. Partington, C. De Waard, and L. G. Notes and correspondence. Isis, 24(1):102–126, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224835.

Sarton:1935:BRBp

[606] George Sarton. Book review: Epiphanius de Gemmis by Robert Pierpont Blake; Henri De Vis. Isis, 24(1):127–129, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224836.

Sarton:1935:BRBq

[607] George Sarton. Book review: Medico-Mystical Work by Simon of T .aibutheh by Alphonse Mingana. Isis, 24(1):129–132, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224837.

Sarton:1935:BRBr

[608] George Sarton. Book review: Al-Farabi’s Arabic-Latin Writings on Music by Henry George Farmer. Isis, 24(1):132–134, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224838.

Kurland:1935:BRB

[609] Samuel Kurland. Book review: Super Oratione Dominica by Magistri Eckardi Opera Latina auspiciis Instituti Sanctae Sabinae ad codicum fidem edita; Raymundus Klibansky; Felicis Meiner. Isis, 24(1):134–136, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224839.

Nunn:1935:BRB

[610] George E. Nunn. Book review: Die Kolumbus-Originale by Fritz Streicher. Isis, 24(1):136–141, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224840.

Pelseneer:1935:BRBb

[611] J. Pelseneer. Book review: Isaac Newton. A Biography by Louis Trenchard More. Isis, 24(1):141–143, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224841.

Prall:1935:BRB

[612] D. W. Prall. Book review: Religion within the Limits of Reason Alone by Immanuel Kant; Theodore M. Greene; Hoyt H. Hudson. Isis, 24 (1):143–145, December 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224842.

Raisz:1935:BRB

[613] E. Raisz. Book review: Bering’s Voyages: An Account of the Efforts of the Russians to Determine the Relation of Asia and America by Frank Alfred Golder; E. P. Bertholf; G. W. Steller; F. A. Golder; Leonhard Stejneger. Isis, 24(1):146–149, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224843.

Davis:1935:BRBc

[614] Tenney L. Davis. Book review: Autobiographical Notes by Jöns Jacob Berzelius; Olof Larsell. Isis, 24(1):149–151, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224844.

Archibald:1935:BRB

[615] R. C. Archibald. Book review: Vorlesungen über Geschichte der antiken mathematischen Wissenschaften. Erster Band, Vorgriechische Mathematik by O. Neugebauer. Isis, 24(1):151–153, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224845.

Welborn:1935:BRB

[616] Mary Catherine Welborn. Book review: Medieval Latin Word-List from British and Irish Sources by J. H. Baxter; Charles Johnson. Isis, 24 (1):153–155, December 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224846.

Hummel:1935:BRB

[617] A. W. Hummel. Book review: Chung-kuo t’ien-chih shih (Agricultural History of China).–Volume I by Wan Kuo-ting. Isis, 24(1):155–157, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224847.

Sakanishi:1935:BRBa

[618] Shio Sakanishi. Book review: Chou pi suan ching no kenkyû (An Inquiry concerning the Chou pi suan ching) by Chûryô Noda. Isis, 24(1):157– 158, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224848.

Sakanishi:1935:BRBb

[619] Shio Sakanishi. Book review: Sai-i bu (Natural Calamities), Part 5: On Earthquakes by Ruijû Kokushi. Isis, 24(1):158–159, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224849.

Sakanishi:1935:BRBc

[620] Shio Sakanishi. Book review: Edo Jidai no Kagaku [Science during the Edo Period, 1600–1867] by Tokyo Science Museum. Isis, 24(1):159–160, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224850.

Sakanishi:1935:BRBd

[621] Shio Sakanishi. Book review: Nihon yakuen-shi no kenkyû [Historical Study of the Japanese Herb Gardens] by Uyeda, Sanpei. Isis, 24(1):160– 162, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224851.

Sarton:1935:BRBs

[622] George Sarton. Book review: Mémoires Scientifiques: Vol. XII: Comptes rendus et analyses. 1896–1904 by Paul Tannery; Mémoires Scientifiques: Vol. XIII: Correspondance by A. Diès. Isis, 24(1):162–163, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224852.

Sarton:1935:BRBt

[623] George Sarton. Book review: A History of Science, Technology, and Philosophy in the 16th & 17th Centuries by A. Wolf; F. Dannemann; A. Armitage. Isis, 24(1):164–167, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224853.

Merton:1935:BRBb

[624] Robert K. Merton. Book review: Cosmogonies of Our Fathers by Katherine Brownell Collier. Isis, 24(1):167–168, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224854.

Leonard:1935:BRB

[625] Henry S. Leonard. Book review: A System of Logistic by Willard Van Orman Quine. Isis, 24(1):168–172, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224855.

Sarton:1935:BRBu

[626] G. Sarton. Book review: Introduction to Modern Physics by Floyd Karker Richtmyer. Isis, 24(1):172–174, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224856.

Wilson:1935:BRB

[627] W. J. Wilson. Book review: Alchemy, Child of Greek Philosophy by Arthur John Hopkins. Isis, 24(1):174–177, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224857.

Merton:1935:BRBc

[628] Robert K. Merton. Book review: A History of Mechanical Inventions by Abbott Payson Usher. Isis, 24(1):177–180, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224858.

Sarton:1935:BRBv

[629] George Sarton. Book review: Geschichte der Cecidologie by Konrad Böhner. Isis, 24(1):180–183, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224859.

Kofoid:1935:BRBn

[630] C. A. Kofoid. Book review: The American Eagle, a Study in Natural and Civil History by Francis Hobart Herrick. Isis, 24(1):183–184, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224860.

Kofoid:1935:BRBo

[631] C. A. Kofoid. Book review: The Discoverers of the Fiji Islands, Tasman, Cook, Bligh, Wilson, Bellingshausen by G. C. Henderson. Isis, 24(1): 184–185, December 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224861.

Kofoid:1935:BRBp

[632] C. A. Kofoid. Book review: The Search for the Northwest Passage by Nellis M. Crouse. Isis, 24(1):185–186, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224862.

Kofoid:1935:BRBq

[633] C. A. Kofoid. Book review: The Drama of Weather by Napier Shaw. Isis, 24(1):186–188, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224863.

Merton:1935:BRBd

[634] Robert K. Merton. Book review: Science and Social Needs by Julian Huxley; H. Levy; Thomas D. Barlow; P. M. S. Blackett. Isis, 24(1): 188–189, December 1935. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224864.

Merton:1935:BRBe

[635] Robert K. Merton. Book review: Mind, Self and Society by George H. Mead; Charles W. Morris. Isis, 24(1):189–191, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224865.

Sarton:1935:FTC

[636] George Sarton and Frances Siegel. Forty-third critical bibliography of the history and philosophy of science and of the history of civilization (to end of February 1935, — with special reference to the Fourteenth Century and Asia). Isis, 24(1):192–303, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224866.

Anonymous:1935:BMd

[637] Anonymous. Back matter. Isis, 24(1):??, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224867.

Anonymous:1935:FDH

[638] Anonymous. Frontispiece: Dr. Harvey Cushing, surgeon, tenth President of the H.S.S. Isis, 24(1):??, December 1935. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224829.

Anonymous:1936:FMa

[639] Anonymous. Front matter. Isis, 24(2):305–307, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225288.

Sarton:1936:SPV

[640] George Sarton. Second preface to volume XXIV: Two fundamental problems. Isis, 24(2):308–309, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225289.

Ruska:1936:SMI

[641] Julius Ruska. Studien zu Muh.ammad ibn umail al-Tamīmī’s Kitāb al-Mā’ al-Waraqī wa’l–Ard. an–Najmīyah. Isis, 24(2):310–342, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225290.

Zoubov:1936:TAL

[642] V. Zoubov. Une théorie aristotélicienne de la lumière du XVIIe siècle. (French) [An Aristotlean theory of light from the 17th Century]. Isis, 24 (2):343–360, February 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225291.

vonHofsten:1936:CD

[643] N. von Hofsten. From Cuvier to Darwin. Isis, 24(2):361–366, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225292.

Nierenstein:1936:DLS

[644] M. Nierenstein. The discovery of lactic sugar. Isis, 24(2):367–369, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225293.

Viets:1936:SI

[645] Henry R. Viets. Stephen d’Irsay (1894–1934). Isis, 24(2):370–374, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225294.

Sarton:1936:MM

[646] George Sarton. Minoan mathematics. Isis, 24(2):375–381, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225295.

Thompson:1936:MMC

[647] Daniel V. Thompson, Jr. More Medieval color-making: Tractatus de Coloribus from Munich, Staatsbibliothek, MS. Latin 444. Isis, 24(2):382– 396, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225296.

Jones:1936:NCI

[648] Charles W. Jones. A note on concepts of the inferior planets in the early Middle Ages. Isis, 24(2):397–399, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225297.

Kinney:1936:DMM

[649] L. C. Mac Kinney. “Dynamidia” in Medieval medical literature. Isis, 24 (2):400–414, February 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225298.

Pearl:1936:HPJ

[650] Raymond Pearl. Historical papers in the journal “human biology”. Isis, 24(2):415–418, February 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225299.

Thorndike:1936:JTA

[651] L. Thorndike. John Tolhopf again. Isis, 24(2):419–421, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225300.

Anonymous:1936:CIWa

[652] Anonymous. Carnegie Institution of Washington. Division of Historical Research. Section of the History of Science. Isis, 24(2):422–428, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225301.

Pogo:1936:NC

[653] A. Pogo, George Sarton, Robert S. Woodbury, and Charles W. Morris. Notes and correspondence. Isis, 24(2):429–436, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225302.

Sarton:1936:BRBa

[654] George Sarton. Book review: A Bibliography of the Poem Syphilis sive morbus Gallicus by Girolamo Fracastoro of Verona by Leona Baumgartner; John F. Fulton. Isis, 24(2):437–439, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225303.

Rosen:1936:BRB

[655] Edward Rosen. Book review: Des Révolutions des Orbes Célestes by Nicolas Copernic; Alexandre Koyré. Isis, 24(2):439–442, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225304.

Merton:1936:BRBa

[656] Robert K. Merton. Book review: The Attitude of Voltaire to Magic and the Sciences by Margaret Sherwood Libby. Isis, 24(2):442–444, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225305.

Sarton:1936:BRBb

[657] George Sarton. Book review: Histoire des sciences. Antiquité by Pierre Brunet; Aldo Mieli. Isis, 24(2):444–447, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225306.

Sarton:1936:BRBc

[658] George Sarton. Book review: Extraits des principaux géographes arabes du moyen âge by R. Blachère. Isis, 24(2):447–449, February 1936. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225307.

Lenzen:1936:BRB

[659] V. F. Lenzen. Book review: The Rise of Modern Physics by Henry Crew. Isis, 24(2):449–450, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225308.

Usher:1936:BRB

[660] Abbott Payson Usher. Book review: Inventing the Ship by S. C. Gilfillan. Isis, 24(2):450–453, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225309.

Durand:1936:BRB

[661] Dana B. Durand. Book review: Quelques éclaircissements épars sur mes Monumenta Cartographica Africae et Aegypti by Prince Youssouf Kamal; E. J. Brill. Isis, 24(2):453–455, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225310.

Leake:1936:BRBa

[662] Chauncey D. Leake. Book review: The Medicine-Man of the American Indian and His Cultural Background by William Thomas Corlett. Isis, 24(2):455–456, February 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225311.

Homans:1936:BRB

[663] George C. Homans. Book review: The Mind and Society [Trattato di Sociologia generale, 1916] by Vilfredo Pareto; Arthur Livingston; Andrew Bongiorno; James Harvey Rogers. Isis, 24(2):456–467, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225312.

Sarton:1936:FFCa

[664] George Sarton, Kurt Vogel, Ralph E. Ockenden, and Frances Siegel. Forty-fourth critical bibliography of the history and philosophy of science and of the history of civilization (to end of May 1935, — with special reference to India). Isis, 24(2):468–536, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225313.

Anonymous:1936:VIa

[665] Anonymous. Volume information. Isis, 24(2):537–607, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225287.

Anonymous:1936:BTCa

[666] Anonymous. Brief table of contents of vol. XXIV (1935–36). Isis, 24 (2):608–610, February 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225314.

Anonymous:1936:BMa

[667] Anonymous. Back matter. Isis, 24(2):??, February 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225315.

Anonymous:1936:VIb

[668] Anonymous. Volume information. Isis, 25(1):1, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224976.

Anonymous:1936:FMb

[669] Anonymous. Front matter. Isis, 25(1):3–5, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224977.

Sarton:1936:PVX

[670] George Sarton. Preface to volume XXV: Humanism versus grammar. Isis, 25(1):6–8, May 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224978.

Macdonald:1936:MPA

[671] D. B. Macdonald. The meanings of the philosophers by al-Ghazzālī. Isis, 25(1):9–15, May 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224979.

Sarton:1936:IDF

[672] George Sarton and Solomon Gandz. The invention of the decimal fractions and the application of the exponential calculus by Immanuel Bonfils of Tarascon (c. 1350). Isis, 25(1):16–45, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224980.

Struik:1936:MND

[673] D. J. Struik. Mathematics in The Netherlands during the first half of the XVIth Century. Isis, 25(1):46–56, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224981.

Tannery:1936:LIP

[674] Paul Tannery. Une lettre inédite de Paul tannery à Frédéric hultsch. Isis, 25(1):57–59, May 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224982.

Meyerhof:1936:GB

[675] Max Meyerhof and G. S. Gotthelf Bergsträsser (1886–1933). Isis, 25(1): 60–62, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224983.

Brachwitz:1936:SMK

[676] Richard Brachwitz. Das Shoso-in Museum der Ksl. Schatz-kammer zu Nara und seine Heilmittel. (German) [The Shoso-in Museum of the Royal Treasure Chamber of Nara and its remedies]. Isis, 25(1):63–72, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224984.

Sarton:1936:CGB

[677] George Sarton. Chinese glass of the beginning of the Confucian Age. Isis, 25(1):73–79, May 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224985.

vonHofsten:1936:ICS

[678] N. von Hofsten. Ideas of creation and spontaneous generation prior to Darwin. Isis, 25(1):80–94, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224986.

Zirkle:1936:AIW

[679] Conway Zirkle. Animals impregnated by the wind. Isis, 25(1):95–130, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224987.

Sarton:1936:NC

[680] George Sarton, Otto Seydl, Robert S. Woodbury, Dana B. Durand, R. E. Ockenden, J. A. Vollgraff, Shio Sakanishi, and Bertha Margaret Frick. Notes and correspondence. Isis, 25(1):131–140, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224988.

Sarton:1936:BRBd

[681] George Sarton. Book review: Encyclopaedia of Philosophical and Natural Sciences as Taught in Baghdad about A.D. 817, or Book of Treasures by Job of Edessa by Job of Edessa; A. Mingana. Isis, 25(1):141–144, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224989.

Sarton:1936:BRBe

[682] George Sarton. Book review: Das Buch der Alaune und Salze by Julius Ruska. Isis, 25(1):144–145, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224990.

Gandz:1936:BRBa

[683] Solomon Gandz. Book review: Die Algebra des Abu Kamil Soga’ ben Aslam by Josef Weinberg. Isis, 25(1):145–147, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224991.

Sarton:1936:BRBf

[684] George Sarton. Book review: The Medical Man and the Witch during the Renaissance by Gregory Zilboorg. Isis, 25(1):147–152, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224992.

D:1936:BRBa

[685] T. L. D. Book review: Collected Works. Vols. VI-VII by Mikhail Vasilyevich Lomonosov. Isis, 25(1):153–154, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224993.

D:1936:BRBb

[686] T. L. D. Book review: A Scientist in the Early Republic, Samuel Latham Mitchell. 1764–1831 by Courtney Robert Hall. Isis, 25(1):154–155, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224994.

Thorndike:1936:BRBa

[687] Lynn Thorndike. Book review: La Pecia dans les manuscrits universitaires du XIIIe et du XIVe siècle by Jean Destrez. Isis, 25(1):155–157, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224995.

Sakanishi:1936:BRBa

[688] Shio Sakanishi. Book review: Shina chûsei igaku shi (History of Chinese Medicine in the Middle Ages) by Liao Wên–Jên. Isis, 25(1):158, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224996.

Gandz:1936:BRBb

[689] Solomon Gandz. Book review: The Romance of Hassidism by Jacob S. Minkin. Isis, 25(1):159–164, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224997.

Farber:1936:BRB

[690] Eduard Färber. Book review: Geschichte der physiologischen Chemie by Fritz Lieben. Isis, 25(1):164–166, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224998.

Merton:1936:BRBb

[691] Robert K. Merton. Book review: The Sociology of Invention by S. C. Gilfillan. Isis, 25(1):166–167, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224999.

Ashley-Montagu:1936:BRBa

[692] M. F. Ashley-Montagu. Book review: Woman. An Historical Gynaecological Compendium by Hermann Heinrich Ploss; Max Bartels; Paul Bartels; Eric John Dingwall; William Heinemann. Isis, 25(1):167–169, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225000.

Sarton:1936:BRBg

[693] George Sarton. Book review: A Documentary History of Primitivism and Related Ideas. Volume I: Primitivism and Related Ideas in Antiquity by Arthur O. Lovejoy; George Boas. Isis, 25(1):169–172, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225001.

Thorndike:1936:BRBb

[694] Lynn Thorndike. Book review: Education of the Founding Fathers of the Republic by James J. Walsh. Isis, 25(1):173–174, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225002.

Ashley-Montagu:1936:BRBb

[695] M. F. Ashley-Montagu. Book review: Classical Contributions to Obstetrics and Gynecology by Herbert Thoms. Isis, 25(1):174–175, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225003.

Sarton:1936:FFCb

[696] George Sarton and Frances Siegel. Forty-fifth critical bibliography of the history and philosophy of science and of the history of civilization (to December 1935). Isis, 25(1):176–317, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225004.

Anonymous:1936:BMb

[697] Anonymous. Back matter. Isis, 25(1):??, May 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225005.

Anonymous:1936:FMc

[698] Anonymous. Front matter. Isis, 25(2):319–322, September 1936. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225370.

Sarton:1936:HDR

[699] George Sarton. A Hindu decimal ruler of the third Millennium. Isis, 25 (2):323–326, September 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225371.

Davis:1936:DCH

[700] Tenney L. Davis. The dualistic cosmogony of Huai-nan-tzŭ and its relations to the background of Chinese and of European alchemy. Isis, 25 (2):327–340, September 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225372.

Ginzburg:1936:DJN

[701] Benjamin Ginzburg. Duhem and Jordanus Nemorarius. Isis, 25(2):341– 362, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225373.

Thorndike:1936:JVH

[702] Lynn Thorndike. Johann Virdung of Hassfurt again. Isis, 25(2):363–371, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225374.

Neave:1936:JBL

[703] E. W. J. Neave. Joseph Black’s lectures on the elements of chemistry. Isis, 25(2):372–390, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225375.

Menschutkin:1936:VVP

[704] Boris N. Menschutkin. Vasilij Vladimirovic Petrov and his physicochemical work. Isis, 25(2):391–398, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225376.

Sarton:1936:CSE

[705] George Sarton. On a curious subdivision of the Egyptian cubit. Isis, 25 (2):399–402, September 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225377.

Pogo:1936:EWC

[706] A. Pogo. Egyptian water clocks. Isis, 25(2):403–425, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225378.

Gandz:1936:BTR

[707] Solomon Gandz. The Babylonian tables of reciprocals. Isis, 25(2):426– 432, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225379.

Calvert:1936:DDS

[708] H. R. Calvert. Decimal division of scales before the metric system. Isis, 25 (2):433–436, September 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225380.

deBray:1936:VL

[709] M. E. J. Gheury de Bray. The velocity of light. Isis, 25(2):437–448, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225381.

Gandz:1936:NC

[710] Solomon Gandz, Henri Bernard, R. E. Ockenden, Kate C. H. Mead, Lynn Thorndike, George Sarton, Ralph C. Benedict, C. D. L., and Edmund O. von Lippmann. Notes and correspondence. Isis, 25(2):449–460, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225382.

Sarton:1936:BRBh

[711] George Sarton. Book review: An Essay on Landscape Painting by Shio Sakanishi; Kuo Hsi. Isis, 25(2):461–464, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225383.

Thorndike:1936:BRBc

[712] Lynn Thorndike. Book review: Die Quaestiones Naturales des Adelardus von Bath by Martin Müller. Isis, 25(2):464–465, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225384.

Johnson:1936:BRB

[713] Francis R. Johnson. Book review: Sir Walter Ralegh: The Last of the Elizabethans by Edward Thompson. Isis, 25(2):465–466, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225385.

Pelseneer:1936:BRB

[714] J. Pelseneer. Book review: The Diary of Robert Hooke (1672–1680) by Henry W. Robinson; Walter Adams; Early Science in Oxford, Vol. X, The Life and Work of Robert Hooke (Part IV) by R. T. Gunther. Isis, 25 (2):466–470, September 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225386.

Leake:1936:BRBb

[715] Chauncey D. Leake. Book review: Count Rumford of Massachusetts by James A. Thompson. Isis, 25(2):470–471, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225387.

Davis:1936:BRBa

[716] T. L. Davis. Book review: Une étape de la notion d’atome et de molécule; Gaudin by Marcel Delépine. Isis, 25(2):471–473, September 1936. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225388.

Gandz:1936:BRBc

[717] Solomon Gandz. Book review: Sternkunde und Sterndienst in Babel by Franz Xaver Kugler; Johann Schaumberger. Isis, 25(2):473–476, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225389.

Sarton:1936:BRBi

[718] George Sarton. Book review: The Library of A. Chester Beatty. Description of a Hieratic Papyrus with a Mythological Story, Love-Songs, and other Miscellaneous Texts by Alan H. Gardiner; Hieratic Papyri in the British Museum. Isis, 25(2):476–478, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225390.

Gandz:1936:BRBd

[719] Solomon Gandz. Book review: History of Hindu Mathematics by B. Datta; A. N. Singh. Isis, 25(2):478–488, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225391.

S:1936:BRB

[720] G. S. Book review: Annals of Science.–A Quarterly Review of the History of Science since the Renaissance by D. McKie; Harcourt Brown; H. W. Robinson. Isis, 25(2):488–489, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225392.

Karpinski:1936:BRB

[721] Louis C. Karpinski. Book review: Die Beziehungen der Mathematik zur Kultur der Gegenwart by A. Voss; Die Verbreitung mathematischen Wissens und mathematischer Auffassung by H. E. Timerding. Isis, 25 (2):489–491, September 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225393.

Davis:1936:BRBb

[722] H. T. Davis. Book review: Trend Analysis of Statistics by Max Sasuly. Isis, 25(2):491–493, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225394.

Davis:1936:BRBc

[723] H. T. Davis. Book review: The Principles of Quantum Mechanics by P. A. M. Dirac. Isis, 25(2):493–496, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225395.

Pogo:1936:BRBa

[724] Alexander Pogo. Book review: Eclipses of the Sun by S. A. Mitchell. Isis, 25(2):496–504, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225396.

Davis:1936:BRBd

[725] Tenney L. Davis. Book review: Origins and Development of Applied Chemistry by J. R. Partington. Isis, 25(2):504–507, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225397.

Sarton:1936:BRBj

[726] George Sarton. Book review: The Beginnings of Plant Hybridization by Conway Zirkle. Isis, 25(2):507–508, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225398.

Larsell:1936:BRBa

[727] O. Larsell. Book review: De Yverbornes Ö. Sextonhundratalsstudier by Johan Nordström. Isis, 25(2):508–510, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225399.

Sarton:1936:BRBk

[728] George Sarton. Book review: Guide Book to the Ruins of Quirigua by Sylvanus G. Morley. Isis, 25(2):511–513, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225400.

Sarton:1936:BRBl

[729] George Sarton. Book review: The Founding of Harvard College by Samuel Eliot Morison; Harvard College in the Seventeenth Century; The Development of Harvard University, 1869–1929. Isis, 25(2):513–520, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225401.

Mead:1936:BRB

[730] Kate C. Mead. Book review: Gyldne Laegeraad saerlight om Diaet by Johanne Christiansen. Isis, 25(2):520–521, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225402.

Sarton:1936:FSCa

[731] George Sarton and Frances Siegel. Forty-sixth critical bibliography of the history and philosophy of science and of the history of civilization (to end of February 1936, — with special reference to China and Japan). Isis, 25 (2):522–613, September 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225403.

Brasch:1936:HSS

[732] Frederick E. Brasch, Henry E. Sigerist, and Lao G. Simons. History of Science Society reports for 1935. Isis, 25(2):614–619, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225404.

Anonymous:1936:BR

[733] Anonymous. Isis report for 1935. Isis, 25(2):620, September 1936. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225405.

Anonymous:1936:BTCb

[734] Anonymous. Brief table of contents of vol. XXV (1936). Isis, 25(2): 621–623, September 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225406.

Anonymous:1936:BMc

[735] Anonymous. Back matter. Isis, 25(2):??, September 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225407.

Anonymous:1936:VIc

[736] Anonymous. Volume information. Isis, 26(1):1, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225044.

Anonymous:1936:FMd

[737] Anonymous. Front matter. Isis, 26(1):3–7, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225045.

Cumont:1936:RPA

[738] Franz Cumont. Regula Philippi Arrhidaei. Isis, 26(1):8–12, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225046.

Stamm:1936:BCT

[739] Edward Stamm. Tractatus de Continuo von Thomas Bradwardina. (German). Isis, 26(1):13–32, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225047.

Thorndike:1936:PMT

[740] Lynn Thorndike. Paravicus: A misprint, not a translator. Isis, 26(1):33– 36, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225048.

vonLippmann:1936:BAG

[741] Edmund O. von Lippmann and J. R. Partington. Biringuccio und Agricola. (German) [Biringuccio and Agricola]. Isis, 26(1):37–38, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225049.

Thorndike:1936:EPB

[742] Lynn Thorndike. Epitomes of Pliny’s Natural History in the Fifteenth Century. Isis, 26(1):39, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225050.

Ockenden:1936:MAD

[743] R. E. Ockenden. Marco Antonio de Dominis and his explanation of the rainbow. Isis, 26(1):40–49, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225051.

Calvert:1936:THL

[744] H. R. Calvert. Thompson of Hosier Lane, instrument makers of the Seventeenth Century. Isis, 26(1):50–52, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225052.

Sarton:1936:RCH

[745] George Sarton. Remarks concerning the history of Twentieth Century science. Isis, 26(1):53–62, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225053.

Archibald:1936:BM

[746] Raymond Clare Archibald. Babylonian mathematics. Isis, 26(1):63–81, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225054.

Gandz:1936:MMT

[747] Solomon Gandz. Mene Mene Tekel Upharsin, a chapter in Babylonian mathematics. Isis, 26(1):82–94, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225055.

Sarton:1936:TND

[748] George Sarton. Third note on date culture in Ancient Babylonia. Isis, 26(1):95–98, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225056.

Farber:1936:AGN

[749] Eduard Färber. Alte Gedanken und neue chemische theorien. Isis, 26 (1):99–126, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225057.

Stapleton:1936:FNA

[750] H. E. Stapleton. Further notes on the Arabic alchemical manuscripts in the libraries of India. Isis, 26(1):127–131, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225058.

Sarton:1936:NHA

[751] George Sarton. Notes on the history of anagrammatism. Isis, 26(1): 132–138, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225059.

Anonymous:1936:CIWb

[752] Anonymous. Carnegie Institution of Washington. Division of Historical Research. Section of the History of Science. Isis, 26(1):139–146, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225060.

Klibansky:1936:NC

[753] Raymond Klibansky, Alexander Pogo, Louis C. Karpinski, and George Sarton. Notes and correspondence. Isis, 26(1):147–156, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225061.

Ashley-Montagu:1936:BRBc

[754] M. F. Ashley-Montagu. Book review: The Historian Ephorus by G. L. Barber. Isis, 26(1):157–158, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225062.

Pogo:1936:BRBb

[755] A. Pogo. Book review: De re aedificatoria libri decem by Leon Battista Alberti; V. P. Zoubov; F. A. Petrovski. Isis, 26(1):158–159, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225063.

Ashley-Montagu:1936:BRBd

[756] M. F. Ashley-Montagu. Book review: The Library of Pico Della Mirandola by Pearl Kibre. Isis, 26(1):159–161, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225064.

Hartner:1936:BRBa

[757] Willy Hartner. Book review: Nogre stycker af thenn frije konst astronomia by Oluf Luth. Isis, 26(1):161–164, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225065.

Sarton:1936:BRBm

[758] George Sarton. Book review: Matteo Ricci’s Scientific Contribution to China by Henry Bernard; Edward Chalmers Werner. Isis, 26(1):164–167, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225066.

Sarton:1936:BRBn

[759] George Sarton. Book review: Euclides danicus, Amsterdam 1672 by Georg Mohr; Johannes Hjelmslev; Beiträge zur Lebensbeschreibung von Georg Mohr (1640–1697) by Johannes Hjelmslev. Isis, 26(1):167–170, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225067.

Merton:1936:BRBc

[760] Robert K. Merton. Book review: Ancients and Moderns: A Study of the Background of the Battle of the Books by Richard Foster Jones. Isis, 26 (1):171–172, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225068.

Larsell:1936:BRBb

[761] O. Larsell. Book review: Georg Stiernhielm, Filosofiska Fragment med Inledning och Kommentar. Vol. I by Johan Nordström. Isis, 26(1):172– 174, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225069.

Leake:1936:BRBc

[762] Chauncey D. Leake. Book review: An Analysis of the De Generatione Animalium of William Harvey by Arthur William Meyer. Isis, 26(1): 174–176, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225070.

Hartner:1936:BRBb

[763] Willy Hartner. Book review: Naturvetenskapens matematiska principer by Isaac Newton. Isis, 26(1):176–177, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225071.

Sarton:1936:BRBo

[764] George Sarton. Book review: Lychnos. Lärdomshistoriska Samfundets Årsbok by Johann Nordstrom; Anders Celsius, Professor i Uppsala, 1701–1744. Lychnos-Bibliotek, Studier och källskrifter utgivna av Lärdomshistoriska Samfundet by N. V. E. Nordenmark. Isis, 26(1):177– 180, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225072.

Weeks:1936:BRB

[765] Mary Elvira Weeks. Book review: Antoine Lavoisier. The Father of Modern Chemistry by Douglas McKie. Isis, 26(1):180–183, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225073.

Kofoid:1936:BRBa

[766] Charles A. Kofoid. Book review: Diary of the Voyage of H. M. S. Rattlesnake (1846–50) by Thomas Henry Huxley; Julian Huxley. Isis, 26 (1):183–185, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225074.

Kofoid:1936:BRBb

[767] C. A. Kofoid. Book review: J. Arthur Harris, Botanist and Biometrician by C. O. Rosendahl; R. A. Gortner; G. O. Burr. Isis, 26(1):185–186, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225075.

Ashley-Montagu:1936:BRBe

[768] M. F. Ashley-Montagu. Book review: Studies in Early Celtic Nature Poetry by Kenneth Jackson. Isis, 26(1):186–187, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225076.

K:1936:BRB

[769] A. C. K. Book review: Dictionnaire biographique des médecins en France au Moyen Age by Ernest Wickersheimer. Isis, 26(1):187–189, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225077.

Sakanishi:1936:BRBb

[770] Shio Sakanishi. Book review: Nihon môjin shi by Nakayama, Taro. Isis, 26(1):189–191, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225078.

Sakanishi:1936:BRBc

[771] Shio Sakanishi. Book review: Zô. Elephants by Sunamoto, Etsujiro. Isis, 26(1):191–192, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225079.

Sakanishi:1936:BRBd

[772] Shio Sakanishi. Book review: Honpô iki-bokora no kenkyû: iki-bokora no shijitsu to sono shinri bunseki. [Shinto Worship of Living Human Gods: Its History and Psychology] by Kato, Genchi. Isis, 26(1):192–193, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225080.

Sakanishi:1936:BRBe

[773] Shio Sakanishi. Book review: Shin kyû ryôhô taisei (Compendium of Moxa Healing) by Sasagawa, Tomooki. Isis, 26(1):193–196, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225081.

Sarton:1936:BRBp

[774] George Sarton. Book review: Tadataka Inô, the Japanese Land Surveyor by Ryôkichi Ôtani; Kazue Sugimura. Isis, 26(1):196–200, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225082.

Minkin:1936:BRB

[775] Jacob S. Minkin. Book review: Studien zum B’estschen H .asidismus by Ysander, Torsten. Isis, 26(1):200–202, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225083.

MacDonald:1936:BRB

[776] D. B. MacDonald. Book review: Grundzüge des islamischen Rechts by Gotthelf Bergsträsser; Joseph Schacht. Isis, 26(1):203, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225084.

Feuer:1936:BRB

[777] Lewis S. Feuer. Book review: Collected Papers of Charles Sanders Peirce (1839–1914).–Volume VI: Scientific Metaphysics by Charles Hartshorne; Paul Weiss. Isis, 26(1):203–208, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225085.

Ashley-Montagu:1936:BRBf

[778] M. F. Ashley-Montagu. Book review: The Tyranny of Greece over Germany by E. M. Butler. Isis, 26(1):208–210, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225086.

Fite:1936:BRB

[779] W. Benjamin Fite. Book review: The Collected Works of George Abram Miller.–Volume I. Isis, 26(1):210–211, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225087.

Struik:1936:BRB

[780] Ruth Ramler Struik. Book review: Zrozeni astronomie (Das Entstehen der Astronomie) by Arnost Dittrich. Isis, 26(1):211–212, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225088.

Sarton:1936:BRBq

[781] George Sarton. Book review: L’expérience barométrique, ses antécédents et ses explications. Étude historique by Cornelis de Waard. Isis, 26(1): 212–215, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225089.

Davis:1936:BRBe

[782] H. T. Davis. Book review: Relativity, Gravitation and World-Structure by E. A. Milne. Isis, 26(1):215–218, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225090.

Ashley-Montagu:1936:BRBg

[783] M. F. Ashley-Montagu. Book review: American Martyrs to Science through the Roentgen Rays by Percy Brown. Isis, 26(1):219–220, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225091.

Davis:1936:BRBf

[784] H. T. Davis. Book review: The Philosophy of a Biologist by John Scott Haldane. Isis, 26(1):220–223, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225092.

OSullivan:1936:BRB

[785] Paul M. O’Sullivan. Book review: Da fitologia sacra by Arlindo Camilo Monteiro. Isis, 26(1):223–224, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225093.

Hartner:1936:BRBc

[786] Willy Hartner. Book review: Geschichte der geodätischen Instrumente und Verfahren im Altertum und Mittelalter by Fritz Schmidt. Isis, 26 (1):224–228, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225094.

Ashley-Montagu:1936:BRBh

[787] M. F. Ashley-Montagu. Book review: Vitamins in Theory and Practice by Leslie J. Harris. Isis, 26(1):228–229, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225095.

Ashley-Montagu:1936:BRBi

[788] M. F. Ashley-Montagu. Book review: A Hundred Years of Anthropology by T. K. Penniman. Isis, 26(1):229–232, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225096.

Ashley-Montagu:1936:BRBj

[789] M. F. Ashley-Montagu. Book review: European Civilization. Its Origin and Development by Edward Eyre. Isis, 26(1):232–234, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225097.

Ashley-Montagu:1936:BRBk

[790] M. F. Ashley-Montagu. Book review: Primitive Law by A. S. Diamond. Isis, 26(1):234–237, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225098.

Merton:1936:BRBd

[791] Robert K. Merton. Book review: Religion and Learning by Olive M. Griffiths. Isis, 26(1):237–239, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225099.

Ashley-Montagu:1936:BRBl

[792] M. F. Ashley-Montagu. Book review: The Origins of Religion by Rafael Karsten. Isis, 26(1):239–243, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225100.

Sarton:1936:FSCb

[793] George Sarton and Frances Siegel. Forty-seventh critical bibliography of the history and philosophy of science and of the history of civilization (to end of April 1936, — with special reference to Israel). Isis, 26(1): 244–298, December 1936. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225101.

Anonymous:1936:BMd

[794] Anonymous. Back matter. Isis, 26(1):??, December 1936. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225102.

Anonymous:1937:FMa

[795] Anonymous. Front matter. Isis, 26(2):299–301, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224911.

Sarton:1937:APA

[796] George Sarton and M. P. Tannery. Appel pour l’achèvement du Système du monde de Duhem. Isis, 26(2):302–303, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224912.

Sarton:1937:HDR

[797] George Sarton. A Hindu decimal ruler of the third Millennium (second note). Isis, 26(2):304–305, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224913.

Thorndike:1937:DTE

[798] Lynn Thorndike. Date of the translation by Ermengaud Blasius of the work on the quadrant by Profatius Judaeus. Isis, 26(2):306–309, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224914.

Thorndike:1937:ICD

[799] Lynn Thorndike. Introduction and canon by Dalmatius to tables of Barcelona for the years 1361–1433 A.D. Isis, 26(2):310–320, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224915.

Thorndike:1937:FJV

[800] Lynn Thorndike. Faust and Johann Virdung of Hassfurt. Isis, 26(2): 321, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224916.

McColley:1937:EFC

[801] Grant McColley. An early friend of the Copernican theory: Gemma frisius. Isis, 26(2):322–325, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224917.

Pogo:1937:INC

[802] Alexander Pogo. Ioannes Nardius (ca. 1580–ca. 1655). Isis, 26(2):326– 329, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224918.

Yamamoto:1937:KHA

[803] Issei Yamamoto. Kunitomo and his astronomical activities in the premeizi era. Isis, 26(2):330–335, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224919.

Sarton:1937:DCT

[804] George Sarton. Darwin’s conception of the theory of natural selection. Isis, 26(2):336–340, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224920.

Cowles:1937:MDB

[805] Thomas Cowles. Malthus, Darwin, and Bagehot: A study in the transference of a concept. Isis, 26(2):341–348, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224921.

Rontgen:1937:DXR

[806] W. C. Rontgen and George Sarton. The discovery of x-rays. Isis, 26(2): 349–369, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224922.

Whitehill:1937:DEL

[807] Walter Muir Whitehill. The date of the earliest Latin-Arabic glossary. Isis, 26(2):370–372, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224923.

Raisz:1937:OHA

[808] Erwin Raisz. Outline of the history of American cartography. Isis, 26 (2):373–391, March 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224924.

McColley:1937:TDR

[809] Grant McColley. The theory of the diurnal rotation of the Earth. Isis, 26(2):392–402, March 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224925.

Sayili:1937:TM

[810] Aydin M. Sayili. Turkish medicine. Isis, 26(2):403–414, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224926.

Mead:1937:LM

[811] Kate C. H. Mead. Lappic medicine. Isis, 26(2):415–422, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224927.

Hume:1937:AML

[812] Edgar Erskine Hume. The Army medical library of Washington the largest medical library that has ever existed. Isis, 26(2):423–447, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224928.

Grossman:1937:NC

[813] Philip Grossman, Jean Capart, George Sarton, and Philip Shorr. Notes and correspondence. Isis, 26(2):448–455, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224929.

Pogo:1937:BRBa

[814] A. Pogo. Book review: Democritos. — Fragments and Testimonials of Antiquity by G. K. Bammel. Isis, 26(2):456–457, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224930.

Sarton:1937:BRBa

[815] George Sarton. Book review: The Ansab al-ashraf of al-Baladhuri by S. D. F. Goitein. Isis, 26(2):457–458, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224931.

Welborn:1937:BRBa

[816] Mary Catherine Welborn. Book review: Johannes de Mirfeld of St. Bartholomew’s Smithfield. His Life and Works by Hartley, Percival Horton-Smith; Aldridge, Harold Richard. Isis, 26(2):458–460, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224932.

Sarton:1937:BRBb

[817] George Sarton. Book review: Opicinus de Canistris by Richard Salomon; A. Heimann; R. Krautheimer. Isis, 26(2):460–463, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224933.

Pogo:1937:BRBb

[818] A. Pogo. Book review: Ephéméride illustrée de la vie d’Erasme by Daniel Van Damme. Isis, 26(2):463–464, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224934.

Pelseneer:1937:BRBa

[819] J. Pelseneer. Book review: Descartes. — Correspondance, publieé avec une introduction et des notes by Ch. Adam; G. Milhaud. Isis, 26(2):464– 465, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224935.

Sarton:1937:BRBc

[820] George Sarton. Book review: Bonaparte, gouverneur d’Égypte by François Charles-Roux. Isis, 26(2):465–470, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224936.

Sarton:1937:BRBd

[821] George Sarton. Book review: Comte. The Founder of Sociology by Francis Sydney Marvin; La vie d’Auguste Comte by Henri Gouhier; La jeunesse d’Auguste Comte et la formation du positivisme. I. Sous le signe de la liberté by Henri Gouhier. Isis, 26(2):470–477, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224937.

Sarton:1937:BRBe

[822] George Sarton. Book review: A Dictionary of Assyrian Chemistry and Geology by R. Campbell Thompson. Isis, 26(2):477–480, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224938.

Sakanishi:1937:BRBa

[823] Shio Sakanishi. Book review: Nihon shakai jigyô dai-nenpyô (Chronology of Social Service Works in Japan) by Keirin Taniyama. Isis, 26(2):480– 481, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224939.

Sakanishi:1937:BRBb

[824] Shio Sakanishi. Book review: Sûgaku-shi kenkyû by Kinnosuke Ogura. Isis, 26(2):481–482, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224940.

Sarton:1937:BRBf

[825] George Sarton. Book review: Catalogus codicum astrologorum graecorum by Mstislav Antonini F. Sangin. Isis, 26(2):483–484, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224941.

Durand:1937:BRB

[826] Dana B. Durand. Book review: De la marine antique à la marine moderne. La révolution du gouvernail, contribution à l’histoire de l’esclavage by Lefebvre des Noettes. Isis, 26(2):484–486, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224942.

P:1937:BRB

[827] A. P. Book review: History of Technics. Vol. I, Part I by B. L. Bogayevski. Isis, 26(2):486–487, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224943.

Sarton:1937:BRBg

[828] George Sarton. Book review: In the Shadow of To-Morrow. A Diagnosis of the Spiritual Distemper of Our Time by Jan Huizinga; J. H. Huizinga. Isis, 26(2):487–489, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224944.

Sarton:1937:FEC

[829] George Sarton and Frances Siegel. Forty-eighth critical bibliography of the history and philosophy of science and of the history of civilization (to end of September 1936 — with special reference to Iran and Islam). Isis, 26(2):490–604, March 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224945.

Anonymous:1937:BTCa

[830] Anonymous. Brief table of contents of vol. XXVI (1936). Isis, 26(2):605– 607, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224946.

Anonymous:1937:BMa

[831] Anonymous. Back matter. Isis, 26(2):??, March 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/224947.

Anonymous:1937:VIa

[832] Anonymous. Volume information. Isis, 27(1):1, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225012.

Anonymous:1937:FMb

[833] Anonymous. Front matter. Isis, 27(1):3–5, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225013.

Sarton:1937:MHH

[834] George Sarton. Materials for the history of the history of science. Isis, 27 (1):6–8, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225014.

Macdonald:1937:NMP

[835] D. B. Macdonald. Note on “the meanings of the philosophers by al-Ghazzālī”. Isis, 27(1):9–10, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225015.

Speter:1937:JCP

[836] Max Speter. Joh. Christ. Polykarp Erxleben, der erfolgreiche Bekämpfer des “Acidum pingue” — Interregnums in der Chemie (1775–76). (German) [Johannes Christian Polycarp Erxleben, the successful fighters of “acidum pingue” — the interregnum in chemistry (1775–76)]. Isis, 27(1): 11–19, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225016.

Sadi:1937:AHC

[837] Lutfi M. Sa’di, George Sarton, and W. T. Van Dyck. Al-Hakîm Cornelius Van Alen Van Dyck (1818–1895). Isis, 27(1):20–45, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225017.

Karpinski:1937:TPM

[838] L. C. Karpinski. Is there progress in mathematical discovery and did the Greeks have analytic geometry? Isis, 27(1):46–52, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225018.

Thorndike:1937:SH

[839] Lynn Thorndike. The secrets of Hermes. Isis, 27(1):53–62, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225019.

K:1937:NC

[840] A. C. K., Kate C. H. Mead, and C. Doris Hellman. Notes and correspondence. Isis, 27(1):63–67, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225020.

Kofoid:1937:BRBa

[841] Charles A. Kofoid. Book review: Die Entwicklung der biologischen Forschungsmethode in der Antike und ihre grundsätzliche Förderung durch Theophrast von Eresos by G. Senn. Isis, 27(1):68–69, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225021.

Ashley-Montagu:1937:BRBa

[842] M. F. Ashley-Montagu. Book review: Lucretius Poet & Philosopher by E. E. Sikes. Isis, 27(1):70–72, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225022.

Meyerhof:1937:BRB

[843] Max Meyerhof. Book review: Memorandum Book of a Tenth-Century Oculist by Casey A. Wood. Isis, 27(1):72–75, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225023.

Ashley-Montagu:1937:BRBb

[844] M. F. Ashley-Montagu. Book review: Georg Wilhelm Steller by Leonhard Stejneger. Isis, 27(1):75–77, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225024.

Kofoid:1937:BRBb

[845] Charles A. Kofoid. Book review: Lamarck et l’interprétation de la nature by Louis Roule. Isis, 27(1):77–78, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225025.

Kofoid:1937:BRBc

[846] Charles A. Kofoid. Book review: Collected Scientific Papers of Sir William Bate Hardy (1864–1934) by William Bate Hardy. Isis, 27(1): 78–79, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225026.

Davis:1937:BRBa

[847] Tenney L. Davis. Book review: Alchemy and Other Chemical Achievements of the Ancient Orient by Masumi Chikashige. Isis, 27(1):79–81, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225027.

Minkin:1937:BRB

[848] Jacob S. Minkin. Book review: The Hebrew Philosophical Genius by Duncan Black Macdonald. Isis, 27(1):81–86, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225028.

Ashley-Montagu:1937:BRBc

[849] M. F. Ashley-Montagu. Book review: Medicine in the Bible by Charles J. Brim. Isis, 27(1):86–89, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225029.

Kofoid:1937:BRBd

[850] Charles A. Kofoid. Book review: Studies in the Literature of Natural Science by Julian M. Drachman. Isis, 27(1):89–90, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225030.

Kofoid:1937:BRBe

[851] Charles A. Kofoid. Book review: Soviet Science by J. G. Crowther. Isis, 27(1):90–92, May 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225031.

Karpinski:1937:BRB

[852] Louis C. Karpinski. Book review: The Teaching of Arithmetic through 400 years, 1535–1935 by Florence A. Yeldham. Isis, 27(1):92–94, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225032.

Kofoid:1937:BRBf

[853] Charles A. Kofoid. Book review: Green Laurels. The Lives and Achievements of the Great Naturalists by Donald Culross Peattie. Isis, 27(1):95, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225033.

K:1937:BRB

[854] A. C. K. Book review: The Herbal of Pseudo-Apuleius by F. W. T. Hunger; Joh. Phil. de Lignamine. Isis, 27(1):96–98, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225034.

Kofoid:1937:BRBg

[855] Charles A. Kofoid. Book review: A History of Embryology by Joseph Needham. Isis, 27(1):98–102, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225035.

Ashley-Montagu:1937:BRBd

[856] M. F. Ashley-Montagu. Book review: Essays in Anthropology Presented to Alfred Louis Kroeber by Robert H. Lowie. Isis, 27(1):102–103, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225036.

Kofoid:1937:BRBh

[857] Charles A. Kofoid. Book review: Primitives and the Supernatural by Lucien Lévy–Bruhl; Lilian A. Clare. Isis, 27(1):104–105, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225037.

Friedrich:1937:BRB

[858] Carl Joachim Friedrich. Book review: Gesetz und Sittengesetz by Herbert Spiegelberg. Isis, 27(1):105–111, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225038.

Davis:1937:BRBb

[859] H. T. Davis. Book review: The Great Chain of Being by Arthur O. Lovejoy. Isis, 27(1):111–114, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225039.

Ashley-Montagu:1937:BRBe

[860] M. F. Ashley-Montagu. Book review: Through Science to God by Nathan A. Smyth. Isis, 27(1):114–116, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225040.

Sarton:1937:FNC

[861] George Sarton and Frances Siegel. Forty-ninth critical bibliography of the history and philosophy of science and of the history of civilization (to end of December 1936 — with special reference to sections 15 to 19). Isis, 27(1):117–200, May 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225041.

Brasch:1937:HSS

[862] Frederick E. Brasch and Lao G. Simons. History of Science Society reports for 1936. Isis, 27(1):201–205, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225042.

Anonymous:1937:BMb

[863] Anonymous. Back matter. Isis, 27(1):??, May 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225043.

Anonymous:1937:FMc

[864] Anonymous. Front matter. Isis, 27(2):207–210, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225408.

Sarton:1937:PVX

[865] George Sarton. Preface to volume XXVII: Unification of good will. Isis, 27(2):211–215, August 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225409.

Renaud:1937:DVD

[866] H. P. J. Renaud. Sur les dates de la vie du mathématicien arabe marocain Ibn al-Bannā’ (XIIe–XIIIe s. J.C.). (French) [On the dates of life of the Maroccan Arab mathematician Ibn al-Bannā’ (12th–13th Centuries A.D.)]. Isis, 27(2):216–218, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225410.

Thomson:1937:UTR

[867] S. Harrison Thomson. An unnoticed treatise of Roger Bacon on time and motion. Isis, 27(2):219–224, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225411.

Dienes:1937:EMH

[868] Mary Dienes. Eastern missions of the Hungarian Dominicans in the first half of the Thirteenth Century. Isis, 27(2):225–241, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225412.

Sarton:1937:RPI

[869] George Sarton. Rumphius, Plinius Indicus (1628–1702). Isis, 27(2):242– 257, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225413.

Lenzen:1937:NTL

[870] V. F. Lenzen. Newton’s third law of motion. Isis, 27(2):258–260, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225414.

McColley:1937:BJW

[871] Grant McColley. Bishop John Wilkins and the definition of science. Isis, 27(2):261–263, August 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225415.

Lowry:1937:CRY

[872] Lucia M. H. Lowry, C. D. Lowry, Jr., and John Rice Miner. Count Rumford, a youthful meteorologist. Isis, 27(2):264–267, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225416.

Lowry:1937:GFD

[873] Lucia M. H. Lowry, C. D. Lowry, Jr., and John Rice Miner. The graveyard fort; a disputed incident in the life of Count Rumford. Isis, 27 (2):268–285, August 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225417.

deSaussure:1937:LS

[874] Raymond de Saussure and Marthe de Saussure. Léopold de saussure (1866–1925). Isis, 27(2):286–305, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225418.

Hyde:1937:GSB

[875] Walter Woodhurn Hyde. The Great St. Bernard Pass and its Hospice. Isis, 27(2):306–320, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225419.

Bologa:1937:NC

[876] V. Bologa, L. van Hée, R. Toussaint, Valeriu L. Bologa, Herbert Noll-Husum, T. L. D., Arthur W. Hummel, and George Sarton. Notes and correspondence. Isis, 27(2):321–331, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225420.

S:1937:BRBa

[877] G. S. Book review: Treatise on Weights and Measures by Epiphanius; James Elmer Dean. Isis, 27(2):332–333, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225421.

Johnson:1937:BRB

[878] Francis R. Johnson. Book review: The Original Writings & Correspondence of the Two Richard Hakluyts by Richard Hakluyt; E. G. R. Taylor. Isis, 27(2):333–334, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225422.

Pelseneer:1937:BRBb

[879] Jean Pelseneer. Book review: Correspondance du P. Marin Mersenne by Mersenne, M.; Cornelis de Waard; René Pintard. Isis, 27(2):334– 336, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225423.

Pelseneer:1937:BRBc

[880] J. Pelseneer. Book review: Memoirs of Sir Isaac Newton’s Life, 1752 by William Stukeley; A. Hastings White. Isis, 27(2):336–337, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225424.

Hartner:1937:BRBa

[881] Willy Hartner. Book review: Papyri in the University of Michigan Collection by John Garret Winter. Isis, 27(2):337–339, August 1937.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225425.

Welborn:1937:BRBb

[882] Mary Catherine Welborn. Book review: The Materials of Medieval Painting by Daniel V. Thompson. Isis, 27(2):340–341, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225426.

Hartner:1937:BRBb

[883] Willy Hartner. Book review: History of Chinese Medicine by K. Chimin Wong; Wu Lien-teh. Isis, 27(2):341–342, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225427.

Sarton:1937:BRBh

[884] George Sarton. Book review: Discours et notices by Emile Picard. Isis, 27(2):342–344, August 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225428.

Hartner:1937:BRBc

[885] Willy Hartner. Book review: Dekane und Dekansternbilder by Wilhelm Gundel. Isis, 27(2):344–348, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225429.

Ashley-Montagu:1937:BRBf

[886] M. F. Ashley-Montagu. Book review: The Behaviour of Animals, an Introduction to Its Study by E. S. Russell. Isis, 27(2):348–349, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225430.

Ashley-Montagu:1937:BRBg

[887] M. F. Ashley-Montagu. Book review: Reports of the Cambridge Anthropological Expedition to the Torres Straits. Volume I by A. C. Haddon. Isis, 27(2):349–353, August 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225431.

Ashley-Montagu:1937:BRBh

[888] M. F. Ashley-Montagu. Book review: Frontier Folkways by James G. Leyburn. Isis, 27(2):353–354, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225432.

Ashley-Montagu:1937:BRBi

[889] M. F. Ashley-Montagu. Book review: Naven, a Survey of the Problems Suggested by a Composite Picture of the Culture of a New Guinea Tribe Drawn from Three Points of View by Gregory Bateson. Isis, 27(2):354– 357, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225433.

Ashley-Montagu:1937:BRBj

[890] M. F. Ashley-Montagu. Book review: A Political and Cultural History of Modern Europe by Carlton J. H. Hayes. Isis, 27(2):357–358, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225434.

Ashley-Montagu:1937:BRBk

[891] M. F. Ashley-Montagu. Book review: Philosophy and History by Raymond Klibansky; H. J. Paton. Isis, 27(2):358–360, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225435.

Sarton:1937:BRBi

[892] George Sarton. Book review: Histoire de la médecine vétérinaire by Emmanuel Leclainche. Isis, 27(2):360–363, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225436.

Sarton:1937:FCB

[893] George Sarton and Frances Siegel. Fiftieth critical bibliography of the history and philosophy of science and of the history of civilization (to end of February 1937–with special reference to mathematics). Isis, 27 (2):364–410, August 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225437.

Sarton:1937:BR

[894] George Sarton. Isis report for 1936. Isis, 27(2):411–412, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225438.

Anonymous:1937:BMc

[895] Anonymous. Back matter. Isis, 27(2):??, August 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225439.

Anonymous:1937:FMd

[896] Anonymous. Front matter. Isis, 27(3):413–415, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225147.

Sarton:1937:SPV

[897] George Sarton. Second preface to volume XXVII: Communion with Erasmus. Isis, 27(3):416–429, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225148.

Jones:1937:MBN

[898] Charles W. Jones. Manuscripts of Bede’s de natura rerum. Isis, 27 (3):430–440, November 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225149.

Diller:1937:GTG

[899] Aubrey Diller. A geographical treatise by Georgius Gemistus Pletho. Isis, 27(3):441–451, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225150.

Cole:1937:BJS

[900] F. J. Cole. The birthplace of Jan Swammerdam 1637–1680. Isis, 27(3): 452, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225151.

Shorr:1937:SJF

[901] Philip Shorr. Sir John Freind (1675–1728) M.D. pioneer historian of medicine. Isis, 27(3):453–474, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225152.

Sarton:1937:CFH

[902] G. Sarton. Charles Fremont historien de la technologie (1855–1930). Isis, 27(3):475–484, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225153.

Sampson:1937:JKF

[903] R. A. Sampson. John Knight Fotheringham (1874–1936). Isis, 27(3): 485–492, November 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225154.

Merton:1937:SK

[904] Robert K. Merton. The sociology of knowledge. Isis, 27(3):493–503, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225155.

Davis:1937:BRBc

[905] Tenney L. Davis. Book review: Tao, the Great Luminant by Evan Morgan. Isis, 27(3):504–506, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225156.

Sarton:1937:BRBj

[906] George Sarton. Book review: Al-Razi’s Buch Geheimnis der Geheimnisse by Julius Ruska. Isis, 27(3):506–507, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225157.

Sakanishi:1937:BRBc

[907] Shio Sakanishi. Book review: Seki-ryû sanpô shichibu sho (Seven Mathematical Treatises) by Seki, Takakazu. Isis, 27(3):507–508, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225158.

Raisz:1937:BRB

[908] Erwin Raisz. Book review: Peary by William Herbert Hobbs. Isis, 27 (3):508–512, November 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225159.

Diller:1937:BRB

[909] Aubrey Diller. Book review: Science in Antiquity by Benjamin Farrington. Isis, 27(3):512–513, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225160.

S:1937:BRBb

[910] S. S. Book review: An Introduction to Zen Buddhism by Suzuki, Daisetz Teitaro. Isis, 27(3):514, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225161.

S:1937:BRBc

[911] S. S. Book review: The Training of the Zen Buddhist Monk by Suzuki, Daisetz Teitaro. Isis, 27(3):514–515, November 1937. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225162.

Davis:1937:BRBd

[912] Tenney L. Davis. Book review: Life and Experiences of a Bengali Chemist. Vol. II by Prafulla Chandra Rây. Isis, 27(3):515–516, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225163.

Bridenbaugh:1937:BRB

[913] Carl Bridenbaugh. Book review: Scientific Interests in the Old South by Thomas Cary Johnson. Isis, 27(3):517–519, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225164.

Davis:1937:BRBe

[914] Tenney L. Davis. Book review: De la méthode dans les sciences expérimentales by Henry Le Chatelier. Isis, 27(3):519–522, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225165.

Hellman:1937:BRB

[915] C. Doris Hellman. Book review: Scritti, conferenze, discorsi sulla storia delle matematchei, raccolti per iniziativa e pubblicati sotto gli auspici della sezione ligure della societá “Mathesis” by Gino Loria. Isis, 27(3): 522–524, November 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225166.

Sakanishi:1937:BRBd

[916] Shio Sakanishi. Book review: Ka Etsu Nô ni okeru sûgaku (Mathematics in the Kaga feudal clan) by Tanaka, Tetsukichi. Isis, 27(3):524–525, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225167.

Sakanishi:1937:BRBe

[917] Shio Sakanishi. Book review: Nihon sûgaku shi (History of Japanese Mathematics) by Mizuki, Kozue. Isis, 27(3):525–526, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225168.

Sakanishi:1937:BRBf

[918] Shio Sakanishi and George Sarton. Book review: Sûgaku-shi by Jacques Boyer; Tsuruichi Hayashi. Isis, 27(3):526–527, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225169.

Lenzen:1937:BRB

[919] V. F. Lenzen. Book review: Mechanics, Molecular Physics, Heat, and Sound by Robert Andrews Millikan; Duane Roller; Earnest Charles Watson. Isis, 27(3):527–528, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225170.

Davis:1937:BRBf

[920] Tenney L. Davis. Book review: Prelude to Chemistry by John Read. Isis, 27(3):528–531, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225171.

Zirkle:1937:BRB

[921] Conway Zirkle. Book review: Herbarius, recherches sur le cérémonial usité chez les anciens pour la cueillette des simples et des plantes magiques by Armand Delatte. Isis, 27(3):531–532, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225172.

Hartner:1937:BRBd

[922] Willy Hartner. Book review: De store opdagelser by Ørjan Olsen. Isis, 27 (3):532–534, November 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225173.

Ashley-Montagu:1937:BRBl

[923] M. F. Ashley-Montagu. Book review: Man Makes Himself by V. Gordon Childe. Isis, 27(3):534–536, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225174.

Ashley-Montagu:1937:BRBm

[924] M. F. Ashley-Montagu. Book review: Storia della medicina by Arturo Castiglioni. Isis, 27(3):536–538, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225175.

Leake:1937:BRB

[925] C. D. Leake. Book review: The Development of Modern Medicine: An Interpretation of the Social and Scientific Factors Involved by Richard H. Shryock. Isis, 27(3):538–539, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225176.

Anonymous:1937:VIb

[926] Anonymous. Volume information. Isis, 27(3):540–600, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225146.

Anonymous:1937:BTCb

[927] Anonymous. Brief table of contents of vol. XXVII (1937). Isis, 27(3): 601–604, November 1937. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225177.

Anonymous:1937:BMd

[928] Anonymous. Back matter. Isis, 27(3):??, November 1937. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225178.

Anonymous:1938:VIa

[929] Anonymous. Volume information. Isis, 28(1):1, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225799.

Anonymous:1938:FMa

[930] Anonymous. Front matter. Isis, 28(1):3–6, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225800.

Sarton:1938:IHS

[931] George Sarton. An institute for the history of science and civilization (third article). Isis, 28(1):7–17, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225801.

Mullett:1938:SWP

[932] Charles F. Mullett. Sir William Petty on the plague. Isis, 28(1):18–25, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225802.

Harris:1938:MBH

[933] Henry Harris. Manchester’s board of health in 1796. Isis, 28(1):26–37, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225803.

Schlueter:1938:FAW

[934] Robert E. Schlueter. Frederick Adolphus Wislizenus (1810–1889), pioneer meteorologist, physician and natural scientist. Isis, 28(1):38–52, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225804.

Thorndike:1938:CHH

[935] Lynn Thorndike. Charles Homer Haskins. Isis, 28(1):53–56, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225805.

Hammett:1938:IAH

[936] Frederik S. Hammett. The ideas of the Ancient Hindus concerning Man. Isis, 28(1):57–72, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225806.

Davis:1938:PRA

[937] Tenney L. Davis. Pictorial representations of alchemical theory. Isis, 28(1):73–86, February 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225807.

Anonymous:1938:CIW

[938] Anonymous. Carnegie Institution of Washington. Division of Historical Research. Section of the History of Science. Isis, 28(1):87–91, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225808.

Zoubov:1938:NC

[939] V. Zoubov, S. Günaltay, Süheyl Ünver, Grant McColley, and G. S. Notes and correspondence. Isis, 28(1):92–100, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225809.

Ashley-Montagu:1938:BRBa

[940] M. F. Ashley-Montagu. Book review: Who Gave the World Syphilis? The Haitian Myth by Richard C. Holcomb. Isis, 28(1):101–102, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225810.

Ashley-Montagu:1938:BRBb

[941] M. F. Ashley-Montagu. Book review: Von der Fürtrefflichkeit und Nutz der Anatomy by Wilhelm Fabry von Hilden; F. de Quervain; H. Bloesch. Isis, 28(1):102–103, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225811.

Needham:1938:BRB

[942] Joseph Needham. Book review: Religious Motives in the Medical Biology of the Seventeenth Century by Walter Pagel. Isis, 28(1):103–105, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225812.

Pelseneer:1938:BRBa

[943] Jean Pelseneer. Book review: David Gregory, Isaac Newton and Their Circle. Extracts from David Gregory’s Memoranda, 1677–1708 by W. G. Hiscock. Isis, 28(1):105–106, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225813.

P:1938:BRBa

[944] A. P. Book review: The Works of M. V. Lomonosov on Physics and Chemistry by B. N. Menshutkin. Isis, 28(1):106–109, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225814.

Ulich:1938:BRB

[945] Robert Ulich. Book review: Noah Webster, Pioneer of Learning by Ervin C. Shoemaker. Isis, 28(1):110–112, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225815.

Sarton:1938:BRBa

[946] George Sarton. Book review: Réaumur [1683–1757]. Un esprit encyclopedique en dehors de l’Encyclopedie. D’après des documents inedits by Jean Torlais. Isis, 28(1):112–115, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225816.

Kofoid:1938:BRBa

[947] Charles A. Kofoid. Book review: Sir Joseph Banks. His Relations with Australia by George Mackaness. Isis, 28(1):116, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225817.

Kofoid:1938:BRBb

[948] Charles A. Kofoid. Book review: Travels and Scientific Collections of Thomas Nuttall by Francis W. Pennell. Isis, 28(1):116–118, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225818.

Sarton:1938:BRBb

[949] George Sarton. Book review: Recollections of My Life by Santiago Ramón y Cajal; E. Horne Craigie; Juan Cano. Isis, 28(1):118–122, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225819.

Kofoid:1938:BRBc

[950] Charles A. Kofoid. Book review: Snow on Cholera by W. H. Frost. Isis, 28(1):122–123, February 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225820.

Diller:1938:BRBa

[951] Aubrey Diller. Book review: Die Frauenheilkunde der alten Welt by Paul Diepgen. Isis, 28(1):123–126, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225821.

Temkin:1938:BRB

[952] Owsei Temkin. Book review: The Papyrus Ebers by B. Ebbell. Isis, 28 (1):126–131, February 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225822.

P:1938:BRBb

[953] A. P. Book review: Agriculture in Western Mediaeval Records by N. A. Dobiash-Rozhdestvenskaya; M. I. Burski. Isis, 28(1):131–134, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225823.

Ashley-Montagu:1938:BRBc

[954] M. F. Ashley-Montagu. Book review: Early Science in Cambridge by R. T. Gunther. Isis, 28(1):134, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225824.

Sarton:1938:BRBc

[955] George Sarton. Book review: IIIe Congrès international d’histoire des sciences. Isis, 28(1):135–138, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225825.

Ashley-Montagu:1938:BRBd

[956] M. F. Ashley-Montagu and H. T. Davis. Book review: Mathematics for the Million by Lancelot Hogben. Isis, 28(1):138–140, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225826.

Pogo:1938:BRBa

[957] A. Pogo. Book review: Le nombre d’or by W. E. van Wijk; De Gregoriaansche kalender by W. E. van Wijk. Isis, 28(1):140–142, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225827.

Durand:1938:BRBa

[958] Dana B. Durand. Book review: L’oeuvre scientifique du Club Alpin Français, (1874–1922) by Club Alpin Français. Isis, 28(1):143–144, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225828.

Durand:1938:BRBb

[959] Dana B. Durand. Book review: Hallucinations scientifiques (Les portulans) by Youssouf Kamal. Isis, 28(1):144–146, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225829.

Kofoid:1938:BRBd

[960] Charles A. Kofoid. Book review: Color Changes of Animals in Relation to Nervous Activity by G. H. Parker. Isis, 28(1):146–149, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225830.

Kofoid:1938:BRBe

[961] Charles A. Kofoid. Book review: The Social History of American Agriculture by Joseph Schafer. Isis, 28(1):149–150, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225831.

Sarton:1938:BRBd

[962] George Sarton. Book review: Africa View by Julian Huxley. Isis, 28 (1):150–151, February 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225832.

Merton:1938:BRBa

[963] Robert K. Merton. Book review: Factors Determining Human Behaviour; Authority and the Individual. Isis, 28(1):151–154, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225833.

Sarton:1938:FFCa

[964] George Sarton and Frances Siegel. Fifty-first critical bibliography of the history and philosophy of science and of the history of civilization (to end of July 1937–with special reference to mechanics, astronomy and physics). Isis, 28(1):155–316, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225834.

Anonymous:1938:BMa

[965] Anonymous. Back matter. Isis, 28(1):??, February 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225835.

Anonymous:1938:FMb

[966] Anonymous. Front matter. Isis, 28(2):317–320, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225689.

Sarton:1938:PVTa

[967] George Sarton. Preface to volume twenty-eight: A story from the Arabian nights. Isis, 28(2):321–329, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225690.

Gohory:1938:LTN

[968] Jacques Gohory and W. H. Bowen. L’histoire de la terre neuve du Peru. Isis, 28(2):330–340, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225691.

Richeson:1938:NRK

[969] A. W. Richeson. Notes on a Rechenbüchlein of Konrad Glaser. Isis, 28(2):341–348, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225692.

Bowen:1938:ETT

[970] Willis Herbert Bowen. The earliest treatise on tobacco: Jacques Gohory’s “instruction sur l’herbe petum”. Isis, 28(2):349–363, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225693.

McColley:1938:JBA

[971] Grant McColley. Josephus Blancanus and the adoption of our word “telescope”. Isis, 28(2):364–365, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225694.

Loria:1938:NSC

[972] Gino Loria. Nel secondo centenario della nascita di G. L. Lagrange 1736– 1936. (Italian) [On the second centennial of the birth of G. L. Lagrange 1736–1936]. Isis, 28(2):366–375, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225695.

Ervine:1938:JTW

[973] St John Ervine. John Turtle Wood, discoverer of the Artemision 1869. Isis, 28(2):376–384, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225696.

Sadi:1938:LWG

[974] Lutfi M. Sa’di and George Sarton. The life and works of George Edward Post (1838–1909). Isis, 28(2):385–417, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225697.

Sarton:1938:ABH

[975] George Sarton. Arthur berry, historian of astronomy (1862–1929). Isis, 28(2):418–420, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225698.

Carpenter:1938:WMW

[976] Frank M. Carpenter. William Morton Wheeler. Isis, 28(2):421–423, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225699.

Hopkins:1938:DEA

[977] Arthur John Hopkins. A defence of Egyptian alchemy. Isis, 28(2):424– 431, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225700.

Meyerhof:1938:MJP

[978] Max Meyerhof. Mediaeval Jewish physicians in the Near East, from Arabic sources. Isis, 28(2):432–460, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225701.

Sarton:1938:NCa

[979] George Sarton, W. Burke-Gaffney, M. Nierenstein, Henry E. Sigerist, R. J. Forbes, and F. S. Marvin. Notes and correspondence. Isis, 28(2): 461–466, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225702.

Ashley-Montagu:1938:BRBe

[980] M. F. Ashley-Montagu. Book review: Andreae Vesalii bruxellensis Icones anatomicae by Andreas Vesalius bruxellensis. Isis, 28(2):467–469, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225703.

Pagel:1938:BRBa

[981] Walter Pagel. Book review: Theophrastus Paracelsus. Idee und Problem seiner Weltanschauung by Franz Strunz. Isis, 28(2):469–471, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225704.

Pagel:1938:BRBb

[982] Walter Pagel. Book review: Paracelsus en Hahnemann, een Renaissance der Geneeskunst by Oosterhuis, R. A. B. Isis, 28(2):471–473, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225705.

Ashley-Montagu:1938:BRBf

[983] M. F. Ashley-Montagu. Book review: John Donne and the New Philosophy by Charles Monroe Coffin. Isis, 28(2):473–475, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225706.

Prall:1938:BRB

[984] D. W. Prall. Book review: Bishop Butler and the Age of Reason: A Study in the History of Thought by Ernest Campbell Mossner. Isis, 28(2):475– 477, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225707.

Ashley-Montagu:1938:BRBg

[985] M. F. Ashley-Montagu. Book review: Edward Gibbon, 1737–1794 by D. M. Low. Isis, 28(2):477–478, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225708.

Ashley-Montagu:1938:BRBh

[986] M. F. Ashley-Montagu. Book review: The Diary of a Surgeon in the Year 1751–1752, by John Knyveton by Ernest Gray. Isis, 28(2):478–480, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225709.

Sarton:1938:BRBe

[987] George Sarton. Book review: Madame Curie (1867–1934). A Biography by Eve Curie; Vincent Sheean; Pierre Curie (1859–1906) by Marie Curie; Charlotte Kellogg; Vernon Kellogg; Marie Skłodowska-Curie, 1867–1934 by Claudius Regaud. Isis, 28(2):480–484, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225710.

Ashley-Montagu:1938:BRBi

[988] M. F. Ashley-Montagu. Book review: Mea Culpa & The Life and Work of Semmelweis by Louis-Ferdinand Celine; Robert Allerton Parker. Isis, 28(2):484–487, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225711.

Ashley-Montagu:1938:BRBj

[989] M. F. Ashley-Montagu. Book review: Recollections and Reflections by J. J. Thomson. Isis, 28(2):487–488, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225712.

Ashley-Montagu:1938:BRBk

[990] M. F. Ashley-Montagu. Book review: Pavlov and His School by Y. P. Frolov. Isis, 28(2):488–489, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225713.

Ashley-Montagu:1938:BRBl

[991] M. F. Ashley-Montagu. Book review: Out of My Life and Work by August Forel. Isis, 28(2):489–490, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225714.

Archibald:1938:BRB

[992] R. C. Archibald. Book review: Mathematische Keilschrift-Texte by O. Neugebauer. Isis, 28(2):490–491, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225715.

Diller:1938:BRBb

[993] Aubrey Diller. Book review: The Frame of the Ancient Greek Maps by William Arthur Heidel. Isis, 28(2):491–493, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225716.

Greene:1938:BRB

[994] Henry Copley Greene. Book review: La musique du Moyen âge by Théophile Gérold; Histoire de la musique des origines à la fin du XIVe siècle. Isis, 28(2):494–502, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225717.

Macdonald:1938:BRB

[995] D. B. Macdonald. Book review: History of the Arabs by Philip K. Hitti. Isis, 28(2):503–504, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225718.

Sakanishi:1938:BRBa

[996] Shio Sakanishi. Book review: I-ka sentetsu shôzô shû (A Collection of Portraits of the Confucian Philosophers Who Practiced Medicine) by Fujinami Gôichi. Isis, 28(2):504–506, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225719.

Ashley-Montagu:1938:BRBm

[997] M. F. Ashley-Montagu. Book review: Men of Science by J. G. Crowther; Famous American Men of Science by J. G. Crowther. Isis, 28(2):507– 508, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225720.

Ashley-Montagu:1938:BRBn

[998] M. F. Ashley-Montagu. Book review: Human Affairs by R. B. Cattell; J. Cohen; R. M. W. Travers. Isis, 28(2):508–510, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225721.

Sarton:1938:BRBf

[999] George Sarton. Book review: Men of Mathematics by Eric Temple Bell. Isis, 28(2):510–513, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225722.

Hellman:1938:BRB

[1000] C. Doris Hellman. Book review: Comets. Their Nature, Origin, and Place in the Science of Astronomy by Mary Proctor; A. C. D. Crommelin. Isis, 28(2):513–514, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225723.

McColley:1938:BRB

[1001] Grant McColley. Book review: Astronomical Thought in Renaissance England by Francis R. Johnson. Isis, 28(2):514–516, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225724.

deSantillana:1938:BRB

[1002] G. de Santillana. Book review: L’elettrologia fine al Volta by Mario Gliozzi. Isis, 28(2):516–520, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225725.

Sarton:1938:BRBg

[1003] George Sarton. Book review: A Meteorological Chronology to A.D. 1450 by C. E. Britton. Isis, 28(2):520–522, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225726.

Ashley-Montagu:1938:BRBo

[1004] M. F. Ashley-Montagu. Book review: Mining and Metallurgy in Negro Africa by Walter Cline. Isis, 28(2):522–528, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225727.

Ashley-Montagu:1938:BRBp

[1005] M. F. Ashley-Montagu. Book review: The Metallurgy and Technology of Gold and Platinum among the Pre-Columbian Indians by Paul Bergsøe. Isis, 28(2):529–531, May 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225728.

Ashley-Montagu:1938:BRBq

[1006] M. F. Ashley-Montagu. Book review: New Frontiers of the Mind by J. B. Rhine. Isis, 28(2):531–534, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225729.

Ashley-Montagu:1938:BRBr

[1007] M. F. Ashley-Montagu. Book review: Archeology of St. Lawrence Island, Alaska by Henry B. Collins. Isis, 28(2):534–536, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225730.

Ashley-Montagu:1938:BRBs

[1008] M. F. Ashley-Montagu. Book review: Witchcraft, Oracles and Magic among the Azande by E. E. Evans-Pritchard. Isis, 28(2):536, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225731.

Ogden:1938:BRB

[1009] Henry V. S. Ogden. Book review: The Great Chain of Being by Arthur O. Lovejoy. Isis, 28(2):537–540, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225732.

Sarton:1938:FSC

[1010] George Sarton and Frances Siegel. Fifty-second critical bibliography of the history and philosophy of science and of the history of civilization (to end of October 1937–with special reference to chemistry, technology and the biological sciences). Isis, 28(2):541–616, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225733.

Anonymous:1938:BTCa

[1011] Anonymous. Brief table of contents of vol. XXVIII (1938). Isis, 28(2): 617–619, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225734.

Anonymous:1938:BMb

[1012] Anonymous. Back matter. Isis, 28(2):620, May 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225735.

Anonymous:1938:VIb

[1013] Anonymous. Volume information. Isis, 29(1):1, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225917.

Anonymous:1938:FMc

[1014] Anonymous. Front matter. Isis, 29(1):3–8, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225918.

Sarton:1938:PVTb

[1015] George Sarton. Preface to volume twenty-nine: … primum herbam deinde spicam deinde plenum frumentum in spica. (Latin) [First the blade, then the ear, after that the full corn in the ear]. Isis, 29(1): 9–14, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225919.

Sarton:1938:TOK

[1016] George Sarton. The tomb of omar Khayyâm. Isis, 29(1):15–19, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225920.

Saltykow:1938:SCG

[1017] N. Saltykow. Souvenirs concernant le Géomètre Yougoslave Marinus Ghetaldi, conservés à Doubrovnik, en Dalmatie. (French) [memoirs concerning the Yugoslav geometer, Marinus Ghetaldi, kept in Dubrovnik in Dalmatia]. Isis, 29(1):20–23, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225921.

Pelseneer:1938:DTP

[1018] Jean Pelseneer. Descartes et le théorème de Poincaré. Isis, 29(1):24– 28, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225922.

Cornog:1938:SSG

[1019] William H. Cornog. Sir Samuel Garth — a court physician of the 18th Century. Isis, 29(1):29–42, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225923.

DuPre:1938:FTN

[1020] A. Mason Du Pré, Jr. The first table of the normal probability integral: Its use by Kramp, who constructed it. Isis, 29(1):43–48, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225924.

Weeks:1938:NGS

[1021] Mary Elvira Weeks. Nils Gabriel Sefström–The sesquicentennial of his birth. Isis, 29(1):49–57, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225925.

Pogo:1938:WJK

[1022] A. Pogo. The writings of J. K. Fotheringham. Isis, 29(1):58–68, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225926.

Thorndike:1938:AAJ

[1023] Lynn Thorndike. Astronomy according to the Jews. Isis, 29(1):69–71, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225927.

Hartner:1938:SNC

[1024] Willy Hartner. Some notes on Chinese musical art. Isis, 29(1):72–94, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225928.

Pelseneer:1938:PS

[1025] Jean Pelseneer. Le primitif et le savant. Isis, 29(1):95–97, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225929.

Sarton:1938:NCb

[1026] George Sarton and H. P. J. Renaud. Notes and correspondence. Isis, 29 (1):98–103, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225930.

Sarton:1938:BRBh

[1027] George Sarton. Book review: The Study of Human Abilities. The Jen wu chih of Liu Shao by John K. Shryock. Isis, 29(1):104–108, July 1938.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225931.

Sarton:1938:BRBi

[1028] George Sarton. Book review: The Medico-Philosophical Controversy between Ibn Butlan of Baghdad and Ibn Ridwan of Cairo. A Contribution to the History of Greek Learning among the Arabs by Joseph Schacht; Max Meyerhof. Isis, 29(1):108–110, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225932.

H:1938:BRB

[1029] A. H. Book review: En islandsk scholasticus fra det 12. århundre by Anne Holtsmark. Isis, 29(1):110, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225933.

Johnson:1938:BRBa

[1030] Francis R. Johnson. Book review: Procedures and Metaphysics: A Study in the Philosophy of Mathematical-Physical Science in the Sixteenth and Seventeenth Centuries by Edward W. Strong. Isis, 29(1):110–113, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225934.

Johnson:1938:BRBb

[1031] Francis R. Johnson. Book review: The School of Night: A Study of the Literary Relationships of Sir Walter Ralegh by M. C. Bradbrook. Isis, 29(1):113–115, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225935.

Ashley-Montagu:1938:BRBt

[1032] M. F. Ashley-Montagu. Book review: Rim of Christendom by Herbert Eugene Bolton. Isis, 29(1):116, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225936.

Watson:1938:BRB

[1033] E. C. Watson. Book review: The Physical Treatises of Pascal: The Equilibrium of Liquids and the Weight of the Mass of the Air by Blaise Pascal; I. H. B. Spiers; A. G. H. Spiers; Pascal, the Life of Genius by Morris Bishop. Isis, 29(1):116–118, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225937.

Ashley-Montagu:1938:BRBu

[1034] M. F. Ashley-Montagu. Book review: Sir Kenelm Digby by John F. Fulton. Isis, 29(1):118–119, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225938.

Merton:1938:BRBb

[1035] Robert K. Merton. Book review: Science and Social Welfare in the Age of Newton by G. N. Clark. Isis, 29(1):119–121, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225939.

Saunders:1938:BRBa

[1036] John B. de C. M. Saunders. Book review: The History of the Surgeons’ Company, 1745–1800 by Cecil Wall. Isis, 29(1):121–123, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225940.

Shorr:1938:BRB

[1037] Philip Shorr. Book review: Jouannet, grand-père de la préhistoire by André Cheynier. Isis, 29(1):123–126, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225941.

Ashley-Montagu:1938:BRBv

[1038] M. F. Ashley-Montagu. Book review: Autobiography of Isaac Jones Wistar, 1827–1905 by Isaac Jones Wistar. Isis, 29(1):126–127, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225942.

Kofoid:1938:BRBf

[1039] Charles A. Kofoid. Book review: Charles Kingsley, 1819–1875 by Margaret Farrand Thorp. Isis, 29(1):127–129, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225943.

Kofoid:1938:BRBg

[1040] Charles A. Kofoid. Book review: Naturalists of the Frontier by Samuel Wood Geiser. Isis, 29(1):129–132, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225944.

Norman:1938:BRB

[1041] Daniel Norman. Book review: Einstein’s Theory of Relativity by Abraham I. Katsh. Isis, 29(1):132–133, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225945.

Kofoid:1938:BRBh

[1042] Charles A. Kofoid. Book review: Memorial Volume by G. Carl Huber. Isis, 29(1):133–134, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225946.

Ashley-Montagu:1938:BRBw

[1043] M. F. Ashley-Montagu. Book review: Short Years. The Life and Letters of John Bruce MacCullum, M. D. by Archibald Malloch. Isis, 29(1): 134–135, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225947.

Sarton:1938:BRBj

[1044] George Sarton. Book review: A Glossary of Greek Birds by D’Arcy Wentworth Thompson. Isis, 29(1):135–138, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225948.

Welborn:1938:BRBa

[1045] Mary Catherine Welborn. Book review: Early Medieval Medicine with Special Reference to France and Chartres by Loren C. MacKinney. Isis, 29(1):138–140, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225949.

Welborn:1938:BRBb

[1046] Mary Catherine Welborn. Book review: A Catalogue of Incipits of Mediaeval Scientific Writings in Latin by Lynn Thorndike; Pearl Kibre. Isis, 29(1):140–141, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225950.

Ashley-Montagu:1938:BRBx

[1047] M. F. Ashley-Montagu. Book review: Mediaeval English Pavingtiles by Loyd Haberly. Isis, 29(1):141–142, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225951.

Thorndike:1938:BRB

[1048] Lynn Thorndike. Book review: Medieval and Historiographical Studies in Honor of James Westfall Thompson by James Lea Cate; Eugene N. Anderson. Isis, 29(1):142–143, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225952.

Sarton:1938:BRBk

[1049] George Sarton. Book review: The Unveiling of Arabia by Reginald Hugh Kiernan; The Syrian Desert by Christina Phelps Grant. Isis, 29(1):143– 145, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225953.

Gandz:1938:BRBa

[1050] Solomon Gandz. Book review: A History of Jewish Literature by Meyer Waxman. Isis, 29(1):145–148, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225954.

Gandz:1938:BRBb

[1051] Solomon Gandz. Book review: A Social and Religious History of the Jews by Salo W. Baron. Isis, 29(1):148–150, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225955.

Gandz:1938:BRBc

[1052] Solomon Gandz. Book review: Judaism and Christianity by W. O. E. Oesterley; H. Loewe. Isis, 29(1):150–153, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225956.

Zirkle:1938:BRBa

[1053] Conway Zirkle. Book review: Food America Gave the World by A. Hyatt Verrill; Otis W. Barrett. Isis, 29(1):153–155, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225957.

Hocking:1938:BRB

[1054] Richard Hocking. Book review: Aspects of Science by Tobias Dantzig. Isis, 29(1):155–157, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225958.

Sarton:1938:BRBl

[1055] George Sarton. Book review: Mémoires scientifiques publiés par J. L. Heiberg & H. G. Zeuthen. XIV, Correspondance by Paul Tannery; A. Diès. Isis, 29(1):157–158, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225959.

Merton:1938:BRBc

[1056] Robert K. Merton. Book review: The Scientist in Action by William H. George. Isis, 29(1):159, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225960.

Ginzburg:1938:BRB

[1057] Benjamin Ginzburg. Book review: Through Science to Philosophy by Herbert Dingle. Isis, 29(1):160–163, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225961.

Leonard:1938:BRB

[1058] Henry S. Leonard. Book review: The Logical Syntax of Language by Rudolf Carnap; Amethe Smeaton. Isis, 29(1):163–167, July 1938. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225962.

Gandz:1938:BRBd

[1059] Solomon Gandz. Book review: Geschichte der Elementar-Mathematik by Johannes Tropfke. Isis, 29(1):167–169, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225963.

Pogo:1938:BRBb

[1060] A. Pogo. Book review: The Romance of the Calendar by P. W. Wilson. Isis, 29(1):169–172, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225964.

Lenzen:1938:BRB

[1061] V. F. Lenzen. Book review: Science and Music by James Jeans. Isis, 29 (1):172, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225965.

Weeks:1938:BRBa

[1062] Mary Elvira Weeks. Book review: Historia de la física by Paul F. Schurmann. Isis, 29(1):172–176, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225966.

Weeks:1938:BRBb

[1063] Mary Elvira Weeks. Book review: A Hundred Years of Chemistry by Alexander Findlay. Isis, 29(1):176–179, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225967.

Davis:1938:BRBa

[1064] Tenney L. Davis. Book review: A Short History of Chemistry by J. R. Partington. Isis, 29(1):179–181, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225968.

Zirkle:1938:BRBb

[1065] Conway Zirkle. Book review: Yearbook of Agriculture, 1936, 1937. Isis, 29(1):182–183, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225969.

Kofoid:1938:BRBi

[1066] Charles A. Kofoid. Book review: Fragments of Entomological History, including Some Personal Recollections of Men and Events by Herbert Osborn. Isis, 29(1):183, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225970.

Z:1938:BRB

[1067] C. Z. Book review: A History of Agricultural Experimentation and Research in the United States, 1607–1925 by Alfred Charles True. Isis, 29 (1):183–184, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225971.

Speter:1938:BRBa

[1068] Max Speter. Book review: Die Zuckererzeugung, 1600–1850 by Jakob Baxa. Isis, 29(1):184–185, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225972.

Zirkle:1938:BRBc

[1069] Conway Zirkle. Book review: Botanic Gardens of the World, Material for a History by C. Stuart Gager. Isis, 29(1):185–186, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225973.

Sarton:1938:BRBm

[1070] George Sarton. Book review: Tobacco by Jerome E. Brooks. Isis, 29(1): 186–188, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225974.

Durand:1938:BRBc

[1071] Dana B. Durand. Book review: Terrae incognitae by Richard Hennig. Isis, 29(1):188–189, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225975.

Sarton:1938:BRBn

[1072] George Sarton. Book review: Historia studii linguae arabicae et collegii missionum ordinis fratrum minorum in conventu ad S. Petrum in Monte Aureo Romae erecti by Arduinus Kleinhans. Isis, 29(1):189–191, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225976.

Ashley-Montagu:1938:BRBy

[1073] M. F. Ashley-Montagu. Book review: Aftermath, a Supplement to the Golden Bough by James George Frazer. Isis, 29(1):191–192, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225977.

Zirkle:1938:BRBd

[1074] Conway Zirkle. Book review: Coming into Being among the Australian Aborigines by Ashley-Montagu, M. F. Isis, 29(1):192–194, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225978.

Kofoid:1938:BRBj

[1075] Charles A. Kofoid. Book review: The Wandering Spirit by Ragnar Numelin. Isis, 29(1):194–196, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225979.

Ashley-Montagu:1938:BRBz

[1076] M. F. Ashley-Montagu. Book review: The Growth of Literature. Volume I The Ancient Literature of Europe by H. Munro Chadwick; N. Kershaw Chadwick. Isis, 29(1):196–197, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225980.

Campbell:1938:BRB

[1077] C. Macfie Campbell. Book review: The Mentally Ill in America. A History of Their Care and Treatment from Colonial Times by Albert Deutsch. Isis, 29(1):197–200, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225981.

Saunders:1938:BRBb

[1078] John B. de C. M. Saunders. Book review: American Medicine by Henry E. Sigerist; Hildegard Wagel; Socialized Medicine in the Soviet Union by Henry E. Sigerist. Isis, 29(1):200–203, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225982.

Sarton:1938:FTC

[1079] George Sarton and Frances Siegel. Fifty-third critical bibliography of the history and philosophy of science and of the history of civilization (to end of January 1938–with special reference to sections 31 to 33). Isis, 29(1): 204–300, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225983.

Leake:1938:IMH

[1080] Chauncey D. Leake. Indianapolis meeting of hss, 1937. Isis, 29(1):301– 302, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225984.

Brasch:1938:HSS

[1081] F. E. Brasch and George Sarton. The History of Science Society. Isis, 29(1):303–305, July 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225985.

Anonymous:1938:BMc

[1082] Anonymous. Back matter. Isis, 29(1):306, July 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225986.

Anonymous:1938:FMd

[1083] Anonymous. Front matter. Isis, 29(2):307–310, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225479.

Sarton:1938:DPA

[1084] George Sarton. Deuxième préface au volume vingt neuf l’histoire de la science et l’organisation internationale. Isis, 29(2):311–325, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225480.

Hopkins:1938:SKP

[1085] Arthur John Hopkins. A study of the kerotakis process as given by Zosimus and later alchemical writers. Isis, 29(2):326–354, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225481.

Kilgour:1938:PJW

[1086] Frederick G. Kilgour. Professor John Winthrop’s notes on Sun spot observations (1739). Isis, 29(2):355–361, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225482.

Weil:1938:SBN

[1087] E. Weil. Some bibliographical notes on the first publication on the Roentgen rays. Isis, 29(2):362–365, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225483.

deSantillana:1938:IGG

[1088] George de Santillana. The idealism of Giovanni Gentile. Isis, 29(2): 366–376, November 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225484.

Thorndike:1938:SAC

[1089] Lynn Thorndike. A study in the analysis of complex scientific manuscripts Sloane 3457: An important alchemical manuscript. Isis, 29 (2):377–392, November 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225485.

Kenney-Herbert:1938:EAP

[1090] A. Kenney-Herbert. The Elephantine Aramaic papyri. Isis, 29(2):393– 397, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225486.

Farber:1938:PC

[1091] Eduard Färber. A philosophy of catalysis. Isis, 29(2):398–402, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225487.

Sarton:1938:NCc

[1092] George Sarton, J. Pelseneer, Lynn Thorndike, Ettore Bortolotti, and P. Sergescu. Notes and correspondence. Isis, 29(2):403–416, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225488.

Gandz:1938:BRBe

[1093] Solomon Gandz and G. S. Book review: Die Sphaerik von Menelaos aus Alexandrien in der Verbesserung von Abu Nasr Mansur b.’Ali b.’Iraq by Max Krause. Isis, 29(2):417–423, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225489.

Sarton:1938:BRBo

[1094] George Sarton. Book review: The Elements of Theology by Proclus; Eric Robinson Dodds. Isis, 29(2):423–428, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225490.

Woodbury:1938:BRB

[1095] Robert S. Woodbury. Book review: The Mind of Latin Christendom by Edward Motley Pickman. Isis, 29(2):428–430, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225491.

S:1938:BRBa

[1096] G. S. Book review: Le livre des questions sur l’oeil de H .onain Ibn Ish.âq by Paul Sbath; Max Meyerhof. Isis, 29(2):430–431, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225492.

Sarton:1938:BRBp

[1097] George Sarton. Book review: OEuvres complètes de Christiaan Huygens by Christiaan Huygens. Isis, 29(2):431–433, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225493.

Pelseneer:1938:BRBb

[1098] Jean Pelseneer. Book review: Hevelius, Flamsteed and Halley. Three Contemporary Astronomers and Their Mutual Relations by Eugene Fairfield MacPike. Isis, 29(2):433–434, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/225494.

Davis:1938:BRBb

[1099] Tenney L. Davis. Book review: Traité Élémentaire de Chimie by Lavoisier. Isis, 29(2):434–435, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225495.

Ashley-Montagu:1938:BRBaa

[1100] M. F. Ashley-Montagu. Book review: Diderot Interpreter of Nature by Jonathan Kemp. Isis, 29(2):435–436, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225496.

Davis:1938:BRBc

[1101] Tenney L. Davis. Book review: Leçons de Philosophie Chimique by J. B. Dumas. Isis, 29(2):436–437, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225497.

Sarton:1938:BRBq

[1102] George Sarton. Book review: Figures de savants by Alfred Lacroix. Isis, 29(2):437–440, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225498.

Ashley-Montagu:1938:BRBab

[1103] M. F. Ashley-Montagu. Book review: Claude Bernard, Physiologist by J. M. D. Olmsted. Isis, 29(2):440–441, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225499.

Speter:1938:BRBb

[1104] Max Speter. Book review: August Kekulé by Richard Anschütz. Isis, 29 (2):442–443, November 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225500.

Kofoid:1938:BRBk

[1105] Charles A. Kofoid. Book review: Baron Constantin von Economo (1876– 1931). His Life and Work by Baroness Constantin von Economo; J. von Wagner-Jauregg; Ramsay Spillman. Isis, 29(2):443–444, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225501.

Gandz:1938:BRBf

[1106] Solomon Gandz. Book review: The Legends of the Jews by Louis Ginzberg; Boaz Cohen. Isis, 29(2):445–449, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225502.

S:1938:BRBb

[1107] S. S. Book review: Tôyô kakke-byô kenkyû. (Historical Study of the Beri-Beri in the Orient) by Liao Wen-jen. Isis, 29(2):449–450, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225503.

S:1938:BRBc

[1108] S. S. Book review: Nihon sugaku no tokushu-sei (Unique Features of Japanese Mathematics) Chuo Koron by Kinnosuke Ogura. Isis, 29(2): 450–451, November 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225504.

Sakanishi:1938:BRBb

[1109] Shio Sakanishi. Book review: Hôken sûgaku no metsubô (Decline of the Japanese Native Mathematics of the Feudal Period) Kaizo by Kinnosuke Ogura. Isis, 29(2):451–452, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225505.

Sakanishi:1938:BRBc

[1110] Shio Sakanishi. Book review: Nihon ganka-gaku shôshi. (A Brief History of Ophthalmology in Japan) by Kensaburo Ogawa. Isis, 29(2):452–453, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225506.

Pagel:1938:BRBc

[1111] Walter Pagel. Book review: Denker der italienischen Renaissance. Gestalten und Probleme by R. Hoenigswald. Isis, 29(2):453–457, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225507.

Sarton:1938:BRBr

[1112] George Sarton. Book review: Adventures in Error by Vilhjalmur Stefansson. Isis, 29(2):457–459, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225508.

Kofoid:1938:BRBl

[1113] Charles A. Kofoid. Book review: Trail-Blazers of Science by Martin Gumpert; Edwin L. Shuman. Isis, 29(2):460–461, November 1938. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225509.

Davis:1938:BRBd

[1114] H. T. Davis. Book review: The Logical Structure of Science by A. Cornelius Benjamin. Isis, 29(2):461–464, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225510.

Davis:1938:BRBe

[1115] H. T. Davis. Book review: An Introduction to the Philosophy of Science by A. Cornelius Benjamin. Isis, 29(2):464–469, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225511.

Kofoid:1938:BRBm

[1116] Charles A. Kofoid. Book review: Josiah Macy, Jr. Foundation by Ludwig Kast. Isis, 29(2):469–470, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225512.

Pogo:1938:BRBc

[1117] A. Pogo. Book review: El calendario maya en la cultura de Tiahuanacu by Fritz Buck. Isis, 29(2):470–472, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225513.

Kofoid:1938:BRBn

[1118] Charles A. Kofoid. Book review: Recent Advances in Cytology by C. D. Darlington. Isis, 29(2):472–474, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225514.

Kofoid:1938:BRBo

[1119] Charles A. Kofoid. Book review: Ecological Animal Geography by Richard Hesse; W. C. Allee; Karl P. Schmidt. Isis, 29(2):474–475, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225515.

Ashley-Montagu:1938:BRBac

[1120] M. F. Ashley-Montagu. Book review: The History of Ethnological Theory by Robert H. Lowie. Isis, 29(2):475–477, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225516.

Saunders:1938:BRBc

[1121] John B. de C. M. Saunders. Book review: A History of Women in Medicine from the Earliest Times to the Beginning of the Nineteenth Century by Kate Campbell Hurd-Mead. Isis, 29(2):477–479, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225517.

Sarton:1938:FFCb

[1122] George Sarton and Frances Siegel. Fifty-fourth critical bibliography of the history and philosophy of science and of the history of civilization (to end of April 1938–with special reference to sections 34 to 49 inclusive). Isis, 29(2):480–601, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225518.

Anonymous:1938:BTCb

[1123] Anonymous. Brief table of contents of vol. XXIX (1938). Isis, 29(2): 602–604, November 1938. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225519.

Anonymous:1938:BMd

[1124] Anonymous. Back matter. Isis, 29(2):??, November 1938. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225520.

Anonymous:1939:VIa

[1125] Anonymous. Volume information. Isis, 30(1):1, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225575.

Anonymous:1939:FMa

[1126] Anonymous. Front matter. Isis, 30(1):3–5, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225576.

Sarton:1939:PVX

[1127] George Sarton. Preface to volume XXX: Ugly ducklings. Isis, 30(1):6– 16, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225577.

al-Dinal-Suyuti:1939:TEP

[1128] Jalal al-Din al Suyut.i and Leon Nemoy. The treatise on the Egyptian pyramids (Tuh.fat al-kirām fī khabar al-ahrām). Isis, 30(1):17–37, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225578.

Roth:1939:EAP

[1129] George B. Roth. An early American pharmacologist: Horatio C. Wood (1841–1920). Isis, 30(1):38–45, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225579.

Speter:1939:PD

[1130] Max Speter. Paul dorveaux (1851–1938). Isis, 30(1):46–51, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225580.

Miller:1939:PSH

[1131] Hugh Miller. Philosophy of science and history of science. Isis, 30(1):52– 64, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225581.

Sayili:1939:AER

[1132] Aydin M. Sayili. The Aristotelian explanation of the rainbow. Isis, 30 (1):65–83, February 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225582.

Vanhee:1939:ECM

[1133] L. Vanhee. Euclide en chinois et mandchou. (French) [Euclid in Chinese and Manchu]. Isis, 30(1):84–88, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225583.

Anonymous:1939:CIW

[1134] Anonymous. Carnegie Institution of Washington. Division of Historical Research. Section of the History of Science. Isis, 30(1):89–94, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225584.

Nierenstein:1939:NC

[1135] M. Nierenstein, Lewis B. Beatty, George Sarton, S. C. Gilfillan, and Albert R. Goldsmith. Notes and correspondence. Isis, 30(1):95–99, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225585.

Sarton:1939:BRBa

[1136] George Sarton. Book review: Hayy ben Yaqdhân, roman philosophique d’Ibn Thofaïl by Léon Gauthier. Isis, 30(1):100–103, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225586.

Durand:1939:BRBa

[1137] Dana B. Durand. Book review: Marco Polo: The Description of the World. Vol. II by A. C. Moule; Paul Pelliot. Isis, 30(1):103–109, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225587.

Pogo:1939:BRBa

[1138] A. Pogo. Book review: Leben und Wirken des Johannes Müller von Königsberg genannt Regiomontanus by Ernst Zinner. Isis, 30(1):109– 111, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225588.

Pogo:1939:BRBb

[1139] A. Pogo. Book review: Der d.eutsche Kalender des Johannes Regiomontan, Nürnberg, um 1474. Faksimiledruck nach dem Exemplar der Preussischen Staatsbibliothek. Mit einer Einleitung von Ernst Zinner by Regiomontanus. Isis, 30(1):111–113, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225589.

Larsell:1939:BRB

[1140] O. Larsell. Book review: Atlantica by Rudbeck Olaus; Axel Nelson. Isis, 30(1):114–119, February 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225590.

Sakanishi:1939:BRB

[1141] Shio Sakanishi. Book review: Tô Sô jidai ni okeru Arab-jin no Shina Tsûshô no gaikyô; kotoni Sô matsu no Teikyo-shihaku Saiiki-jin Ho Jukô no jiseki by Jitsizô Kuwabara. Isis, 30(1):120–121, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225591.

Pagel:1939:BRB

[1142] W. Pagel. Book review: Unendliche Sphaere und Allmittelpunkt by Dietrich Mahnke. Isis, 30(1):121–124, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225592.

Lenzen:1939:BRE

[1143] V. F. Lenzen. Book review: The Evolution of Physics by Albert Einstein; Leopold Infeld. Isis, 30(1):124–125, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225593.

Speter:1939:BRB

[1144] Max Speter. Book review: Chemisch-Pharmazeutisches Bio- und Bibliographikon by Fritz Ferchl. Isis, 30(1):125–127, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225594.

Norman:1939:BRB

[1145] Daniel Norman. Book review: The Museum of Modern Art.– Photography, 1839–1937. Isis, 30(1):127–128, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225595.

Zirkle:1939:BRBa

[1146] Conway Zirkle. Book review: Genetics and the Origin of Species by Theodosius Dobzhansky. Isis, 30(1):128–131, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225596.

Zirkle:1939:BRBb

[1147] Conway Zirkle. Book review: Herbals. Their Origin and Evolution. A Chapter in the History of Botany 1470–1670 by Agnes Arber. Isis, 30 (1):131–132, February 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225597.

Hartner:1939:BRB

[1148] Willy Hartner. Book review: Die Schrift in Vergangenheit und Gegenwart by Hans Jensen. Isis, 30(1):132–137, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225598.

Sarton:1939:FFC

[1149] George Sarton and Frances Siegel. Fifty-fifth critical bibliography of the history and philosophy of science and of the history of civilization (to end of July 1938, — with special reference to sections 50 to 58). Isis, 30 (1):138–218, February 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225599.

Anonymous:1939:BMa

[1150] Anonymous. Back matter. Isis, 30(1):??, February 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225600.

Anonymous:1939:FMb

[1151] Anonymous. Front matter. Isis, 30(2):219–221, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226288.

Sarton:1939:SPV

[1152] George Sarton. Second preface to vol. xxx: Iconographic honesty. Isis, 30(2):222–235, May 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226289.

Davis:1939:APK

[1153] Tenney L. Davis and Chao Yün-Ts’ung. An alchemical poem by Kao Hsiang-Hsien. Isis, 30(2):236–240, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226290.

Seares:1939:GEH

[1154] Frederick H. Seares. George ellery hale: The scientist afield. Isis, 30(2): 241–267, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226291.

Dobell:1939:DU

[1155] Clifford Dobell. Dr. O. Uplavici (1887–1938). Isis, 30(2):268–272, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226292.

Ockenden:1939:NHL

[1156] R. E. Ockenden. A note on Hyoscyamus Luteus. Isis, 30(2):273–276, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226293.

Diller:1939:NC

[1157] Aubrey Diller. Notes and correspondence. Isis, 30(2):277–283, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226294.

Johnson:1939:BRB

[1158] Francis R. Johnson. Book review: The Three Voyages of Martin Frobisher by Martin Frobisher; Vilhjalmur Stefansson; Eloise McCaskill. Isis, 30(2):284–289, May 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226295.

Kofoid:1939:BRBa

[1159] Charles A. Kofoid. Book review: Sir Grafton Elliot Smith. (1871– 1937) by Warren R. Dawson. Isis, 30(2):289–291, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226296.

Sarton:1939:BRBb

[1160] George Sarton. Book review: La science arabe et son rôle dans l’évolution scientifique by Aldo Mieli; Henry-Paul-Joseph Renaud; Max Meyerhof; Julius Ruska. Isis, 30(2):291–295, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226297.

Ashley-Montagu:1939:BRBa

[1161] M. F. Ashley-Montagu. Book review: Architects of Ideas by Ernest R. Trattner. Isis, 30(2):295–297, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226298.

Morris:1939:BRB

[1162] Charles W. Morris. Book review: Organon.–International Review by Sanislaw Michalski. Isis, 30(2):297–298, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226299.

Davis:1939:BRBa

[1163] H. T. Davis. Book review: The Principles of Mathematics by Bertrand Russell. Isis, 30(2):298–302, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226300.

Davis:1939:BRBb

[1164] H. T. Davis. Book review: The Human Value of Biology by Johan Hjort. Isis, 30(2):302–304, May 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226301.

Kofoid:1939:BRBb

[1165] Charles A. Kofoid. Book review: Geschichte der Botanik von den ersten Anfängen bis zur Gegenwart by Martin Möbius. Isis, 30(2):304–306, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226302.

Sarton:1939:BRBc

[1166] George Sarton. Book review: La psychologie de la conversion chez les peuples non-civilisés by Raoul Allier. Isis, 30(2):306–310, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226303.

Ashley-Montagu:1939:BRBb

[1167] M. F. Ashley-Montagu. Book review: The Intelligent Individual and Society by P. W. Bridgman. Isis, 30(2):310–312, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226304.

Ashley-Montagu:1939:BRBc

[1168] M. F. Ashley-Montagu. Book review: The New Vision of Man by F. S. Marvin. Isis, 30(2):312–313, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226305.

Sarton:1939:FSCa

[1169] George Sarton and Frances Siegel. Fifty-sixth critical bibliography of the history and philosophy of science and the history of civilization (to November 1938 — with special reference to B.C. centuries). Isis, 30(2): 314–420, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226306.

Anonymous:1939:BMb

[1170] Anonymous. Back matter. Isis, 30(2):??, May 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226307.

Anonymous:1939:FMc

[1171] Anonymous. Front matter. Isis, 30(3):421–424, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225522.

Sarton:1939:TPV

[1172] George Sarton. Third preface to volume XXX: Comenius Redivivus. Isis, 30(3):425–451, August 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225523.

McColley:1939:UR

[1173] Grant McColley. The universe of de revolutionibus. Isis, 30(3):452– 472, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225524.

Kilgour:1939:TRC

[1174] Frederick G. Kilgour. Thomas robie (1689–1729), colonial scientist and physician. Isis, 30(3):473–490, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225525.

Falls:1939:BPB

[1175] William F. Falls. Buffon et les premières bêtes du jardin du roi: Histoire ou Légende? Isis, 30(3):491–494, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225526.

Sarton:1939:LIL

[1176] George Sarton. Lunar influences on living things. Isis, 30(3):495–507, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225527.

Loria:1939:NC

[1177] Gino Loria, G. S., Albert Lejeune, Bertha Frick, J. L. B. Engelhard, G. van Rijnberk, D. Schoute, Aldo Mieli, Arnold Reymond, O. Neugebauer, and R. C. A. Notes and correspondence. Isis, 30(3):508–519, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225528.

Diller:1939:BRBa

[1178] Aubrey Diller. Book review: L’Optique et la Catoptrique by Euclide. Isis, 30(3):520–521, August 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225529.

Kofoid:1939:BRBc

[1179] Charles A. Kofoid. Book review: The Mind of the Ancient World by H. N. Wethered. Isis, 30(3):521–523, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225530.

P:1939:BRB

[1180] A. P. Book review: De re aedificatoria libri decem by Leon Battista Alberti. Isis, 30(3):523–525, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225531.

Pogo:1939:BRBc

[1181] A. Pogo. Book review: The Discovery of Abyssinia by the Portuguese in 1530. A Facsimile of the Relation Entitled Carta das novas que vieram a el Rey nosso Senhor do descobrimento do preste Joham by Henry Thomas; Armando Cortesao. Isis, 30(3):525–529, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225532.

Cohen:1939:BRBa

[1182] I. Bernard Cohen. Book review: The Mathematical Work of John Wallis, D. D., F. R. S, (1616–1703) by J. F. Scott. Isis, 30(3):529–532, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225533.

Kofoid:1939:BRBd

[1183] Charles A. Kofoid. Book review: Man and Beast in French Thought of the Eighteenth Century by Hester Hastings. Isis, 30(3):532–533, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225534.

Kofoid:1939:BRBe

[1184] C. A. Kofoid. Book review: Voyage de Lapérouse sur les côtes de l’Alaska et de la Californie (1786) avec une introduction et des notes by Gilbert Chinard. Isis, 30(3):533–535, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225535.

Kofoid:1939:BRBf

[1185] Charles A. Kofoid. Book review: Birds of America by John James Audubon. Isis, 30(3):535–536, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225536.

Kofoid:1939:BRBg

[1186] Charles A. Kofoid. Book review: The Life and Convictions of William Sidney Thayer, Physician by Edith Gittings Reid. Isis, 30(3):536–538, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225537.

Kofoid:1939:BRBh

[1187] Charles A. Kofoid. Book review: The Conquest of the North Pole by J. Gordon Hayes. Isis, 30(3):538–539, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225538.

Diller:1939:BRBb

[1188] Aubrey Diller. Book review: Wesen und Entstehung des Sexus im Denken der Antike by Konrad Blersch. Isis, 30(3):539–540, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225539.

Durand:1939:BRBb

[1189] Dana B. Durand. Book review: Geography in the Middle Ages by George H. T. Kimble. Isis, 30(3):540–542, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225540.

Ashley-Montagu:1939:BRBd

[1190] M. F. Ashley-Montagu. Book review: The Gilding Process and the Metallurgy of Copper and Lead among the Pre-Columbian Indians by Paul Bergsoe. Isis, 30(3):542–543, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225541.

Kofoid:1939:BRBi

[1191] Charles A. Kofoid. Book review: Cartography of the Northwest Coast of America to the Year 1800 by Henry H. Wagner. Isis, 30(3):543–545, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225542.

Kofoid:1939:BRBj

[1192] Charles A. Kofoid. Book review: The Story of Human Error by Joseph Jastrow. Isis, 30(3):545–547, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225543.

Ashley-Montagu:1939:BRBe

[1193] M. F. Ashley-Montagu. Book review: A Basis of Opinion by Adrian Coates. Isis, 30(3):547–549, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225544.

Kofoid:1939:BRBk

[1194] Charles A. Kofoid. Book review: Genetics and the Origin of Species by Theodosius Dobzhansky. Isis, 30(3):549–551, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225545.

Zirkle:1939:BRBc

[1195] Conway Zirkle. Book review: The American Plant Migration: I. The Potato by Berthold Laufer. Isis, 30(3):551–552, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225546.

Kofoid:1939:BRBl

[1196] C. A. Kofoid. Book review: The Emergence of Human Culture by Carl J. Warden. Isis, 30(3):552–553, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225547.

Ashley-Montagu:1939:BRBf

[1197] M. F. Ashley-Montagu. Book review: The Mind of Primitive Man by Franz Boas. Isis, 30(3):553–554, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225548.

Ashley-Montagu:1939:BRBg

[1198] M. F. Ashley-Montagu. Book review: Towards the Twentieth Century by H. V. Routh. Isis, 30(3):554–555, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225549.

Merton:1939:BRB

[1199] Robert K. Merton. Book review: Political Arithmetic: A Symposium of Population Studies by Lancelot Hogben. Isis, 30(3):555–557, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225550.

Ashley-Montagu:1939:BRBh

[1200] M. F. Ashley-Montagu. Book review: The Culture of Cities by Lewis Mumford. Isis, 30(3):557–558, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225551.

Ashley-Montagu:1939:BRBi

[1201] M. F. Ashley-Montagu. Book review: Inquisition and Liberty by G. G. Coulton. Isis, 30(3):558–560, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225552.

Simpson:1939:BRB

[1202] Lesley Byrd Simpson. Book review: Historia del Protomedicato de Buenos Aires by Juan Ramón Beltrán. Isis, 30(3):560–561, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225553.

Guttentag:1939:BRB

[1203] Otto E. Guttentag. Book review: Geschichte der Gesellschaft der Aerzte in Wien 1837–1937. Isis, 30(3):561–562, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225554.

Mayer:1939:AP

[1204] Joseph Mayer and Henry R. Viets. Administrative papers. Isis, 30(3): 563–572, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225555.

Anonymous:1939:VIb

[1205] Anonymous. Volume information. Isis, 30(3):573–642, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225521.

Anonymous:1939:BTC

[1206] Anonymous. Brief table of contents of vol. XXX (1939). Isis, 30(3):643– 645, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225556.

Anonymous:1939:BMc

[1207] Anonymous. Back matter. Isis, 30(3):??, August 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225557.

Anonymous:1939:VIc

[1208] Anonymous. Volume information. Isis, 31(1):1, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226010.

Anonymous:1939:FMd

[1209] Anonymous. Front matter. Isis, 31(1):3–5, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226011.

Sarton:1939:PVT

[1210] George Sarton. Preface to volume thirty-one: Quo usque tandem? (Latin) [How long?]. Isis, 31(1):6–7, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226012.

Sayili:1939:WIS

[1211] Aydin M. Sayili. Was ibni sina an Iranian or a turk? Isis, 31(1):8–24, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226013.

Jager:1939:AAD

[1212] Robert Jager and Boyd C. Patterson. The artificial arithmetick in decimals of Robert Jager (London, 1651). Isis, 31(1):25–31, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226014.

Stimson:1939:ASC

[1213] Dorothy Stimson. Amateurs of science in 17th Century England. Isis, 31(1):32–47, November 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226015.

Crommelin:1939:AHJ

[1214] C. A. Crommelin. Eine astronomische Handschrift aus dem 17. Jahrhundert. (German) [An astronomical manuscript from the 17th Century]. Isis, 31(1):48–50, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226016.

Winter:1939:SNE

[1215] H. J. J. Winter. Scientific notes from the early minutes of the Peterborough Society 1730–1745. Isis, 31(1):51–59, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226017.

Bigelow:1939:HWT

[1216] Robert P. Bigelow. Harry Walter Tyler. Isis, 31(1):60–64, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226018.

Jones:1939:PSC

[1217] Richard F. Jones. Puritanism, science, and Christ Church. Isis, 31(1):65– 67, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226019.

Thorndike:1939:HUC

[1218] Lynn Thorndike. A hitherto unnoticed criticism of astrology: Liber de reprobatione iudiciorum astrologiae. Isis, 31(1):68–78, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226020.

S:1939:NC

[1219] G. S. Notes and correspondence. Isis, 31(1):79–82, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226021.

Edelstein:1939:BRB

[1220] Ludwig Edelstein. Book review: Theophrastea, Studien zur botanischen Begriffsbildung by Reinhold Strömberg. Isis, 31(1):83–85, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226022.

Kofoid:1939:BRBm

[1221] Charles A. Kofoid. Book review: Phineas Fletcher, Man of Letters, Science, and Divinity by Abram Barnett Langdale. Isis, 31(1):86–87, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226023.

Goldsmith:1939:BRB

[1222] Albert Robin Goldsmith. Book review: The Rôle of Scientific Societies in the Seventeenth Century by Martha Ornstein. Isis, 31(1):87–89, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226024.

Saunders:1939:BRB

[1223] John B. de C. M. Saunders. Book review: Observationes anatomicae selectiores, Amstelodamensium. 1667–1673 by F. J. Cole. Isis, 31(1):90– 91, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226025.

Cohen:1939:BRBb

[1224] I. Bernard Cohen. Book review: Benjamin Franklin (1706–1790) by Carl Van Doren. Isis, 31(1):91–94, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226026.

Jackson:1939:BRB

[1225] Dugald C. Jackson. Book review: A Hundred Years of Mechanical Engineering by Edward Cressy. Isis, 31(1):94–95, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226027.

Sarton:1939:BRBd

[1226] George Sarton. Book review: Les missions scientifiques à l’époque de Muh.ammad ’Alî, de ’Abbâs al-Awwal et de Sa’îd by Le Prince Omar Toussoun. Isis, 31(1):95–99, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226028.

Gandz:1939:BRBa

[1227] Solomon Gandz. Book review: La Médecine en Assyrie et en Babylonie by Georges Contenau. Isis, 31(1):99–101, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226029.

Gandz:1939:BRBb

[1228] Solomon Gandz. Book review: Medieval Number Symbolism by Vincent F. Hopper. Isis, 31(1):101–103, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226030.

Pogo:1939:BRBd

[1229] A. Pogo. Book review: Decimal Tables for the Reduction of Hindu Dates from the Data of the Surya-Siddhanta by W. E. Van Wijk. Isis, 31(1): 103–104, November 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226031.

Macdonald:1939:BRB

[1230] Duncan B. Macdonald. Book review: Islam in the World by Zaki Ali. Isis, 31(1):104–105, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226032.

Ashley-Montagu:1939:BRBj

[1231] M. F. Ashley-Montagu. Book review: What Science Really Means by Julius W. Friend; James Feibleman. Isis, 31(1):105–108, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226033.

Hellman:1939:BRB

[1232] C. Doris Hellman. Book review: Le matematiche nella storia e nella cultura by Federigo Enriques. Isis, 31(1):108–109, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226034.

Bobrovnikoff:1939:BRB

[1233] N. T. Bobrovnikoff. Book review: A Hundred Years of Astronomy by Reginald L. Waterfield. Isis, 31(1):109–112, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226035.

Kofoid:1939:BRBn

[1234] Charles A. Kofoid. Book review: March of the Iron Men by Roger Burlingame. Isis, 31(1):112–113, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226036.

Kofoid:1939:BRBo

[1235] Charles A. Kofoid. Book review: Swedish Contributions to the Development of Plant Breeding by Åke Åkerman; I. Granhall; G. Nilsson-Leissner; A. Müntzing; O. Tedin. Isis, 31(1):113–114, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226037.

Kofoid:1939:BRBp

[1236] Charles A. Kofoid. Book review: Sixty Years of Botany in Great Britain (1875–1935) by F. O. Bower. Isis, 31(1):114–115, November 1939. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226038.

Kofoid:1939:BRBq

[1237] Charles A. Kofoid. Book review: A Bibliography of Eastern Asiatic Botany by Elmer D. Merrill; Egbert H. Walker. Isis, 31(1):115–116, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226039.

Kofoid:1939:BRBr

[1238] C. A. Kofoid. Book review: Feeling and Emotion. A History of Theories by H. M. Gardiner; Ruth Clark Metcalf; John G. Beebe-Center. Isis, 31 (1):116–117, November 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226040.

Sarton:1939:BR

[1239] George Sarton. Isis report for 1938. Isis, 31(1):118–120, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226041.

Sarton:1939:FSCb

[1240] George Sarton and Frances Siegel. Fifty-seventh critical bibliography of the history and philosophy of science and of the history of civilization (to May 1939 — with special reference to cent. I to VII inclusive). Isis, 31 (1):121–298, November 1939. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226042.

Anonymous:1939:BMd

[1241] Anonymous. Back matter. Isis, 31(1):??, November 1939. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

      www.jstor.org/stable/226043.